### v1

# An algebraic theory of Łojasiewicz exponents

Abstract.
We develop a unified algebraic and valuative theory of Łojasiewicz exponents for pairs of graded families and filtrations of ideals. Within this framework, local Łojasiewicz exponents, gradient exponents, and exponents at infinity are all realized as asymptotic containment thresholds between filtrations, governed by integral closure. This reformulation shows that Łojasiewicz exponents are fundamentally valuative optimization problems.
The central structural contribution of the paper is a finite–max principle. Under verifiable algebraic hypotheses, the a priori infinite valuative supremum bounding the Łojasiewicz exponent reduces to a finite maximum, and computes the Łojasiewicz exponent precisely. We identify two complementary mechanisms leading to this phenomenon: finite testing arising from normalized blowups and Noetherian Rees algebras, and attainment via compactness of normalized valuation spaces under linear boundedness assumptions.
This finite–max framework yields strong structural consequences. We prove rigidity results showing that common extremal valuations force equality of Łojasiewicz ratios, and we establish stratification and stability phenomena for Łojasiewicz exponents in families, including fractional linearity and wall-chamber behavior along natural one–parameter deformations.
The theory recovers and explains classical results in toric and Newton–polyhedral settings, particularly, for Newton nondegenerate case, where the Łojasiewicz exponent is computed by finitely many toric divisorial valuations corresponding to facet data.
Finally, we illustrate why the hypotheses underlying the finite–max principle are essential, delineating the precise scope of the theory.
Key words and phrases:
Łojasiewicz exponent, singularity, integral closure, valuation, ideal containment
2020 Mathematics Subject Classification:
13H15, 32B10, 32S05, 14B05, 58K30
Part I Motivations, Goals and Statement of Main Results
We start the paper by describing motivation, conceptual background, and an overview of the main results.
1. Introduction
The Łojasiewicz exponent is a fundamental invariant measuring the rate at which one analytic object dominates another near a singular locus. More precisely, given analytic function germs
$f_{1},\dots,f_{s},\;g_{1},\dots,g_{t}\in\mathcal{O}_{\mathbb{C}^{d},0},$
the classical Łojasiewicz inequality [39] asserts that there exist constants $C>0$, $\theta>0$, and a neighborhood $U$ of the origin such that
$\max_{1\leq j\leq t}|g_{j}(x)|^{\theta}\;\leq\;C\max_{1\leq i\leq s}|f_{i}(x)|\quad\text{for all }x\in U.$
The infimum of such exponents $\theta$ is called the (analytic) Łojasiewicz exponent of $(f_{1},\dots,f_{s})$ with respect to $(g_{1},\dots,g_{t})$.
Introduced originally by Łojasiewicz in the study of real and complex analytic sets, this exponent captures subtle asymptotic information about singularities: it quantifies relative vanishing rates, governs gradient inequalities for isolated critical points, controls stability and equisingularity phenomena, and plays a central role in the study of global polynomial mappings and optimization problems. Several important variants arise naturally, including the gradient Łojasiewicz exponent of hypersurfaces with isolated singularities (cf. [8, 13, 20, 37]) and the Łojasiewicz exponent at infinity for polynomial mappings (cf. [14, 23, 53, 54]). Despite their close conceptual relationship, these invariants have traditionally been studied using markedly different techniques, ranging from analytic arc methods and curve selection to Newton polyhedra and projective compactifications.
A key unifying insight already appears in the classical work of Lejeune-Jalabert and Teissier [36]. They showed that analytic Łojasiewicz inequalities are equivalent to integral–closure containments between powers of the corresponding ideals. Particularly, if $\mathfrak{a}=(f_{1},\dots,f_{s})$ and $\mathfrak{b}=(g_{1},\dots,g_{t})$, then inequalities of the form above hold if and only if there exist integers $p,q$ such that
$\mathfrak{b}^{q}\subseteq\overline{\mathfrak{a}^{p}}.$
This result places the analytic Łojasiewicz exponent squarely within the framework of ideal containment problems and reveals its intrinsic valuative nature; see also [7, 52]. This viewpoint has been generalized by Bivià-Ausina [1], Bivià-Ausina–Encinas [2, 3] and Bivià-Ausina–Fukui [4] in their studies of Łojasiewicz exponent of ideals via integral closures and mixed multiplicites.
Independently, ideal containment problems have played a central role in commutative algebra and algebraic geometry; notably, the uniform symbolic–ordinary containment theorems of Ein–Lazarsfeld–Smith [15] and Hochster–Huneke [27], together with generalization by Ma–Schwede [40]. Motivated by the fact that such containments are often not sharp, Harbourne and Huneke [25] proposed improved containment conjectures, leading to a large body of work during the last fifteen years; particularly, the introduction of numerical invariants measuring failure of containment, namely, the resurgence, introduced by Bocci–Harbourne [6] and further developed by Guardo–Harbourne–Van Tuyl [19]. Morerecently, these ideas have been extended to graded families and filtrations of ideals, allowing symbolic powers, integral closures, and other asymptotic constructions to be treated in a unified way; see [21].
The purpose of this paper is to bring these two threads together. We develop a unified algebraic and valuative theory of Łojasiewicz exponents for pairs of graded filtrations of ideals, in which local, gradient, and at–infinity exponents all arise as asymptotic containment thresholds. This reformulation places classical Łojasiewicz inequalities into a single containment framework and reveals that Łojasiewicz exponents are governed by valuative optimization problems. In general, Łojasiewicz exponent is bounded by a supremum of asymptotic valuative ratios associated to the underlying filtrations. While such variational descriptions are a priori infinite, a central theme of this paper is that, under verifiable algebraic or geometric hypotheses, this supremum is in fact a maximum attained by finitely many valuations and computes the Łojasiewicz exponent. Identifying when this finite–max phenomenon occurs and exploiting structural consequences, including rigidity, stability, stratification and wall-chamber behaviors, is the main objective of this paper.
2. Goals and statement of main results
This paper develops a unified algebraic and valuative theory of Łojasiewicz exponents. It has two overarching aims. The first aim is conceptual: to introduce a general notion of algebraic Łojasiewicz exponent attached to a pair of graded filtrations of ideals and show that all classical analytic Łojasiewicz exponents (local, gradient, and at infinity) arise uniformly as instances of this invariant. From this perspective, Łojasiewicz exponents are interpreted as asymptotic containment thresholds between filtrations, placing them within the framework of integral closure and ideal containment theory. The second aim is structural: to show that valuations are the computational mechanism governing these thresholds, and to isolate verifiable hypotheses under which the resulting valuative optimization problems reduce to finite maxima over explicit candidate sets. This finite–max philosophy underlies the rigidity, stability, and polyhedral phenomena developed later in the paper.
2.1. From analytic inequalities to algebraic thresholds
After fixing analytic conventions in Section 3, we introduce in Section 4 the algebraic Łojasiewicz exponent $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$ attached to a pair of filtrations ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ and ${\mathfrak{b}_{\bullet}}=\{\mathfrak{b}_{q}\}_{q\geq 1}$. By definition, $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$ records the smallest asymptotic slope at which the containments
$\mathfrak{b}_{qt}\subseteq\overline{\mathfrak{a}_{pt}}\quad\text{hold for }t\gg 1.$
The key point is that, for families of ordinary powers of two ideals, this algebraic invariant exactly recovers the analytic one (Theorems 4.5 and 4.9). Consequently, local, gradient, and at–infinity Łojasiewicz exponents all arise as instances of $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$, distinguished only by the choice of filtrations, by the class of valuations allowed to appear, and possibly by inverting the invariant.
2.2. Valuative formulas and the finite–max principle
A central feature of the algebraic Łojasiewicz exponent $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$ is that it admits a valuative description (Section 5). Using the valuative criterion for integral closure, we show that $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$ is bounded below by a supremum of ratios
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\;\geq\;\sup_{v}\frac{v({\mathfrak{a}_{\bullet}})}{v({\mathfrak{b}_{\bullet}})},$
where $v$ ranges over valuations centered at $\mathfrak{m}$ (Theorem 5.6).
A fundamental question is whether this supremum is in fact a maximum, and whether it computes the Łojasiewicz exponent. We isolate two independent mechanisms that force such a finite–max phenomenon. The first is a finite testing principle: under suitable Noetherian hypotheses, containment of integral closures can be tested on a fixed finite set of divisorial valuations arising from a normalized Rees algebra or a fixed birational model (Theorems 5.8 and 5.9). The second mechanism relies on compactness properties of normalized valuation spaces together with uniform convergence of the functions $v\mapsto v(\mathfrak{a}_{p})/p$ and $v\mapsto v(\mathfrak{b}_{q})/q$, which ensures attainment of the supremum even in the absence of finite generation (Theorem 5.14). We also illustrate that considering only divisorial valuations may not be enough to realize the supremum in general (Theorem 5.16).
2.3. Rigidity, stability, and stratification
When the finite–max principle applies, the valuative formula for $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$ provides an effective description of the invariant, and the extremal valuation(s) attaining the maximum become geometrically meaningful data. A further contribution of the paper is to show that these extremal valuations govern the behavior of Łojasiewicz exponents in families.
We prove that common extremal valuations force rigidity of Łojasiewicz ratios and that, under finite–max hypotheses, parameter spaces admit a natural stratification according to the valuation computing the exponent (Theorems 8.1 and 9.9). Along one–parameter families, this yields fractional linearity of $1/\mathcal{L}$ and shows that affineness forces persistence of the extremal valuation (Theorem 9.14, Corollary 9.16). These results provide a unified algebraic explanation for stability and wall–chamber behavior observed in analytic and polyhedral settings.
2.4. Polyhedral models and explicit computations
We describe explicit formulas and finite computation methods in combinatorial situations. For monomial and toric ideals, the exponent is determined by polyhedral data and toric valuations (Section 10). In these situations, integral–closure containments are encoded by linear inequalities, toric divisorial valuations correspond to linear functionals, and the valuative optimization problem reduces to a finite maximization over facet data of Newton polyhedra (Theorems 10.18 and 10.11).
We also discuss in more details how Łojasiewicz exponent at infinity behaves in our algebraic framework. We prove valuative and Rees–algebraic formulas at infinity (Theorem 11.3), and obtain finite computation and Newton polyhedral formulas in the nondegenerate case (Theorems 11.10 and 11.7).
2.5. Uniformity and necessity
The final part of the paper illustrates why the hypotheses used to obtain finite testing and finite–max formulas are essential (Theorems 12.1 and 12.3). Outside situations admitting verifiable finite testing, the set of valuations computing $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$ need not admit any a priori finite reduction, delineating the precise scope of the theory developed here (Theorem 12.4).
[…]
Part II Łojasiewicz Exponents and the Algebraic Framework
In this part, we develop the algebraic framework underlying Łojasiewicz exponents. We begin by recalling analytic Łojasiewicz exponents and their standard variants, and then generalize these invariants to a purely algebraic notion attached to graded families and filtrations of ideals. This perspective allows Łojasiewicz exponents to be considered in a more general context, interpreted as asymptotic containment thresholds, and sets the stage for the valuative analysis carried out in the subsequent parts of the paper.
3. Analytic Łojasiewicz exponents
3.1. Background
In this section we briefly fix the analytic setting and conventions for Łojasiewicz exponents that will be used throughout the paper. Our goal is not to survey the analytic theory, but to record a formulation that is compatible with the algebraic containment framework developed later.
Given two systems of analytic function germs near a point, Łojasiewicz-type inequalities compare their relative rates of vanishing. Several equivalent formulations appear in the literature, differing mainly by normalization. We adopt a convention that aligns the analytic exponent with integral–closure containments and the algebraic Łojasiewicz exponent defined in Section 4.
3.2. Analytic definition
Let
$\mathfrak{a}=(f_{1},\ldots,f_{s}),\quad\mathfrak{b}=(g_{1},\ldots,g_{t})\subseteq\mathcal{O}_{\mathbb{C}^{d},0},$
where $0\in f_{i}^{-1}(0)\cap g_{j}^{-1}(0)$ for all $i,j$. Define
$\|f(x)\|:=\max_{1\leq i\leq s}|f_{i}(x)|,\quad\|g(x)\|:=\max_{1\leq j\leq t}|g_{j}(x)|.$
Definition 3.1 (Analytic Łojasiewicz exponent).
The analytic Łojasiewicz exponent of $\mathfrak{a}$ with respect to $\mathfrak{b}$ is
$\mathcal{L}^{an}_{\mathfrak{b}}(\mathfrak{a}):=\inf\left\{\theta>0:\exists\,U\ni 0,\ \exists\,C>0\ \text{such that}\ C\|f(x)\|\geq\|g(x)\|^{\theta}\ \forall x\in U\right\}.$
Remark 3.2 (Convention).
The inequality is written with $\geq$”, so that larger values of $\mathcal{L}^{an}_{\mathfrak{b}}(\mathfrak{a})$ correspond to stronger domination of (the vanishing of) $\mathfrak{a}$ by $\mathfrak{b}$ near the origin. Other formulations in the literature may also differ by inversion of the exponent.
3.3. The gradient Łojasiewicz exponent
Let $f:(\mathbb{C}^{d},0)\to(\mathbb{C},0)$ be a holomorphic function with an isolated critical point at the origin, and let $J(f)$ denote its Jacobian ideal. The gradient Łojasiewicz exponent of $f$ at 0, usually denoted by $L(f,0)$, is the optimal (infimum) value $\eta>0$ such that
$\|\nabla f(x)\|\geq C\|x\|^{\eta}$
for $x$ near 0. In light of Definition 3.1, the gradient Łojasiewicz exponent of $f$ at 0 is given by $L(f,0)=\mathcal{L}^{an}_{\mathfrak{m}}(J(f)).$
3.4. The Łojasiewicz exponent at infinity
In addition to local exponents, one also studies Łojasiewicz inequalities at infinity for polynomial mappings
$F:\mathbb{C}^{n}\to\mathbb{C}^{m},$
typically of the form
$\|x\|^{\alpha}\leq C\|F(x)\|\quad\text{for }\|x\|\gg 1,$
or variants involving $\mathrm{dist}(x,F^{-1}(0))$. The optimal (i.e., supremum) exponent $\alpha$ in such an inequality is called the Łojasiewicz exponent of $F$ at infinity, and denoted by $L_{\infty}(F)$.
Although inequalities at infinity are global on $\mathbb{C}^{n}$, their meaning is local after projective compactification: the condition $\|x\|\to\infty$ corresponds to approaching the divisor at infinity, and the growth of $F$ is encoded by the relative vanishing orders of the source and target defining functions along the boundary.
4. Algebraic Łojasiewicz exponents of filtrations
In this section we introduce the algebraic Łojasiewicz exponent $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$ attached to a pair of filtrations ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ and ${\mathfrak{b}_{\bullet}}=\{\mathfrak{b}_{q}\}_{q\geq 1}$ of ideals in a Noetherian ring, viewed as the asymptotic containment threshold for $\mathfrak{b}_{qt}\subseteq\overline{\mathfrak{a}_{pt}}$ for $t\gg 1$ (Definition 4.3). We show that this invariant recovers the classical analytic exponent in the local analytic setting (Theorem 4.5). Finally, we package Łojasiewicz exponents at infinity into the same algebraic framework by working with the canonical growth filtrations induced along the boundary divisor of a compactification, leading to an equivalence for the local and global “at infinity” invariants in Theorem 4.9.
4.1. Graded families and Łojasiewicz exponent
The algebraic definition naturally involves graded families and filtrations of ideals, encompassing powers, integral closures, symbolic powers, and growth filtrations arising in geometric contexts. Throughout this section, $R$ denotes a Noetherian ring.
Definition 4.1 (Graded family and filtration).
A graded family of ideals ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ in $R$ is a collection of proper ideals satisfying
$\mathfrak{a}_{p}\mathfrak{a}_{p^{\prime}}\subseteq\mathfrak{a}_{p+p^{\prime}}\quad\text{for all }p,p^{\prime}\geq 1.$
A graded family ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ is called a filtration if
$\mathfrak{a}_{p+1}\subseteq\mathfrak{a}_{p}\quad\text{for all }p\geq 1.$
Remark 4.2.
All graded families appearing naturally in this paper — including powers of ideals, integral closures of powers, symbolic powers — are filtrations. By ideals, we always mean proper (non-unit) ideals.
The Łojasiewicz exponent records the smallest asymptotic slope at which one filtration is eventually contained in another.
Definition 4.3 (Algebraic Łojasiewicz exponent for filtrations).
Let ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ and ${\mathfrak{b}_{\bullet}}=\{\mathfrak{b}_{q}\}_{q\geq 1}$ be filtrations of ideals in $R$. The algebraic Łojasiewicz exponent of $\mathfrak{a}_{\bullet}$} with respect to $\mathfrak{b}_{\bullet}$} is
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\;:=\;\inf\left\{\frac{q}{p}\in\mathbb{Q}_{>0}\ :\ \mathfrak{b}_{qt}\subseteq\overline{\mathfrak{a}_{pt}}\text{ for }t\gg 1\right\}\ \in\ \mathbb{R}_{\geq 0}\cup\{\infty\},$
with the convention that $\inf\emptyset=\infty$. For ideals $\mathfrak{a},\mathfrak{b}\subseteq R$ we write $\mathcal{L}_{\mathfrak{b}}(\mathfrak{a})$ for $\mathcal{L}_{\{\mathfrak{b}^{q}\}_{q\geq 1}}\bigl(\{\mathfrak{a}^{p}\}_{p\geq 1}\bigr)$.
4.2. Algebraic and analytic Łojasiewicz exponents
We now compare the algebraic Łojasiewicz exponent with its analytic counterpart in the local analytic setting. As before, assume $R=\mathcal{O}_{\mathbb{C}^{d},0}$, $\mathfrak{a}=(f_{1},\dots,f_{s}),\mathfrak{b}=(g_{1},\dots,g_{t})$ where $0\in f_{i}^{-1}(0)\cap g_{j}^{-1}(0)$ for any $i$ and $j$, and $\|f(x)\|:=\max_{i}|f_{i}(x)|,\quad\|g(x)\|:=\max_{j}|g_{j}(x)|.$
The comparison rests on the following analytic criterion for integral closure stated by Lejeune-Jalabert–Teissier [36].
Lemma 4.4 (Lejeune-Jalabert–Teissier).
Let $p,q\in\mathbb{N}$. The following are equivalent:
  1. (1)
    $\mathfrak{b}^{q}\subseteq\overline{\mathfrak{a}^{p}}$.
  2. (2)
    There exist a neighborhood $U$ of 0 and a constant $C>0$ such that
    $\|g(x)\|^{q}\leq C\|f(x)\|^{p}\quad\text{for all }x\in U.$
Proof.
This is the Lejeune-Jalabert–Teissier analytic criterion for integral closure; see [36, Théorème 7.2]. For the reader’s convenience and to fix normalization conventions, we include a self-contained argument in the appendix. ∎
Our first result shows that, for two ideals $\mathfrak{a}$ and $\mathfrak{b}$, the analytic and algebraic Łojasiewicz exponents agree.
Theorem 4.5.
We have $\mathcal{L}^{an}_{\mathfrak{b}}(\mathfrak{a})=\mathcal{L}_{\mathfrak{b}}(\mathfrak{a}).$
Proof.
Let $\theta>0$ be admissible for $\mathcal{L}^{an}_{\mathfrak{b}}(\mathfrak{a})$ (see Definition 3.1). Then, there exist a neighborhood $U$ of 0 and a constant $C>0$ such that
$\|g(x)\|^{\theta}\leq C\|f(x)\|\quad\text{for all }x\in U.$
Shrinking $U$ if necessary, we may assume $\|g(x)\|\leq 1$ on $U$. Fix $p\in\mathbb{N}$ and set $q=\lceil p\theta\rceil$. Raising the above inequality to the $p$-th power yields
$\|g(x)\|^{p\theta}\leq C^{p}\|f(x)\|^{p}\quad(x\in U).$
Since $q\geq p\theta$ and $\|g(x)\|\leq 1$, we have $\|g(x)\|^{q}\leq\|g(x)\|^{p\theta}$, and so
$\|g(x)\|^{q}\leq C^{p}\|f(x)\|^{p}\quad(x\in U).$
By Lemma 4.4, this implies $\mathfrak{b}^{q}\subseteq\overline{\mathfrak{a}^{p}}.$ It follows that, for any $t\in\mathbb{N}$, we have
$\mathfrak{b}^{qt}\subseteq\left(\overline{\mathfrak{a}^{p}}\right)^{t}\subseteq\overline{\mathfrak{a}^{pt}}.$
Thus, $\frac{q}{p}$ is admissible for $\mathcal{L}_{\mathfrak{b}}(\mathfrak{a})$. Therefore,
$\mathcal{L}_{\mathfrak{b}}(\mathfrak{a})\leq\frac{q}{p}<\theta+\frac{1}{p}.$
By replacing $\frac{q}{p}$ with $\frac{qt}{pt}$, which is also admissible for $\mathcal{L}_{\mathfrak{b}}(\mathfrak{a})$, and letting $t\rightarrow\infty$ (so $pt\rightarrow\infty$), we conclude that $\mathcal{L}_{\mathfrak{b}}(\mathfrak{a})\leq\theta$. Taking the infimum over admissible $\theta$ gives
$\mathcal{L}_{\mathfrak{b}}(\mathfrak{a})\leq\mathcal{L}^{an}_{\mathfrak{b}}(\mathfrak{a}).$
Conversely, let $\frac{q}{p}\in\mathbb{Q}_{>0}$ be admissible for $\mathcal{L}_{\mathfrak{b}}(\mathfrak{a})$. Then, for $t\gg 1$, we have $\mathfrak{b}^{qt}\subseteq\overline{\mathfrak{a}^{pt}}.$ By Lemma 4.4, there exist a neighborhood $U$ of 0 and a constant $C>0$ such that
$\|g(x)\|^{qt}\leq C\|f(x)\|^{pt}\quad\text{for all }x\in U.$
Taking $(pt)$-th roots gives
$\|g(x)\|^{q/p}\leq C^{1/(pt)}\|f(x)\|\quad(x\in U).$
Hence, $\theta=\frac{q}{p}$ is admissible for $\mathcal{L}^{an}_{\mathfrak{b}}(\mathfrak{a})$, and so $\mathcal{L}^{an}_{\mathfrak{b}}(\mathfrak{a})\leq\frac{q}{p}$. Taking the infimum over admissible $\frac{q}{p}$ yields
$\mathcal{L}^{an}_{\mathfrak{b}}(\mathfrak{a})\leq\mathcal{L}_{\mathfrak{b}}(\mathfrak{a}).$
This completes the proof. ∎
Observe, in particular, that when $\mathfrak{a}$ and $\mathfrak{b}$ are $\mathfrak{m}$–primary, both the analytic and algebraic Łojasiewicz exponents are finite and governed by integral-closure containments.
Example 4.6.
Let $R=\mathbb{C}[[x_{1},\dots,x_{n}]]$, $\mathfrak{m}=(x_{1},\dots,x_{n})$ and consider
$\mathfrak{a}=(x_{1}^{\alpha_{1}},\dots,x_{n}^{\alpha_{n}})\text{ and }\mathfrak{b}=\mathfrak{m}.$
It can be seen that the inequality
$\|(x_{1},\dots,x_{n})\|^{\theta}\leq C\max(|x_{1}|^{\alpha_{1}},\dots,|x_{n}|^{\alpha_{n}})$
fails for $\theta<\max\{\alpha_{1},\dots,\alpha_{n}\}$ along the coordinate axis, and holds for $\theta=\max\{\alpha_{1},\dots,\alpha_{n}\}$. Thus, by definition, $\mathcal{L}^{an}_{\mathfrak{b}}(\mathfrak{a})=\max\{\alpha_{1},\dots,\alpha_{n}\}.$
On the other hand, one can check on pure powers of the $x_{i}$’s that
$\mathfrak{m}^{qt}\subseteq\overline{\mathfrak{a}^{pt}}\text{ for }t\gg 1\Longleftrightarrow\frac{q}{p}\geq\max\{\alpha_{1},\dots,\alpha_{n}\}.$
That is,
$\mathcal{L}_{\mathfrak{b}}(\mathfrak{a})=\max\{\alpha_{1},\dots,\alpha_{n}\}=\mathcal{L}^{an}_{\mathfrak{b}}(\mathfrak{a}).$
4.3. Exponents at infinity in the algebraic framework
We shall explore an algebraic formulation of Łojasiewicz exponents at infinity, understand precisely how the analytic condition $\|x\|\gg 1$ is encoded by divisors at infinity on a suitable compactification, and show that the resulting analytic invariants agree with the algebraic containment invariants used in our valuative framework.
Let $F=(f_{1},\dots,f_{m}):\mathbb{A}^{n}_{\mathbb{C}}\to\mathbb{A}^{m}_{\mathbb{C}}$ be a polynomial mapping. We fix the standard projective compactifications $\mathbb{A}^{n}_{\mathbb{C}}\hookrightarrow\mathbb{P}^{n}_{\mathbb{C}}$ and $\mathbb{A}^{m}_{\mathbb{C}}\hookrightarrow\mathbb{P}^{m}_{\mathbb{C}}$, with hyperplanes at infinity
$H_{X}:=\{Z_{0}=0\}\subseteq\mathbb{P}^{n}_{\mathbb{C}}\text{ and }H_{Y}:=\{W_{0}=0\}\subseteq\mathbb{P}^{m}_{\mathbb{C}},$
where $Z_{0},\dots,Z_{n}$ and $W_{0},\dots,W_{m}$ denote the homogeneous coordinates of $\mathbb{P}^{n}_{\mathbb{C}}$ and $\mathbb{P}^{m}_{\mathbb{C}}$, respectively. Let $\Gamma_{F}\subseteq\mathbb{A}^{n}_{\mathbb{C}}\times\mathbb{A}^{m}_{\mathbb{C}}$ be the graph of $F$ and let $\overline{\Gamma}_{F}\subseteq\mathbb{P}^{n}_{\mathbb{C}}\times\mathbb{P}^{m}_{\mathbb{C}}$ be the normalization of the Zariski closure of $\Gamma_{F}$. Let
$\pi:\overline{\Gamma}_{F}\to\mathbb{P}^{n}_{\mathbb{C}}\text{ and }\rho:\overline{\Gamma}_{F}\to\mathbb{P}^{m}_{\mathbb{C}}$
be the natural projections and set
$D_{X}:=\pi^{-1}(H_{X})\text{ and }D_{Y}:=\rho^{-1}(H_{Y}).$
Denote by $\mathscr{I}_{X}:=\mathcal{I}_{D_{X}}$ and $\mathscr{I}_{Y}:=\mathcal{I}_{D_{Y}}$ the corresponding ideal sheaves of $D_{X}$ and $D_{Y}$ on $\overline{\Gamma}_{F}$. Approaching infinity in the source along the graph of $F$ corresponds precisely to approaching $D_{X}$ inside $\overline{\Gamma}_{F}$.
What does $\|x\|\gg 1$ mean near a boundary point. Fix $\xi\in D_{X}$. Choose indices $i\in\{1,\dots,n\}$ and $\ell\in\{0,1,\dots,m\}$ such that $\pi(\xi)\in\{Z_{i}\neq 0\}$ and $\rho(\xi)\in\{W_{\ell}\neq 0\}$. Define holomorphic functions near $\xi$ by
$u_{X}:=\Big(\frac{Z_{0}}{Z_{i}}\Big)\circ\pi\text{ and }u_{Y}:=\Big(\frac{W_{0}}{W_{\ell}}\Big)\circ\rho.$
Then, $u_{X}$ is a local defining equation of $D_{X}$ at $\xi$. If $\xi\in D_{Y}$ and $\ell\geq 1$, then $u_{Y}$ is a local defining equation of $D_{Y}$ at $\xi$. If $\xi\notin D_{Y}$, then we may take $\ell=0$ and then $u_{Y}\equiv 1$ near $\xi$.
Lemma 4.7.
With notation as above, there exist an analytic neighborhood $U$ of $\xi$ in $\overline{\Gamma}_{F}$ and constants $C_{X}\geq 1$ and $C_{Y}\geq 1$ such that for all $(x,F(x))\in U\cap\Gamma_{F}$ one has
(4.1) $\frac{1}{|u_{X}|}\leq\|x\|\leq\frac{C_{X}}{|u_{X}|}.$
Moreover, if $\xi\in D_{X}\cap D_{Y}$ and $\ell\geq 1$, then for all $(x,F(x))\in U\cap\Gamma_{F}$ one has
(4.2) $\frac{1}{|u_{Y}|}\leq\|F(x)\|\leq\frac{C_{Y}}{|u_{Y}|}.$
If $\xi\in D_{X}\setminus D_{Y}$, then $u_{Y}$ is a unit near $\xi$ and $\|F(x)\|$ is bounded on $U\cap\Gamma_{F}$. In particular, for $\xi\in D_{X}\cap D_{Y}$ and $\theta\geq 0$, the inequality
$\|x\|^{\theta}\leq C\|F(x)\|$
holds on $(U\cap\Gamma_{F})\setminus D_{X}$ if and only if
(4.3) $|u_{Y}|\leq C^{\prime}|u_{X}|^{\theta}$
holds on $(U\cap\Gamma_{F})\setminus D_{X}$.
Proof.
Work in the affine chart $\{Z_{i}\neq 0\}\subseteq\mathbb{P}^{n}_{\mathbb{C}}$ with coordinates $z_{0}:=Z_{0}/Z_{i}$ and $z_{j}:=Z_{j}/Z_{i}$ for $j\neq i$. On $\mathbb{A}^{n}_{\mathbb{C}}=\{Z_{0}\neq 0\}$ we have $x_{j}=Z_{j}/Z_{0}=z_{j}/z_{0}$. Thus, on $U\cap\Gamma_{F}$ we have $u_{X}=z_{0}\circ\pi$ and
$|x_{j}|=\frac{|z_{j}\circ\pi|}{|u_{X}|}.$
Shrinking $U$ if necessary, we may assume $U$ is relatively compact. Since each $z_{j}\circ\pi$ is holomorphic on $U$, it is bounded on $U$. Hence, there exists $M_{X}\geq 1$ such that $|z_{j}\circ\pi|\leq M_{X}$ on $U$ for all $j$. This gives $\|x\|\leq M_{X}/|u_{X}|$, which is the upper bound in (4.1). For the lower bound, note that $x_{i}=Z_{i}/Z_{0}=1/(Z_{0}/Z_{i})=1/z_{0}$, so on $U\cap\Gamma_{F}$ we have $\|x\|\geq|x_{i}|=1/|u_{X}|$.
For the target side, the argument is analogous in the affine chart $\{W_{\ell}\neq 0\}\subseteq\mathbb{P}^{m}_{\mathbb{C}}$. If $\ell\geq 1$ and $\xi\in D_{Y}$, then $u_{Y}=(W_{0}/W_{\ell})\circ\rho$ vanishes at $\xi$ and $F_{j}=W_{j}/W_{0}=(W_{j}/W_{\ell})/(W_{0}/W_{\ell})$, so bounding the affine coordinates yields (4.2). If $\xi\in D_{X}\setminus D_{Y}$, we may take $\ell=0$, so $u_{Y}\equiv 1$ near $\xi$ and $\|F(x)\|$ is bounded. Finally, when $\xi\in D_{X}\cap D_{Y}$, combining (4.1) and (4.2) shows that $\|x\|^{\theta}\leq C\|F(x)\|$ holds on $(U\cap\Gamma_{F})\setminus D_{X}$ if and only if $|u_{Y}|\leq C^{\prime}|u_{X}|^{\theta}$ holds there. ∎
Local and global analytic exponents at infinity. Lemma 4.7 allows us to recast the definition of Łojasiewicz exponent at infinity in terms of divisors at infinity.
Definition 4.8.
  1. (1)
    Fix $\xi\in D_{X}$ and let $u_{X}$ and $u_{Y}$ be as above. The local analytic Łojasiewicz exponent at infinity of $F$ along $\xi$ is
    $\mathcal{L}^{\text{an}}_{\infty,\xi}(F):=\sup\left\{\theta\geq 0\ \middle|\ \exists\,C>0\text{ such that }|u_{Y}|\leq C|u_{X}|^{\theta}\text{ on }(U_{\xi}\cap\Gamma_{F})\setminus D_{X}\right\},$
    where $U_{\xi}$ is a sufficiently small analytic neighborhood of $\xi$ in $\overline{\Gamma}_{F}$. (Note that, if $\xi\in D_{X}\setminus D_{Y}$, then $u_{Y}$ is a unit near $\xi$ and therefore $\mathcal{L}^{\text{an}}_{\infty,\xi}(F)=0$.)
  2. (2)
    The global analytic Łojasiewicz exponent at infinity of $F$ is
    $\mathcal{L}^{\text{an}}_{\infty}(F):=\inf_{\xi\in D_{X}}\mathcal{L}^{\text{an}}_{\infty,\xi}(F).$
We work with the convention $\inf\varnothing=\infty$, $1/\infty=0$ and $1/0=\infty$. The next result shows that Łojasiewicz exponent at infinity can also be understood in our algebraic framework.
Theorem 4.9.
For every $\xi\in D_{X}$, consider the graded families of ideals in $\mathcal{O}_{\overline{\Gamma}_{F},\xi}$ given by
$\mathfrak{a}_{\bullet,\xi}:=\{\mathscr{I}_{X,\xi}^{p}\}_{p\geq 1}\text{ and }\mathfrak{b}_{\bullet,\xi}:=\{\mathscr{I}_{Y,\xi}^{q}\}_{q\geq 1}.$
Then,
$\mathcal{L}^{\text{an}}_{\infty,\xi}(F)=\frac{1}{\mathcal{L}_{\mathfrak{b}_{\bullet,\xi}}(\mathfrak{a}_{\bullet,\xi})}=\frac{1}{\mathcal{L}_{\mathscr{I}_{Y,\xi}}(\mathscr{I}_{X,\xi})}.$
Equivalently,
$\mathcal{L}^{\text{an}}_{\infty,\xi}(F)=\frac{1}{\inf\left\{\frac{q}{p}\in\mathbb{Q}_{>0}\ \middle|\ \mathscr{I}_{Y,\xi}^{\,qt}\subseteq\overline{\mathscr{I}_{X,\xi}^{\,pt}}\text{ for all }t\gg 1\right\}}.$
Proof.
If $\xi\in D_{X}\setminus D_{Y}$, then $\mathscr{I}_{Y,\xi}=\mathcal{O}_{\overline{\Gamma}_{F},\xi}$, so the admissible set in Definition 4.3 for $\mathcal{L}_{\mathfrak{b}_{\bullet,\xi}}(\mathfrak{a}_{\bullet,\xi})$ is empty and $\mathcal{L}_{\mathfrak{b}_{\bullet,\xi}}(\mathfrak{a}_{\bullet,\xi})=\infty$, i.e., $1/\mathcal{L}_{\mathfrak{b}_{\bullet,\xi}}(\mathfrak{a}_{\bullet,\xi})=0$. On the other hand, $u_{Y}$ is a unit near $\xi$, so $\mathcal{L}^{\text{an}}_{\infty,\xi}(F)=0$ by Definition 4.8. Thus, the equality holds in this case.
Assume now $\xi\in D_{X}\cap D_{Y}$. Let $\theta=p/q\in\mathbb{Q}_{>0}$. By Lemma 4.7, the condition $\theta\leq\mathcal{L}^{\text{an}}_{\infty,\xi}(F)$ is equivalent to the existence of $C>0$ such that
$|u_{Y}|\leq C|u_{X}|^{p/q}\text{ on }(U_{\xi}\cap\Gamma_{F})\setminus D_{X}.$
Equivalently, $|u_{Y}|^{q}\leq C^{q}|u_{X}|^{p}$ holds in a punctured neighborhood of $\xi$. By the Lejeune-Jalabert–Teissier criterion, Lemma 4.4, this is equivalent to the containment
$(u_{Y})^{q}\subseteq\overline{(u_{X})^{p}}\text{ in }\mathcal{O}_{\overline{\Gamma}_{F},\xi}.$
Since $D_{X}$ and $D_{Y}$ are Cartier divisors, we have $\mathscr{I}_{X,\xi}=(u_{X})$ and $\mathscr{I}_{Y,\xi}=(u_{Y})$. Thus, the above containment is equivalent to
$\mathscr{I}_{Y,\xi}^{\,q}\subseteq\overline{\mathscr{I}_{X,\xi}^{\,p}}\text{ in }\mathcal{O}_{\overline{\Gamma}_{F},\xi}.$
Raising to $t$-th powers yields
$\mathscr{I}_{Y,\xi}^{\,qt}\subseteq\overline{\mathscr{I}_{X,\xi}^{\,pt}}\text{ for all }t\geq 1.$
Therefore,
$\frac{1}{\mathcal{L}^{\text{an}}_{\infty,\xi}(F)}=\inf\left\{\frac{q}{p}\in\mathbb{Q}_{>0}\ \middle|\ \mathscr{I}_{Y,\xi}^{\,qt}\subseteq\overline{\mathscr{I}_{X,\xi}^{\,pt}}\text{ for all }t\gg 1\right\}=\mathcal{L}_{\mathfrak{b}_{\bullet,\xi}}(\mathfrak{a}_{\bullet,\xi}).\qed$
Part III Valuations as the Computational Engine
In this part, we describe the Łojasiewicz exponent in valuative terms and identify conditions under which the lower bound supremum over valuations is attained, reduces to a finite maximum, and computes the Łojasiewicz exponent.
5. Valuative formulas and the finite–max principle
In this section we establish the valuative framework that serves as the computational engine of the paper. For a pair of filtrations $({\mathfrak{a}_{\bullet}},{\mathfrak{b}_{\bullet}})$ on a Noetherian local domain, we first express $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$ in terms of the asymptotic valuative data $v({\mathfrak{a}_{\bullet}})$ and $v({\mathfrak{b}_{\bullet}})$ and obtain the fundamental bound
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\ \geq\ \sup_{v}\frac{v({\mathfrak{a}_{\bullet}})}{v({\mathfrak{b}_{\bullet}})},$
where $v$ ranges over valuations centered at $\mathfrak{m}$ (Theorem 5.6). The main structural contribution is to isolate verifiable hypotheses under which this supremum both equals $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$ and is attained by a valuation, so the optimization problem reduces to a finite maximum: first via finite divisorial testing sets arising from Noetherian/finitely generated Rees algebras (Theorems 5.8 and 5.9), and second via compactness of normalized valuation spaces together with uniform convergence of the functions $v\mapsto v(\mathfrak{a}_{p})/p$ and $v\mapsto v(\mathfrak{b}_{q})/q$ (Theorems 5.14, 5.16 and 5.17).
5.1. Valuations and asymptotic values of filtrations
We recall basic notions on valuations and the asymptotic value of a graded family with respect to a valuation. For a geometric perspective on spaces of valuations and their role in controlling singularity invariants – especially in dimension two – we refer to Favre–Jonsson [16] monograph.
Throughout this section, $(R,\mathfrak{m})$ is a Noetherian local domain. A (semi)valuation on $R$ is a map
$v:R\setminus\{0\}\to\mathbb{R}_{\geq 0}$
satisfying $v(xy)=v(x)+v(y)$ and $v(x+y)\geq\min\{v(x),v(y)\}$, extended by $v(0)=\infty$. When discussing valuations on $R$, we always assume $v$ is centered at $\mathfrak{m}$, i.e., $v(x)>0$ for all $x\in\mathfrak{m}$. For an ideal $I\subseteq R$ we write
$v(I):=\min\{v(x):x\in I\}.$
For a graded family ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$, set
$v({\mathfrak{a}_{\bullet}}):=\lim_{p\to\infty}\frac{v(\mathfrak{a}_{p})}{p}=\inf_{p\geq 1}\frac{v(\mathfrak{a}_{p})}{p}.$
Remark 5.1 (Subadditivity and existence of $v({\mathfrak{a}_{\bullet}})$).
Let ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ be a graded family and let $v$ be a valuation centered at $\mathfrak{m}$. Then, the sequence $\{v(\mathfrak{a}_{p})\}_{p\geq 1}$ is subadditive:
$v(\mathfrak{a}_{p+p^{\prime}})\leq v(\mathfrak{a}_{p})+v(\mathfrak{a}_{p^{\prime}})\quad\forall p,p^{\prime}\geq 1.$
Thus, by Fekete’s lemma, the limit $v({\mathfrak{a}_{\bullet}})=\lim_{p\to\infty}v(\mathfrak{a}_{p})/p$ exists and equals $\inf_{p\geq 1}v(\mathfrak{a}_{p})/p$.
5.2. Valuative criterion for integral closure
We recall classical valuative criteria for integral closure, which we will repeatedly used in the paper.
Lemma 5.2 (Classical valuative criterion for integral closure).
Let $(R,\mathfrak{m})$ be a Noetherian local domain, $I\subseteq R$ an ideal, and $x\in R$. Then, $x\in\overline{I}$ if and only if
$v(x)\geq v(I)$
for every valuation $v$ of $\operatorname{Frac}(R)$ centered at $\mathfrak{m}$. Moreover, it suffices to test the inequalities on the finitely many Rees valuations of $I$, equivalently on the divisorial valuations associated to prime divisors on the normalized blowup of $I$.
Proof.
This result is classical and goes back to Rees [45] and Lipman [38]; see also Huneke–Swanson [30, Chapters 6 and 10] for a modern account. ∎
Remark 5.3 (Analytic arc criterion).
When $R=\mathcal{O}_{\mathbb{C}^{d},0}$, the same condition can be tested on analytic arcs via the curve criterion for integral closure; see, for instance [36].
On a fixed birational model, the classical valuative criterion gives rise to the following reduction statement.
Lemma 5.4 (Integral–closure reduction on a fixed model).
Let $(R,\mathfrak{m})$ be an excellent Noetherian local domain. Let $\pi:Y\to\operatorname{Spec}(R)$ be a proper birational morphism such that $Y$ is normal. Let $\mathfrak{a},\mathfrak{b}\subseteq R$ be ideals, and assume that $\mathfrak{a}\mathcal{O}_{Y}$ and $\mathfrak{b}\mathcal{O}_{Y}$ are invertible. Write
$\mathfrak{a}\mathcal{O}_{Y}=\mathcal{O}_{Y}(-A),\quad\mathfrak{b}\mathcal{O}_{Y}=\mathcal{O}_{Y}(-B),$
with $A,B$ effective Cartier divisors supported on the exceptional locus.
Then, for all $p,q\geq 1$, the following are equivalent:
  1. (1)
    $\mathfrak{b}^{q}\subseteq\overline{\mathfrak{a}^{p}}$ in $R$;
  2. (2)
    $\pi^{*}(\mathfrak{b}^{q})\subseteq\overline{\pi^{*}(\mathfrak{a}^{p})}$ in $\mathcal{O}_{Y}$;
  3. (3)
    $q\,B\geq p\,A$ as divisors on $Y$;
  4. (4)
    $\operatorname{\mathrm{ord}}_{E}(\mathfrak{b}^{q})\geq\operatorname{\mathrm{ord}}_{E}(\mathfrak{a}^{p})$ for every prime divisor $E$ on $Y$.
In particular, integral–closure containments are completely determined by the finitely many divisorial valuations $\{\operatorname{\mathrm{ord}}_{E}\}$ arising from the exceptional divisors of $\pi$.
Proof.
Observe that $\mathfrak{a}^{p}\mathcal{O}_{Y}=\mathcal{O}_{Y}(-pA)$ and $\mathfrak{b}^{q}\mathcal{O}_{Y}=\mathcal{O}_{Y}(-qB)$ are invertible, and so, integrally closed on $Y$. Furthermore, since $Y$ is normal, $\pi$ factors through the normalized blowup of $\mathfrak{a}$. Thus, we have the following standard equality (cf. [30, Chapters 10 and 19]):
$\overline{\mathfrak{a}^{p}}=\pi_{*}(\mathfrak{a}^{p}\mathcal{O}_{Y}).$
Particularly, if $\mathfrak{b}^{q}\mathcal{O}_{Y}\subseteq\mathfrak{a}^{p}\mathcal{O}_{Y}$, then pushing forward gives $\mathfrak{b}^{q}\subseteq\pi_{*}(\mathfrak{a}^{p}\mathcal{O}_{Y})=\overline{\mathfrak{a}^{p}}.$ On the other hand, if $\mathfrak{b}^{q}\subseteq\overline{\mathfrak{a}^{p}}=\pi_{*}(\mathfrak{a}^{p}\mathcal{O}_{Y})$ then $\mathfrak{b}^{q}\mathcal{O}_{Y}\subseteq(\pi_{*}(\mathfrak{a}^{p}\mathcal{O}_{Y}))\mathcal{O}_{Y}\subseteq\mathfrak{a}^{p}\mathcal{O}_{Y}$, where the latter containment follows from the natural evaluation map $\pi^{*}\pi_{*}(\mathfrak{a}^{p}\mathcal{O}_{Y})\rightarrow\mathfrak{a}^{p}\mathcal{O}_{Y}$. Hence, we have the equivalence (1) $\Longleftrightarrow$ (2).
The equivalences with (3) and (4) follow from the fact that $\mathfrak{a}^{p}\mathcal{O}_{Y}$ and $\mathfrak{b}^{q}\mathcal{O}_{Y}$ are invertible sheaves, and so $\mathfrak{b}^{q}\mathcal{O}_{Y}\subseteq\mathfrak{a}^{p}\mathcal{O}_{Y}\Longleftrightarrow\mathcal{O}_{Y}(-qB)\subseteq\mathcal{O}_{Y}(-pA)$, which is equivalent to $qB\geq pA\Longleftrightarrow q\operatorname{\mathrm{ord}}_{E}(\mathfrak{b})\geq p\operatorname{\mathrm{ord}}_{E}(\mathfrak{a})$ for all prime divisors $E\subseteq Y$. ∎
Remark $5.$
In the Newton nondegenerate setting, the toric modification associated to the Newton fan of $f$ provides such a morphism $\pi$, and the divisors $E$ correspond to compact facets of the Newton polyhedron (cf. [11, 33, 42]).
5.3. Valuative formula for $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$
We derive a general valuative lower bound for the algebraic Łojasiewicz exponent in terms of ratios of asymptotic valuation values. This is the basic computational engine that allows us to explore the finiteness mechanisms later.
Theorem 5.6 (Valuative lower bound).
Assume that $(R,\mathfrak{m})$ is a Noetherian local domain. Let ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ and ${\mathfrak{b}_{\bullet}}=\{\mathfrak{b}_{q}\}_{q\geq 1}$ be graded families of ideals in $R$. Then
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\ \geq\ \sup_{v}\ \frac{v({\mathfrak{a}_{\bullet}})}{v({\mathfrak{b}_{\bullet}})},$
where the supremum is taken over all valuations $v$ centered at $\mathfrak{m}$ with $v({\mathfrak{b}_{\bullet}})>0$.
Proof.
Set $M=\sup_{v}{v({\mathfrak{a}_{\bullet}})}/{v({\mathfrak{b}_{\bullet}})}.$ Fix a rational number $\frac{q}{p}\in\mathbb{Q}_{>0}$ which is admissible in the definition of $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$, i.e. such that
$\mathfrak{b}_{qt}\subseteq\overline{\mathfrak{a}_{pt}}\quad\text{for all integers }t\gg 1.$
Let $v$ be a valuation of $\operatorname{Frac}(R)$ centered at $\mathfrak{m}$ with $v({\mathfrak{b}_{\bullet}})>0$. By the valuative criterion for integral closure (Lemma 5.2), the containment $\mathfrak{b}_{qt}\subseteq\overline{\mathfrak{a}_{pt}}$ implies $v(\mathfrak{b}_{qt})\geq v(\mathfrak{a}_{pt})$ for all $t\gg 1$. Thus,
(5.1) $\frac{v(\mathfrak{b}_{qt})}{t}\ \geq\ \frac{v(\mathfrak{a}_{pt})}{t}\quad\text{for all }t\gg 1.$
By Remark 5.1, the limits $v({\mathfrak{a}_{\bullet}})=\lim_{r\to\infty}\frac{v(\mathfrak{a}_{r})}{r}$ and $v({\mathfrak{b}_{\bullet}})=\lim_{r\to\infty}\frac{v(\mathfrak{b}_{r})}{r}$ exist. In particular, we have
[…]
This, together with (5.1), implies that $q\,v({\mathfrak{b}_{\bullet}})\ \geq\ p\,v({\mathfrak{a}_{\bullet}}).$ Since $v({\mathfrak{b}_{\bullet}})>0$, we conclude that
$\frac{q}{p}\ \geq\ \frac{v({\mathfrak{a}_{\bullet}})}{v({\mathfrak{b}_{\bullet}})}.$
As this holds for every valuation $v$ centered at $\mathfrak{m}$, it follows that $\frac{q}{p}\geq M.$ Taking the infimum over all admissible $\frac{q}{p}$ in the definition of $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$ yields the desired inequality $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\geq M$. ∎
Example 5.7.
Let $R=\mathbb{C}[[x,y]],\mathfrak{m}=(x,y)$, and consider
$\mathfrak{a}=(x^{\alpha},y^{\beta})\text{ and }\mathfrak{b}=\mathfrak{m}.$
For a monomial valuation $v_{u}$, corresponding to $u=(u_{1},u_{2})\in\mathbb{Z}^{2}_{\geq 0}$, with $v_{u}(x)=u_{1}$ and $v_{u}(y)=u_{2}$, we get
$\frac{v_{u}(\mathfrak{a})}{v_{u}(\mathfrak{b})}=\frac{\min\{\alpha\cdot u_{1},\beta\cdot u_{2}\}}{\min\{u_{1},u_{2}\}}.$
Taking the maximum value over $u$ gives $\max\{\alpha,\beta\}$. Thus, Theorem 5.6 gives
$\mathcal{L}_{\mathfrak{b}}(\mathfrak{a})\geq\max\{\alpha,\beta\}.$
On the other hand, by Example 4.6, we know that $\mathcal{L}_{\mathfrak{b}}(\mathfrak{a})=\max\{\alpha,\beta\}.$ This example illustrates that the bound given in Theorem 5.6 is sharp.
Theorem 5.6 shows that $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$ is governed by a valuative optimization problem. In general, however, the inquality may be strict, the resulting supremum need not be a maximum, and the relevant candidate valuations need not be controlled uniformly without additional finiteness hypotheses. The next subsections isolate two mechanisms that force the supremum to be attained and to reduce to a finite maximum.
5.4. From supremum to maximum: finite testing and Noetherian Rees algebra
We identify the first mechanism that forces the valuative supremum to define the Łojasiewicz exponent and to be a finite maximum.
Theorem 5.8 (Finite testing set $\Rightarrow$ finite maximum formula).
Let $(R,\mathfrak{m})$ be a local domain and let ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ and ${\mathfrak{b}_{\bullet}}=\{\mathfrak{b}_{q}\}_{q\geq 1}$ be graded families of ideals in $R$. Fix a finite set of (divisorial) valuations $V=\{v_{1},\dots,v_{N}\}$ of $\operatorname{Frac}(R)$ centered at $\mathfrak{m}$. Assume:
  1. (H1)
    For all $p,q\geq 1$ one has the equivalence
    (5.2) $\mathfrak{b}_{q}\subseteq\overline{\mathfrak{a}_{p}}\quad\Longleftrightarrow\quad v_{i}(\mathfrak{b}_{q})\geq v_{i}(\mathfrak{a}_{p})\ \text{ for all }1\leq i\leq N.$
  2. (H2)
    For each $i$ we have $v_{i}({\mathfrak{b}_{\bullet}})>0$.
Then,
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})=\max_{1\leq i\leq N}\frac{v_{i}({\mathfrak{a}_{\bullet}})}{v_{i}({\mathfrak{b}_{\bullet}})}.$
Proof.
Set
$M\ :=\ \max_{1\leq i\leq N}\ \frac{v_{i}({\mathfrak{a}_{\bullet}})}{v_{i}({\mathfrak{b}_{\bullet}})}\ \in\ (0,\infty),$
which is finite by (H2). We prove the two inequalities separately.
Step 1: Lower bound $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\geq M$. Fix $\frac{q}{p}\in\mathbb{Q}_{>0}$ admissible in the definition of $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$, so that
$\mathfrak{b}_{qt}\ \subseteq\ \overline{\mathfrak{a}_{pt}}\quad\text{for all }t\gg 1.$
By the same arguments as in the proof of Theorem 5.6, we get
$\frac{q}{p}\ \geq\ \frac{v_{i}({\mathfrak{a}_{\bullet}})}{v_{i}({\mathfrak{b}_{\bullet}})}\quad\text{for all }1\leq i\leq N.$
Thus, $\frac{q}{p}\geq M$. Taking the infimum over all admissible $\frac{q}{p}$ gives $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\geq M$.
Step 2: Upper bound $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\leq M$. Fix $\varepsilon>0$ and set $\beta\ :=\ \min_{1\leq i\leq N}v_{i}({\mathfrak{b}_{\bullet}})\ >\ 0.$ Choose $p\gg 1$ and set
[…]
For each $i$, by the definition of $v_{i}({\mathfrak{a}_{\bullet}})$ as a limit along the subsequence $\{pt\}$, there exists $t_{0}(i)\in\mathbb{N}$ such that for all integers $t\geq t_{0}(i)$,
$\frac{v_{i}(\mathfrak{a}_{pt})}{pt}\ \leq\ v_{i}({\mathfrak{a}_{\bullet}})+\varepsilon.$
Let $t_{0}:=\max_{i}t_{0}(i)$, so the above estimate holds for all $i$ and all $t\geq t_{0}$.
On the other hand, since $v_{i}(\mathfrak{b}_{qt})\ \geq\ qt\,v_{i}({\mathfrak{b}_{\bullet}})$, for every $t\geq t_{0}$ and every $i$, we have
$v_{i}(\mathfrak{b}_{qt})\ \geq\ qt\,v_{i}({\mathfrak{b}_{\bullet}})\ \geq\ \left(M+\frac{\varepsilon}{\beta}\right)pt\,v_{i}({\mathfrak{b}_{\bullet}})\ \geq\ pt\,(M\,v_{i}({\mathfrak{b}_{\bullet}})+\varepsilon)\ \geq\ pt\,(v_{i}({\mathfrak{a}_{\bullet}})+\varepsilon)\ \geq\ v_{i}(\mathfrak{a}_{pt}).$
That is, for all integers $t\geq t_{0}$, $v_{i}(\mathfrak{b}_{qt})\geq v_{i}(\mathfrak{a}_{pt})$ for every $i$. By (H1) applied to $(pt,qt)$, it follows that
$\mathfrak{b}_{qt}\ \subseteq\ \overline{\mathfrak{a}_{pt}}\quad\text{for all }t\geq t_{0}.$
Thus, $\frac{q}{p}$ is admissible in the definition of $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$, and so
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\ \leq\ \frac{q}{p}\ \leq\ M+\frac{\varepsilon}{\beta}+\frac{1}{p}.$
Letting $p\to\infty$ gives $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\leq M+\frac{\varepsilon}{\beta}$, and then letting $\varepsilon\to 0$ yields $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\leq M$. This completes the proof. ∎
In most applications below, the equivalence in (H1) arises from integral-closure testing on a fixed model (e.g., a normalized blowup or toric modification). While the forward implication alone suffices for the lower bound, the equivalence ensures that admissible containments can be produced uniformly.
Hypotheses for the finite testing principle (Theorem 5.8) are satisfied when the Rees algebra is Noetherian, as shown in the next result.
Theorem 5.9 (Finite Rees valuations for a Noetherian filtration).
Let $(R,\mathfrak{m})$ be an excellent Noetherian local domain, and let ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ be a filtration of ideals in $R$. Assume that the Rees algebra $\mathcal{R}({\mathfrak{a}_{\bullet}}):=\bigoplus_{p\geq 0}\mathfrak{a}_{p}T^{p}$ is finitely generated as an RR–algebra. Let $\overline{\mathcal{R}({\mathfrak{a}_{\bullet}})}$ denote its integral closure, and set $\overline{Y}:=\operatorname{Proj}\overline{\mathcal{R}({\mathfrak{a}_{\bullet}})}.$ Then, there exists a finite set of divisorial valuations
$\mathcal{V}({\mathfrak{a}_{\bullet}})=\{v_{1},\dots,v_{r}\},$
arising from prime divisors on $\overline{Y}$ dominating $\operatorname{Spec}R$, such that for every filtration ${\mathfrak{b}_{\bullet}}=\{\mathfrak{b}_{q}\}_{q\geq 1}$ of ideals in $R$ satisfying $v_{i}({\mathfrak{b}_{\bullet}})>0$ for all $1\leq i\leq r$, one has
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})=\max_{1\leq i\leq r}\frac{v_{i}({\mathfrak{a}_{\bullet}})}{v_{i}({\mathfrak{b}_{\bullet}})}.$
Proof.
We break the proof into two steps.
Step 1: a finite divisorial testing set for the integral closures $\overline{\mathfrak{a}_{p}}$. Let
$\mathcal{R}:=\mathcal{R}({\mathfrak{a}_{\bullet}})=\bigoplus_{p\geq 0}\mathfrak{a}_{p}T^{p}\subseteq R[T]\quad\text{and}\quad\overline{\mathcal{R}}:=\overline{\mathcal{R}({\mathfrak{a}_{\bullet}})}$
be the integral closure of $\mathcal{R}$ in its total ring of fractions. Since $\mathcal{R}$ is finitely generated over the excellent Noetherian domain $R$, it is Noetherian, and $\overline{\mathcal{R}}$ is a finite $\mathcal{R}$–module; in particular $\overline{\mathcal{R}}$ is again a finitely generated RR–algebra.
Let $\pi:\overline{Y}\to\operatorname{Spec}R$ the induced proper birational morphism. Write $E\subseteq\overline{Y}$ for the exceptional locus of $\pi$. Since $\overline{Y}$ is of finite type over $R$, the exceptional locus has only finitely many irreducible components of codimension one; denote them by
$E_{1},\dots,E_{r}.$
Each $E_{i}$ defines a divisorial valuation $v_{i}$ of $\operatorname{Frac}(R)$ by
$v_{i}(h):=\operatorname{\mathrm{ord}}_{E_{i}}(\pi^{*}h)\quad(0\neq h\in\operatorname{Frac}(R)),$
normalized so that $v_{i}(R)\geq 0$ and $v_{i}(\mathfrak{m})>0$ whenever $\pi(E_{i})\subseteq V(\mathfrak{m})$.
We claim that these finitely many valuations test the integral closure of each $\mathfrak{a}_{p}$ in the sense that
(5.3) $\overline{\mathfrak{a}_{p}}=\{\,h\in R:\ v_{i}(h)\geq v_{i}(\mathfrak{a}_{p})\ \text{for all }i=1,\dots,r\,\}.$
Indeed, for $h\in R$ we have
$h\in\overline{\mathfrak{a}_{p}}\ \Longleftrightarrow\ hT^{p}\in\overline{\mathfrak{a}_{p}T^{p}}\subseteq\overline{\mathcal{R}}_{p}\ \Longleftrightarrow\ hT^{p}\in\overline{\mathcal{R}},$
where $\overline{\mathcal{R}}_{p}$ denotes the degree-$p$ piece of $\overline{\mathcal{R}}$. Now, $\overline{\mathcal{R}}$ is a normal Noetherian domain, so a Krull domain; therefore,
$\overline{\mathcal{R}}=\bigcap_{\begin{subarray}{c}\mathfrak{P}\in\operatorname{Spec}\overline{\mathcal{R}}\\ \mathrm{ht}(\mathfrak{P})=1\end{subarray}}(\overline{\mathcal{R}})_{\mathfrak{P}}\ \subseteq\ \operatorname{Frac}(\overline{\mathcal{R}}).$
Consequently, $hT^{p}\in\overline{\mathcal{R}}$ if and only if $hT^{p}$ lies in each DVR $(\overline{\mathcal{R}})_{\mathfrak{P}}$ for $\mathrm{ht}(\mathfrak{P})=1$. Among the height-one primes $\mathfrak{P}$, only those corresponding to the codimension-one components of the exceptional locus contribute nontrivially to membership in $\mathfrak{a}_{p}$; the corresponding valuations are precisely $v_{1},\dots,v_{r}$, and the resulting condition is exactly (5.3). This proves the claim.
We set
$\mathcal{V}({\mathfrak{a}_{\bullet}}):=\{v_{1},\dots,v_{r}\}.$
Step 2: computation of $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$. Fix any filtration ${\mathfrak{b}_{\bullet}}=\{\mathfrak{b}_{q}\}_{q\geq 1}$ of ideals in $R$. Set
$\lambda\ :=\ \max_{1\leq i\leq r}\ \frac{v_{i}({\mathfrak{a}_{\bullet}})}{v_{i}({\mathfrak{b}_{\bullet}})}\ \in\ \mathbb{R}_{\geq 0}.$
Let $\frac{q}{p}\in\mathbb{Q}_{>0}$ be admissible in Definition 4.3, i.e.,
$\mathfrak{b}_{qt}\subseteq\overline{\mathfrak{a}_{pt}}\quad\text{for all integers }t\gg 1.$
By (5.3) applied to $pt$, this containment is equivalent to
$v_{i}(\mathfrak{b}_{qt})\ \geq\ v_{i}(\mathfrak{a}_{pt})\quad\text{for all }i=1,\dots,r\ \text{and all }t\gg 1.$
By the same arguments as in the proof of Theorem 5.6, we get $q\,v_{i}({\mathfrak{b}_{\bullet}})\geq p\,v_{i}({\mathfrak{a}_{\bullet}})$, i.e.,
$\frac{q}{p}\ \geq\ \frac{v_{i}({\mathfrak{a}_{\bullet}})}{v_{i}({\mathfrak{b}_{\bullet}})}\quad\text{for all }i.$
Hence, $\frac{q}{p}\geq\lambda$. Taking the infimum over admissible $\frac{q}{p}$ yields $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\geq\lambda$.
Conversely, fix $\varepsilon>0$ sufficiently small. By the definition of $v_{i}({\mathfrak{a}_{\bullet}})$ and $v_{i}({\mathfrak{b}_{\bullet}})$, there exists $p_{0}$ such that for all $p\geq p_{0}$ and all $i$,
$\frac{v_{i}(\mathfrak{a}_{p})}{p}\ \leq\ v_{i}({\mathfrak{a}_{\bullet}})+\varepsilon\text{ and }\frac{v_{i}(\mathfrak{b}_{p})}{p}\ \geq\ v_{i}({\mathfrak{b}_{\bullet}})-\varepsilon>0.$
Set
$\delta=\epsilon\cdot\max\limits_{1\leq i\leq r}\frac{v_{i}({\mathfrak{a}_{\bullet}})+v_{i}({\mathfrak{b}_{\bullet}})}{v_{i}({\mathfrak{b}_{\bullet}})(v_{i}({\mathfrak{b}_{\bullet}})-\epsilon)}.$
Then, $\delta\rightarrow 0$ as $\epsilon\rightarrow 0$ and $(\lambda+\delta)(v({\mathfrak{b}_{\bullet}})-\epsilon)\geq v_{i}({\mathfrak{a}_{\bullet}})+\epsilon.$ Choose $p\geq p_{0}$ large and set
$q:=\left\lceil(\lambda+\delta)p\right\rceil.$
Now choose $t_{0}$ large enough so that $qt\geq p_{0}$ for all $i$ and all $t\geq t_{0}$. Then, for every $i$ and every integer $t\geq t_{0}$ we have
$v_{i}(\mathfrak{b}_{qt})\ \geq\ qt\,(v_{i}({\mathfrak{b}_{\bullet}})-\varepsilon)\ \geq\ (\lambda+\delta)pt\,(v_{i}({\mathfrak{b}_{\bullet}})-\varepsilon)\ \geq\ pt\,(v_{i}({\mathfrak{a}_{\bullet}})+\varepsilon)\ \geq\ v_{i}(\mathfrak{a}_{pt}).$
Hence, by (5.3),
$\mathfrak{b}_{qt}\subseteq\overline{\mathfrak{a}_{pt}}\quad\text{for all integers }t\geq t_{0}.$
Thus, $\frac{q}{p}$ is admissible in Definition 4.3, so
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\ \leq\ \frac{q}{p}\ \leq\ \lambda+\delta+\frac{1}{p}.$
Letting $p\to\infty$ and then $\varepsilon\to 0$ ($\delta\to 0$) gives $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\leq\lambda$. The theorem is proved. ∎
Remark 5.10.
  1. (1)
    The finite set $\mathcal{V}({\mathfrak{a}_{\bullet}})$ depends only on the filtration $\mathfrak{a}_{\bullet}$} and not on $\mathfrak{b}_{\bullet}$}. When ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}^{p}\}_{p\geq 1}$ for an $\mathfrak{m}$–primary ideal $\mathfrak{a}$, the set $\mathcal{V}({\mathfrak{a}_{\bullet}})$ coincides with the classical set of Rees valuations of $\mathfrak{a}$.
  2. (2)
    The condition that $v_{i}({\mathfrak{b}_{\bullet}})>0$ in Theorem 5.9 is often satisfied in practice. For instance, if $\mathfrak{b}_{\bullet}$} is linearly bounded from below, i.e., there exists a constant $c$ such that $\mathfrak{m}^{\lceil cq\rceil}\subseteq\mathfrak{b}_{q}.$ Particularly, if ${\mathfrak{b}_{\bullet}}=\{\mathfrak{m}^{q}\}_{q\in\mathbb{N}}$, then the hypothesis $v_{i}({\mathfrak{b}_{\bullet}})>0$ in Theorem 5.9 is automatically satisfied.
  3. (3)
    The conclusion of Theorem 5.9 still holds under a slightly weaker condition that $v_{i}({\mathfrak{b}_{\bullet}})=0$ implies $v_{i}({\mathfrak{a}_{\bullet}})>0$ for any $i=1,\dots,r$. This is satisfied particularly if $v_{i}({\mathfrak{a}_{\bullet}})>0$ for all $i=1,\dots,r$.
Corollary 5.11 (Finite testing set from finite generation).
Let $(R,\mathfrak{m})$ be an excellent Noetherian local domain, and let ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ be a filtration whose Rees algebra $\mathcal{R}({\mathfrak{a}_{\bullet}})$ is finitely generated. Then, there exists a finite set of divisorial valuations $\mathcal{V}$ with the following property:
For any filtration ${\mathfrak{b}_{\bullet}}=\{\mathfrak{b}_{q}\}_{q\geq 1}$, with $v({\mathfrak{b}_{\bullet}})>0\ \forall\ v\in\mathcal{V}$, and any $p,q\geq 1$,
$\mathfrak{b}_{q}\subseteq\overline{\mathfrak{a}_{p}}\quad\Longleftrightarrow\quad v(\mathfrak{b}_{q})\geq v(\mathfrak{a}_{p})\ \text{for all }v\in\mathcal{V}.$
In particular, $\mathcal{V}$ is a finite testing set for asymptotic containment problems involving $\mathfrak{a}_{\bullet}$}.
Proof.
By Theorem 5.9, there exists a finite set of divisorial valuations $\mathcal{V}({\mathfrak{a}_{\bullet}})$ computing $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$ for all filtrations $\mathfrak{b}_{\bullet}$}. The valuative criterion for integral closure then implies that the same set of valuations detects all containments $\mathfrak{b}_{q}\subseteq\overline{\mathfrak{a}_{p}}$, proving the claim. ∎
Example 5.12.
Let $R:=k[x,y]_{(x,y)},\ \mathfrak{m}:=(x,y)$, and let
$\mathfrak{a}_{1}:=(x^{4},\,x^{2}y^{3},\,y^{5})\subseteq R.$
Consider ${\mathfrak{a}_{\bullet}}:=\{\mathfrak{a}_{p}\}_{p\geq 1}\quad\text{with }\mathfrak{a}_{p}=\mathfrak{a}_{1}^{p},\quad{\mathfrak{b}_{\bullet}}:=\{\mathfrak{b}_{q}\}_{q\geq 1}\quad\text{with }\mathfrak{b}_{q}=\mathfrak{m}^{q}.$
Since the Rees algebra $R({\mathfrak{a}_{\bullet}})$ is finitely generated, Corollary 5.11 applies. Hence, there exists a finite set of divisorial valuations $V({\mathfrak{a}_{\bullet}})$ such that, for all $p,q\geq 1$,
$\mathfrak{b}_{q}\subseteq\overline{\mathfrak{a}_{p}}\quad\Longleftrightarrow\quad v(\mathfrak{b}_{q})\geq v(\mathfrak{a}_{p})\ \text{for all }v\in V({\mathfrak{a}_{\bullet}}),$
and
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})=\max_{v\in V({\mathfrak{a}_{\bullet}})}\frac{v({\mathfrak{a}_{\bullet}})}{v({\mathfrak{b}_{\bullet}})}.$
In this two–dimensional monomial setting, the valuations in $V({\mathfrak{a}_{\bullet}})$ admit an explicit description: they are precisely the divisorial (toric) valuations corresponding to the compact edges of the Newton polygon of $\mathfrak{a}_{1}$. Concretely, one may take the monomial valuations $v_{(1,1)}$ and $v_{(2,1)}$, defined by
$v_{(u,v)}(x)=u,\quad v_{(u,v)}(y)=v.$
A direct computation gives $v_{(1,1)}({\mathfrak{a}_{\bullet}})=4,v_{(2,1)}({\mathfrak{a}_{\bullet}})=5$, and $v_{(1,1)}({\mathfrak{b}_{\bullet}})=v_{(2,1)}({\mathfrak{b}_{\bullet}})=1.$ Consequently,
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})=5.$
This illustrates concretely how the valuative supremum reduces to a finite maximum on a fixed normalized blowup.
5.5. From supremum to maximum: normalized valuation space and linearly boundedness
We explain how compactness of normalized valuation spaces yields attainment of the valuative supremum under linear boundedness hypotheses.
Let $\operatorname{Val}_{\mathfrak{m}}$ denote the set of real (semi)valuations $v$ of $R$ centered at $\mathfrak{m}$. Fix a nonzero $\mathfrak{m}$-primary ideal $\mathfrak{q}\subseteq R$ and consider the normalized slice
$\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q}):=\{v\in\operatorname{Val}_{\mathfrak{m}}:v(\mathfrak{q})=1\}.$
Lemma 5.13.
Let $(R,\mathfrak{m})$ be a Noetherian local ring and let $\mathfrak{q}\subseteq R$ be an $\mathfrak{m}$–primary ideal. Then, $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})$ is compact in the topology of pointwise convergence on $R$. Moreover, for any graded family $\mathfrak{a}_{\bullet}$} of ideals in $R$, the function
$v\longmapsto v({\mathfrak{a}_{\bullet}})$
is upper semicontinuous on $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})$.
Proof.
By the discussion following [9, Definition 1.3] (see also [31, Proposition 5.9] and the discussion that followed), the normalized space $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$ is compact in the topology of pointwise convergence on $R$.
Note that, by [31, Lemma 4.1], $v\mapsto v(\mathfrak{m})$ and $v\mapsto v(\mathfrak{q})$ are continuous on $\operatorname{Val}_{\mathfrak{m}}$. Define
$\Phi:$ $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})\longrightarrow\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m}),\quad\Phi(v):=\frac{1}{v(\mathfrak{m})}\,v,\text{ and }$
$\Psi:$ $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})\longrightarrow\operatorname{Val}_{\mathfrak{m}}(q),\quad\Psi(w):=\frac{1}{w(\mathfrak{q})}\,w.$
These maps are well-defined since $\mathfrak{q}$ is $\mathfrak{m}$-primary, so $v(\mathfrak{m})>0$ and $v(\mathfrak{q})>0$ for all $v\in\operatorname{Val}_{\mathfrak{m}}$. Moreover, for each $g\in R$ one has
$(\Phi(v))(g)=\frac{v(g)}{v(\mathfrak{m})},\quad(\Psi(w))(g)=\frac{w(g)}{w(\mathfrak{q})},$
so $\Phi$ and $\Psi$ are continuous. It is immediate that $\Phi$ and $\Psi$ are inverse to each other. Hence, $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})$ is homeomorphic to the compact space $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$ and, therefore, $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})$ is compact.
Finally, by definition of the topology on $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})$, the map $v\mapsto v(g)$ is continuous for every $g\in R$. Thus, for a finitely generated ideal $I\subseteq R$, the map $v\mapsto v(I)$ is continuous. In particular,
$v(\mathfrak{a}_{\bullet})=\inf_{p\geq 1}\frac{v(\mathfrak{a}_{p})}{p}$
is an infimum of continuous functions. It follows that $v\mapsto v(\mathfrak{a}_{\bullet})$ is upper semicontinuous on $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})$. ∎
Theorem 5.14 (Attainment and valuative equality for linearly bounded filtrations).
Let $R$ be a Noetherian local domain. Let ${\mathfrak{a}_{\bullet}},{\mathfrak{b}_{\bullet}}$ be filtrations such that there exists an $\mathfrak{m}$-primary ideal $\mathfrak{q}$ and constants $c,C,d,D>0$ with
$\mathfrak{q}^{\lceil cp\rceil}\subseteq\mathfrak{a}_{p}\subseteq\mathfrak{q}^{\lfloor Cp\rfloor}\quad\text{and}\quad\mathfrak{q}^{\lceil dp\rceil}\subseteq\mathfrak{b}_{p}\subseteq\mathfrak{q}^{\lfloor Dp\rfloor}\quad\text{for all }p\geq 1.$
Then, the supremum in the valuative formula
$\sup_{v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})}\frac{v({\mathfrak{a}_{\bullet}})}{v({\mathfrak{b}_{\bullet}})}$
is finite and is attained by some valuation $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})$.
Suppose, in addition, that the functions
$v\longmapsto\frac{v(\mathfrak{a}_{p})}{p}\quad\text{and}\quad v\longmapsto\frac{v(\mathfrak{b}_{p})}{p}$
converge uniformly on $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})$ to $v({\mathfrak{a}_{\bullet}})$ and $v({\mathfrak{b}_{\bullet}})$, respectively. Then,
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\ =\ \sup_{v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})}\frac{v({\mathfrak{a}_{\bullet}})}{v({\mathfrak{b}_{\bullet}})},$
and the right-hand side is attained by some valuation in $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})$.
Proof.
By the linear boundedness assumptions,
$v({\mathfrak{a}_{\bullet}})\leq C\,v(\mathfrak{q}),\quad v({\mathfrak{b}_{\bullet}})\geq d\,v(\mathfrak{q})$
for all $v\in\operatorname{Val}_{\mathfrak{m}}$. In particular, on the compact slice $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})$ one has $v(\mathfrak{q})=1$ and, therefore, $v({\mathfrak{a}_{\bullet}})\leq C$ and $v({\mathfrak{b}_{\bullet}})\geq d>0$. Set
$s:=\sup_{v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})}\frac{v({\mathfrak{a}_{\bullet}})}{v({\mathfrak{b}_{\bullet}})}<\infty.$
By definition, there exists a sequence $\{v_{j}\}_{j\in\mathbb{N}}\subseteq\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})$ such that $s=\lim\limits_{j\rightarrow\infty}\frac{v_{j}({\mathfrak{a}_{\bullet}})}{v_{j}({\mathfrak{b}_{\bullet}})}.$ Since $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})$ is compact, by Lemma 5.13, a subsequence $v_{j_{k}}$ of this sequence converges to an element $v_{*}\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})$. The upper semicontinuity in Lemma 5.13 then gives
$v_{*}({\mathfrak{a}_{\bullet}})\geq\limsup_{k\rightarrow\infty}v_{j_{k}}({\mathfrak{a}_{\bullet}})\text{ and }v_{*}({\mathfrak{b}_{\bullet}})\geq\limsup_{k\rightarrow\infty}v_{j_{k}}({\mathfrak{b}_{\bullet}})\geq d.$
For any $\eta>0$, set
$E_{\eta}=\{v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q}):v({\mathfrak{b}_{\bullet}})\geq v_{*}({\mathfrak{b}_{\bullet}})-\eta\}.$
It follows from Lemma 5.13, that $v\mapsto v({\mathfrak{b}_{\bullet}})$ is an upper semicontinuous function. This implies that $E_{\eta}$ is a closed subset or, equivalently, its complement in $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})$ is an open subset. We then conclude, since $v_{j_{k}}\longrightarrow v_{*}$, that $v_{j_{k}}\in E_{\eta}$ for $k\gg 1$. Therefore, for $k\gg 1$,
$s=\lim_{k\rightarrow\infty}\frac{v_{j_{k}}({\mathfrak{a}_{\bullet}})}{v_{j_{k}}({\mathfrak{b}_{\bullet}})}\leq\frac{\limsup_{k\rightarrow\infty}v_{j_{k}}({\mathfrak{a}_{\bullet}})}{v_{*}({\mathfrak{b}_{\bullet}})-\eta}\leq\frac{v_{*}({\mathfrak{a}_{\bullet}})}{v_{*}({\mathfrak{b}_{\bullet}})-\eta}.$
This inequality holds for any $\eta>0$. Hence, by taking $\eta\rightarrow 0^{+}$, we have
$s\leq\frac{v_{*}({\mathfrak{a}_{\bullet}})}{v_{*}({\mathfrak{b}_{\bullet}})}\leq s.$
This forces the desired equality $s=\frac{v_{*}({\mathfrak{a}_{\bullet}})}{v_{*}({\mathfrak{b}_{\bullet}})}.$
Assume now that $\frac{v(\mathfrak{a}_{p})}{p}\to v({\mathfrak{a}_{\bullet}})$ and $\frac{v(\mathfrak{b}_{p})}{p}\to v({\mathfrak{b}_{\bullet}})$ uniformly on $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})$. Let $\lambda>s$. By definition of $s$, we have $v({\mathfrak{a}_{\bullet}})\leq s\,v({\mathfrak{b}_{\bullet}})$ for all $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})$. Thus, for all $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})$,
$\lambda v({\mathfrak{b}_{\bullet}})-v({\mathfrak{a}_{\bullet}})\ \geq\ (\lambda-s)\,v({\mathfrak{b}_{\bullet}})\ \geq\ (\lambda-s)d\ =:\ \delta.$
Choose $\varepsilon>0$ with $\varepsilon<\min\{d/2,\ \delta/(\lambda+1)\}$. Then, for every $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})$ one has $v({\mathfrak{b}_{\bullet}})-\varepsilon>0$ and
$v({\mathfrak{a}_{\bullet}})+\varepsilon\leq\lambda v({\mathfrak{b}_{\bullet}})-\delta+\varepsilon=\lambda(v({\mathfrak{b}_{\bullet}})-\epsilon)-\underbrace{(\delta-(\lambda+1)\epsilon)}_{=:\delta^{\prime}>0},$
so
$\frac{v({\mathfrak{a}_{\bullet}})+\varepsilon}{v({\mathfrak{b}_{\bullet}})-\varepsilon}\leq\lambda-\frac{\delta^{\prime}}{v({\mathfrak{b}_{\bullet}})-\epsilon}\leq\lambda-\frac{\delta^{\prime}}{d-\epsilon}<\lambda\quad\text{for all }v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q}).$
Set
$M:=\sup_{v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})}\frac{v({\mathfrak{a}_{\bullet}})+\varepsilon}{v({\mathfrak{b}_{\bullet}})-\varepsilon},$
so $M\leq\lambda-\frac{\delta^{\prime}}{d-\epsilon}<\lambda$.
By uniform convergence, there exists $N$ such that for all $p\geq N$ and all $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})$,
$\left|\frac{v(\mathfrak{a}_{p})}{p}-v({\mathfrak{a}_{\bullet}})\right|\leq\varepsilon\quad\text{and}\quad\left|\frac{v(\mathfrak{b}_{p})}{p}-v({\mathfrak{b}_{\bullet}})\right|\leq\varepsilon.$
Choose integers $P\geq N$ and $Q\geq N$ with $M\leq\frac{Q}{P}<\lambda.$
Claim. For all integers $t\geq 1$ one has
$\mathfrak{b}_{Qt}\subseteq\overline{\mathfrak{a}_{Pt}}.$
Proof of the claim. For $t\geq 1$, $Pt\geq N$ and $Qt\geq N$, so for all $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})$,
$\frac{v(\mathfrak{a}_{Pt})}{Pt}\leq v({\mathfrak{a}_{\bullet}})+\varepsilon,\quad\frac{v(\mathfrak{b}_{Qt})}{Qt}\geq v({\mathfrak{b}_{\bullet}})-\varepsilon.$
Therefore, for every $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})$ and every $t\geq 1$,
$\frac{v(\mathfrak{b}_{Qt})}{Pt}=\frac{Q}{P}\cdot\frac{v(\mathfrak{b}_{Qt})}{Qt}\ \geq\ \frac{Q}{P}\bigl(v({\mathfrak{b}_{\bullet}})-\varepsilon\bigr)\ \geq\ M\bigl(v({\mathfrak{b}_{\bullet}})-\varepsilon\bigr)\ \geq\ v({\mathfrak{a}_{\bullet}})+\varepsilon\ \geq\ \frac{v(\mathfrak{a}_{Pt})}{Pt}.$
Multiplying by $Pt$ gives
[…]
Now let $w$ be any divisorial valuation centered at $\mathfrak{m}$. Since $w(\mathfrak{q})>0$, the normalized valuation $\tilde{w}:=w/w(\mathfrak{q})$ lies in $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})$, so the previous inequality applied to $\tilde{w}$ yields $w(\mathfrak{b}_{Qt})\geq w(\mathfrak{a}_{Pt})$ for all $t\geq 1$. By the valuative criterion for integral closure (Lemma 5.2), this implies $\mathfrak{b}_{Qt}\subseteq\overline{\mathfrak{a}_{Pt}}$ for all $t\geq 1$, proving the claim. $\square$
By the claim, $\frac{Q}{P}$ is admissible $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$, and so
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\leq\frac{Q}{P}<\lambda.$
Since $\lambda>s$ was arbitrary, we obtain $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\leq s$.
On the other hand, Theorem 5.6 gives $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\geq s$. Thus, $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})=s$, and the right-hand side is attained by the first part of the theorem. ∎
Example 5.15.
Let $(R,\mathfrak{m})=k[[x,y]]$ and $\mathfrak{m}=(x,y)$. Define two filtrations ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ and ${\mathfrak{b}_{\bullet}}=\{\mathfrak{b}_{p}\}_{p\geq 1}$ by
$\mathfrak{a}_{p}:=\big(x^{i}y^{j}\in R\mid i+2j\geq 3p\big),\quad\mathfrak{b}_{p}:=\big(x^{i}y^{j}\in R\mid 2i+j\geq 5p\big).$
It can be seen that, for $p\geq 1$, $\mathfrak{a}_{p}$ is integrally closed, and
$\mathfrak{m}^{3p}\subseteq\mathfrak{a}_{p}\subseteq\mathfrak{m}^{p},\quad\mathfrak{m}^{5p}\subseteq\mathfrak{b}_{p}\subseteq\mathfrak{m}^{p}.$
It can also be checked that the functions $v\mapsto v(\mathfrak{a}_{p})/p$ and $v\mapsto v(\mathfrak{b}_{p})/p$ converge uniformly to $v({\mathfrak{a}_{\bullet}})$ and $v({\mathfrak{b}_{\bullet}})$, respectively, on the space of normalized real valuations of $\operatorname{Frac}(R)$ centered at $\mathfrak{m}$.
We claim that
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})=\frac{6}{5}.$
Indeed, fix $p,q$ with $q/p\geq 6/5$ and let $t\geq 1$. For any monomial $x^{i}y^{j}\in\mathfrak{b}_{qt}$ we have $2i+j\geq 5qt$. Minimizing $i+2j$ subject to $2i+j\geq 5qt$ and $i,j\geq 0$ yields
$i+2j\geq\frac{5}{2}qt\geq\frac{5}{2}\cdot\frac{6}{5}pt=3pt,$
and so $x^{i}y^{j}\in\mathfrak{a}_{pt}$. Thus, $\mathfrak{b}_{qt}\subseteq\mathfrak{a}_{pt}=\overline{\mathfrak{a}_{pt}}$ for all $t$, showing $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\leq 6/5$.
Conversely, assume $q/p<6/5$ and set $m_{t}=\lceil\frac{5}{2}qt\rceil$. Then, $x^{m_{t}}\in\mathfrak{b}_{qt}$ since $2m_{t}\geq 5qt$. However, $x^{m_{t}}\in\mathfrak{a}_{pt}$ would require $m_{t}\geq 3pt$, which fails for all $t\gg 1$ since $\frac{5}{2}q<3p$. Hence, $\mathfrak{b}_{qt}\nsubseteq\mathfrak{a}_{pt}=\overline{\mathfrak{a}_{pt}}$ for $t\gg 1$, proving $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\geq 6/5$, giving the claimed value of $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$.
Furthermore, consider the monomial valuation $v$ defined by $v(x)=1$, $v(y)=2$. Then, $v(\mathfrak{m})=1$, $v(\mathfrak{a}_{\bullet})=3$, and $v(\mathfrak{b}_{\bullet})=5/2$, so
$\frac{v(\mathfrak{a}_{\bullet})}{v(\mathfrak{b}_{\bullet})}=\frac{6}{5}=\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}}).$
Thus, the valuative supremum is attained by $v$, as predicted by Theorem 5.14.
The following theorem shows that the valuation attaining supremum value in Theorem 5.14 is not necessarily divisorial, indicating that focusing on divisorial valuations may not be sufficient to study Łojasiewicz exponents in general.
Theorem 5.16 (Non-divisorial attainment in dimension two).
Let $R=\mathbb{C}[[x,y]]$ with maximal ideal $\mathfrak{m}=(x,y)$, and set $\mathfrak{q}=\mathfrak{m}$. Then, there exist filtrations ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ and ${\mathfrak{b}_{\bullet}}=\{\mathfrak{b}_{p}\}_{p\geq 1}$ satisfying the linear boundedness hypotheses of Theorem 5.14 such that
$\sup_{v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})}\frac{v({\mathfrak{a}_{\bullet}})}{v({\mathfrak{b}_{\bullet}})}$
is attained by a valuation $v_{*}\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})$ which is not divisorial, and moreover no divisorial valuation in $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})$ attains this supremum.
Proof.
We work on the normalized slice $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})=\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})$, so $v(\mathfrak{m})=1$ for all $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$. We shall also use notations and terminology from Favre–Jonsson [16] monograph.
Step 1: a non-divisorial end and divisorial approximants. By Favre–Jonsson’s classification of the valuative tree of $R=\mathbb{C}[[x,y]]$ in [16], there exist non-divisorial ends (for instance curve valuations or infinitely singular valuations). Fix such an end valuation $\nu\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$. Then, $\nu$ is not divisorial, and there exists an increasing sequence of divisorial valuations
$\nu_{1}\prec\nu_{2}\prec\cdots\prec\nu_{i}\prec\cdots\prec\nu$
converging to $\nu$ in $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$, where $\prec$ denotes the tree order. We use the fact that the skewness function $\alpha(\cdot)$ on the valuative tree is strictly increasing along any nontrivial segment (see [16]).
Step 2: realizing $\alpha(v\wedge\nu_{i})$ as $v(J_{i})$. For valuations $v,\mu\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$, Favre–Jonsson [16] define the intersection product by
$v\cdot\mu:=\alpha(v\wedge\mu),$
and for fixed $\mu$ the associated extremal tree potential is $g_{\mu}(v):=v\cdot\mu=\alpha(v\wedge\mu)$. If $\mu$ is divisorial, then it was proved (see the proof of [17, Prop. 5.15]; although this was written for the ring $\mathbb{C}\{x,y\}$ of convergent formal power series, the argument applies verbatim to the formal power series ring $\mathbb{C}[[x,y]]$) that there exist irreducible elements $\phi_{\mu,1},\phi_{\mu,2}\in\mathfrak{m}$ such that
$g_{\mu}(v)=\min\{v(\phi_{\mu,1}),v(\phi_{\mu,2})\}\quad\text{for all }v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m}).$
Applying this to $\mu=\nu_{i}$, we may choose irreducibles $\phi_{i,1},\phi_{i,2}\in\mathfrak{m}$ with
$\alpha(v\wedge\nu_{i})=\min\{v(\phi_{i,1}),v(\phi_{i,2})\}\quad\forall\,v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m}).$
Set
$J_{i}:=(\phi_{i,1},\phi_{i,2})\subseteq R.$
Since $\phi_{i,1},\phi_{i,2}$ are distinct irreducible germs through the origin, the ideal $J_{i}$ has height $2$ in the regular local ring $R=\mathbb{C}[[x,y]]$ and is therefore $\mathfrak{m}$–primary. For all $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$ we have the exact identity
(5.4) $v(J_{i})=\alpha(v\wedge\nu_{i}).$
Set $A_{i}:=\alpha(\nu_{i})\in(0,\infty)$.
Step 3: normalized coefficients and definition of the filtrations. Set
$c_{i}:=\frac{2^{-i}}{A_{i}}.$
Since $\alpha(\operatorname{\mathrm{ord}}_{\mathfrak{m}})=1$ on the normalized slice and $\alpha$ is increasing, we have $A_{i}\geq 1$ and so $c_{i}\leq 2^{-i}$. Using (5.4), for all $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$,
$0\leq c_{i}\,v(J_{i})=c_{i}\,\alpha(v\wedge\nu_{i})\leq c_{i}\,\alpha(\nu_{i})=2^{-i}.$
Define ${\mathfrak{b}_{\bullet}}:=\mathfrak{m}^{\bullet}$, i.e., $\mathfrak{b}_{p}=\mathfrak{m}^{p}$. Define ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ by
(5.5) $\mathfrak{a}_{p}:=\mathfrak{m}^{p}\cdot\prod_{i\geq 1}J_{i}^{\lfloor p\,c_{i}\rfloor}.$
Since $c_{i}\leq 2^{-i}$, we have $\lfloor p\,c_{i}\rfloor=0$ for all $i\gg 1$, so the product is finite for each $p$. In particular, each $\mathfrak{a}_{p}$ is $\mathfrak{m}$–primary. Observe further that, for all $p,p^{\prime}\geq 1$ and all $i$, $\lfloor(p+p^{\prime})c_{i}\rfloor\geq\lfloor pc_{i}\rfloor+\lfloor p^{\prime}c_{i}\rfloor.$ Therefore,
$\mathfrak{a}_{p}\mathfrak{a}_{p^{\prime}}=\mathfrak{m}^{p+p^{\prime}}\cdot\prod_{i}J_{i}^{\lfloor pc_{i}\rfloor+\lfloor p^{\prime}c_{i}\rfloor}\subseteq\mathfrak{m}^{p+p^{\prime}}\cdot\prod_{i}J_{i}^{\lfloor(p+p^{\prime})c_{i}\rfloor}=\mathfrak{a}_{p+p^{\prime}}.$
Moreover, $\mathfrak{a}_{p+1}\subseteq\mathfrak{a}_{p}$, so $\mathfrak{a}_{\bullet}$} is a filtration of ideals.
Step 4: computation of $v({\mathfrak{a}_{\bullet}})$. Fix $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$. By definition, we have
$v(\mathfrak{a}_{p})=v(\mathfrak{m}^{p})+\sum_{i\geq 1}\lfloor p\,c_{i}\rfloor\,v(J_{i})=p\,v(\mathfrak{m})+\sum_{i\geq 1}\lfloor p\,c_{i}\rfloor\,v(J_{i}).$
Dividing by $p$,
$\frac{v(\mathfrak{a}_{p})}{p}=v(\mathfrak{m})+\sum_{i\geq 1}\frac{\lfloor p\,c_{i}\rfloor}{p}\,v(J_{i}).$
For each $i$, $\frac{\lfloor p\,c_{i}\rfloor}{p}\to c_{i}$ as $p\to\infty$, and
$0\leq\frac{\lfloor p\,c_{i}\rfloor}{p}\,v(J_{i})\leq c_{i}\,v(J_{i})\leq 2^{-i}.$
Since $\sum_{i\geq 1}2^{-i}<\infty$, dominated convergence applies and yields
$v({\mathfrak{a}_{\bullet}})=\lim_{p\to\infty}\frac{v(\mathfrak{a}_{p})}{p}=v(\mathfrak{m})+\sum_{i\geq 1}c_{i}\,v(J_{i})=v(\mathfrak{m})+\sum_{i\geq 1}c_{i}\,\alpha(v\wedge\nu_{i}).$
Since ${\mathfrak{b}_{\bullet}}=\mathfrak{m}^{\bullet}$ and $v(\mathfrak{m})=1$ on $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$, we have $v({\mathfrak{a}_{\bullet}})/v({\mathfrak{b}_{\bullet}})=v({\mathfrak{a}_{\bullet}})$ on this slice.
Step 5: unique maximizer and non-attainment by divisorial valuations. Because $\nu\wedge\nu_{i}=\nu_{i}$ for all $i$,
$\nu({\mathfrak{a}_{\bullet}})=1+\sum_{i\geq 1}c_{i}\,\alpha(\nu_{i})=1+\sum_{i\geq 1}2^{-i}=2.$
Let $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$ with $v\neq\nu$. Since $\nu$ is an end and $\nu_{i}\uparrow\nu$, there exists $i_{0}$ such that $v\not\succeq\nu_{i_{0}}$; otherwise $v\succeq\nu_{i}$ for all $i$ and so $v\succeq\nu$, forcing $v=\nu$. For this $i_{0}$ we have $v\wedge\nu_{i_{0}}\prec\nu_{i_{0}}$, and strict monotonicity of $\alpha$ along segments implies
$\alpha(v\wedge\nu_{i_{0}})<\alpha(\nu_{i_{0}}).$
For all $i$, $\alpha(v\wedge\nu_{i})\leq\alpha(\nu_{i})$, with strict inequality at $i=i_{0}$. Therefore,
$v({\mathfrak{a}_{\bullet}})=1+\sum_{i\geq 1}c_{i}\,\alpha(v\wedge\nu_{i})<1+\sum_{i\geq 1}c_{i}\,\alpha(\nu_{i})=2=\nu({\mathfrak{a}_{\bullet}}).$
Thus, $\nu$ is the unique maximizer of $v({\mathfrak{a}_{\bullet}})$ on $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$ and, hence, of $v({\mathfrak{a}_{\bullet}})/v({\mathfrak{b}_{\bullet}})$. Since $\nu$ is not divisorial, no divisorial valuation attains the supremum.
Step 6: linear boundedness. Since ${\mathfrak{b}_{\bullet}}=\mathfrak{m}^{\bullet}$, it is linearly bounded with respect to $\mathfrak{q}=\mathfrak{m}$. For $\mathfrak{a}_{\bullet}$}, the definition (5.5) gives $\mathfrak{m}^{p}\subseteq\mathfrak{a}_{p}$ for all $p$. Using $J_{i}\subseteq\mathfrak{m}$ and $\sum_{i}c_{i}\leq 1$, we obtain
$\mathfrak{a}_{p}\subseteq\mathfrak{m}^{p}\cdot\prod_{i}\mathfrak{m}^{\lfloor p\,c_{i}\rfloor}=\mathfrak{m}^{p+\sum_{i}\lfloor p\,c_{i}\rfloor}\subseteq\mathfrak{m}^{p+\lfloor p\sum_{i}c_{i}\rfloor}\subseteq\mathfrak{m}^{2p}.$
Hence, $\mathfrak{a}_{\bullet}$} is linearly bounded with respect to $\mathfrak{q}=\mathfrak{m}$, and the hypotheses of Theorem 5.14 are satisfied. This completes the proof. ∎
5.6. Finite maximizers in the ideal case
In the ideal case, i.e., when the two graded families of ideals are those of ordinary powers of given two ideals, we recover the classical stronger statement that the maximum is attained by finitely many Rees valuations.
Theorem 5.17 (Finite set of maximizers for ideals).
Let $(R,\mathfrak{m})$ be an excellent Noetherian local domain. Let $\mathfrak{a}$ be an ideal and let $\mathfrak{b}$ be any proper ideal with $\mathfrak{b}\neq 0$. Then,
$\mathcal{L}_{\mathfrak{b}}(\mathfrak{a})=\sup_{\begin{subarray}{c}\text{divisorial valuation }\\ v\text{ centered at }\mathfrak{m}\end{subarray}}\frac{v(\mathfrak{a})}{v(\mathfrak{b})}$
and, if finite, is a maximum and achieved by one of the Rees valuations of $\mathfrak{a}$. In particular, if finite, $\mathcal{L}_{\mathfrak{b}}(\mathfrak{a})$ is computed by finitely many divisorial valuations.
Proof.
This is a classical consequence of Rees’s theory of valuations (cf. [30, 49]). It also follows from Theorem 5.9 as we shall illustrate below.
Since $\mathcal{R}(\mathfrak{a})=\bigoplus_{p\geq 0}\mathfrak{a}^{p}T^{p}$ is the usual Rees algebra of $\mathfrak{a}$, the hypothesis of finite generation is automatic. Hence, by Theorem 5.9, there exists a finite set of divisorial valuations $\mathcal{V}(\mathfrak{a})=\{v_{1},\dots,v_{r}\}$, arising from prime divisors on the normalization of $\operatorname{Proj}\overline{\mathcal{R}(\mathfrak{a})}$, such that for any filtration $\mathfrak{b}_{\bullet}$}, with $v_{i}({\mathfrak{b}_{\bullet}})>0$, one has
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}}(\mathfrak{a}^{\bullet})=\max_{1\leq i\leq r}\frac{v_{i}(\mathfrak{a})}{v_{i}({\mathfrak{b}_{\bullet}})}.$
Specializing to ${\mathfrak{b}_{\bullet}}=\{\mathfrak{b}^{q}\}_{q\geq 1}$, noticing that $\mathfrak{b}\subseteq\mathfrak{m}$ so $\mathfrak{b}_{\bullet}$} is linearly bounded from below (as in Remark 5.10), yields the desired finite–maximum formula.
Finally, the valuations coming from the codimension–one components of the exceptional locus of the normalized blowup $\operatorname{Proj}\overline{\mathcal{R}(\mathfrak{a})}\to\operatorname{Spec}R$ are precisely the classical Rees valuations of $\mathfrak{a}$; see [30, Chapter 10]. Therefore, the maximum above may be taken over the Rees valuations of $\mathfrak{a}$, as claimed. ∎
Example 5.18.
Consider $\mathfrak{a}=(x^{5},x^{2}y^{3},y^{7})\subseteq R=\mathbb{C}[[x,y]]$. The Newton polyhedron of $\mathfrak{a}$ has two compact faces with inward normals $(1,1)$ and $(3,1)$. These correspond to Rees valuations $v_{(1,1)}$ and $v_{(3,1)}$ of $\mathfrak{a}$. The values of interest are:
$v_{(1,1)}(\mathfrak{a})=5,v_{(3,1)}(\mathfrak{a})=7,v(\mathfrak{m})=1.$
Hence, $\mathcal{L}_{\mathfrak{m}}(\mathfrak{a})=7$ is computed by a Rees valuations of $\mathfrak{a}$.
Part IV Structural Inequalities and Dualities
In this part, we derive structural inequalities for Łojasiewicz exponents and related invariants as consequences of the valuative framework developed in Part III.
6. Mixed multiplicities, Jacobian polygons, and Teissier–type inequalities
This section links Łojasiewicz-type containment thresholds to mixed multiplicities and Teissier’s convex-geometric inequalities. We restate Teissier–Rees–Sharp/Alexandrov–Fenchel type bounds for mixed multiplicities of ideals (Theorem 6.3). We then introduce mixed Łojasiewicz ratios and establish monotonicity/concavity properties that organize these inequalities (Proposition 6.6, Proposition 6.11). Finally, we specialize to Jacobian ideals and the algebraic avatars of Milnor and Tjurina numbers, obtaining Teissier-type bounds for the gradient Łojasiewicz exponent (Theorem 6.13), and we close with an extension principle that generalizes statements for ideals to filtrations.
6.1. Hilbert–Samuel multiplicity and mixed multiplicities
Throughout this section, $(R,\mathfrak{m})$ is a Noetherian local ring of dimension $d$. Assume, in addition, that $\dim\text{Nil}(\widehat{R})<d$, where $\text{Nil}(\cdot)$ represents the nil-radical of the ring (see [12]). This condition is satisfied, for instance, if $R$ is analytically unramified.
We recall the mixed multiplicities of ideals. Let $I_{1},\dots,I_{r}\subseteq R$ be ideals such that $I_{1}+\dots+I_{r}$ is $\mathfrak{m}$–primary (equivalently, $I_{1}^{n_{1}}\cdots I_{r}^{n_{r}}$ has finite colength for all large $(n_{i})$). Then, the function
$(n_{1},\dots,n_{r})\longmapsto\operatorname{\mathrm{length}}\!\left(R/(I_{1}^{n_{1}}\cdots I_{r}^{n_{r}})\right)\quad(n_{i}\in\mathbb{N})$
agrees with a polynomial of total degree $d$ for all $n_{i}\gg 0$. The degree–$d$ part of this polynomial can be uniquely written as
$\sum_{a_{1}+\cdots+a_{r}=d}\frac{e(I_{1}^{[a_{1}]},\dots,I_{r}^{[a_{r}]})}{a_{1}!\cdots a_{r}!}\,n_{1}^{a_{1}}\cdots n_{r}^{a_{r}},$
where the coefficients $e(I_{1}^{[a_{1}]},\dots,I_{r}^{[a_{r}]})\in\mathbb{N}$ are the mixed multiplicities of $I_{1},\dots,I_{r}$.
We will primarily use the two-ideal case $(I,J)$, where we write $e(I^{[i]},J^{[d-i]})\quad(0\leq i\leq d).$ The following are standard facts:
  1. (a)
    (Scaling) For $p,q\geq 1$,
    $e((I^{p})^{[i]},(J^{q})^{[d-i]})=p^{\,i}q^{\,d-i}\,e(I^{[i]},J^{[d-i]}).$
  2. (b)
    (Monotonicity) If $I\subseteq I^{\prime}$ and $J\subseteq J^{\prime}$ are $\mathfrak{m}$–primary, then
    $e(I^{[i]},J^{[d-i]})\ \geq\ e((I^{\prime})^{[i]},(J^{\prime})^{[d-i]})\quad\text{for all }i.$
  3. (c)
    (Integral closure invariance) If $\overline{I}=\overline{I^{\prime}}$ and $\overline{J}=\overline{J^{\prime}}$, then
    $e(I^{[i]},J^{[d-i]})=e((I^{\prime})^{[i]},(J^{\prime})^{[d-i]})\quad\text{for all }i.$
Definition 6.1 (Mixed multiplicities of filtrations).
Let ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ and ${\mathfrak{b}_{\bullet}}=\{\mathfrak{b}_{q}\}_{q\geq 1}$ be graded families of $\mathfrak{m}$-primary ideals. For $1\leq i\leq d$, set
$e_{i}({\mathfrak{a}_{\bullet}}^{[i]},{\mathfrak{b}_{\bullet}}^{[d-i]})=\limsup_{p\rightarrow\infty}\frac{e(\mathfrak{a}_{p}^{[i]},\mathfrak{b}_{p}^{[d-i]})}{n^{d}}.$
It was shown by Cid-Ruiz–Montaño [10] that, under the condition that $\dim\text{Nil}(\widehat{R})<d$, $e_{i}({\mathfrak{a}_{\bullet}}^{[i]},{\mathfrak{b}_{\bullet}}^{[d-i]})$ exists as an actual limit. When ${\mathfrak{b}_{\bullet}}=\{\mathfrak{m}^{q}\}_{q\geq 1}$, we write $e_{i}({\mathfrak{a}_{\bullet}})$ for $e({\mathfrak{a}_{\bullet}}^{[i]},{\mathfrak{b}_{\bullet}}^{[d-i]})$.
Definition 6.2 (Mixed Łojasiewicz rations).
For a graded family $\mathfrak{a}_{\bullet}$} of $\mathfrak{m}$-primary ideals, set
$\Lambda_{i}({\mathfrak{a}_{\bullet}})=\frac{e_{i}({\mathfrak{a}_{\bullet}})}{e_{i-1}({\mathfrak{a}_{\bullet}})},\ 1\leq i\leq d.$
6.2. Teissier–Rees–Sharp and Alexandrov–Fenchel inequalities
We begin with the Alexandrov–Fenchel type inequalities for mixed multiplicities in the form best suited for graded families (or filtrations).
Theorem 6.3 (Teissier–Rees–Sharp inequality).
Assume that $R$ is quasi–unmixed. Let ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ and ${\mathfrak{b}_{\bullet}}=\{\mathfrak{b}_{q}\}_{q\geq 1}$ be graded families such that $\mathfrak{a}_{p}+\mathfrak{b}_{q}$ is $\mathfrak{m}$–primary for all $p,q\geq 1$. Then, for all $p,q\geq 1$, and $1\leq i\leq d-1$, we have
$e\!\left(\mathfrak{a}_{p}^{[i]},\mathfrak{b}_{q}^{[d-i]}\right)^{2}\ \geq\ e\!\left(\mathfrak{a}_{p}^{[i-1]},\mathfrak{b}_{q}^{[d-i+1]}\right)\,e\!\left(\mathfrak{a}_{p}^{[i+1]},\mathfrak{b}_{q}^{[d-i-1]}\right).$
Equivalently, for fixed $p,q$, the sequence $i\ \longmapsto\ e\!\left(\mathfrak{a}_{p}^{[i]},\mathfrak{b}_{q}^{[d-i]}\right)$ is log-concave.
Proof.
The stated inequality is precisely the Alexandrov-Fenchel type inequality for mixed multiplicities. It was first established by Teissier [50] in the $\mathfrak{m}$-primary case via Minkowski inequalities for multiplicities, and was subsequently extended by Rees and Sharp [46] to the general situation where $\mathfrak{a}_{p}+\mathfrak{b}_{q}$ is $\mathfrak{m}$-primary, under the quasi-unmixed hypothesis of $R$. ∎
6.3. Containments $\Rightarrow$ mixed-multiplicity inequalities
Following the foundational work of Teissier on mixed multiplicities, several authors – most notably Bivià-Ausina and Encinas [1, 3] – have related Lojasiewicz-type invariants to mixed multiplicity data, typically through inequalities involving integral closures of powers of ideals. The proposition below isolates the underlying numerical mechanism in a containment-theoretic form.
Proposition 6.4 (Containment implies mixed-multiplicity bounds).
Let ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ and ${\mathfrak{b}_{\bullet}}=\{\mathfrak{b}_{q}\}_{q\geq 1}$ be $\mathfrak{m}$–primary graded families. Fix $p,q\geq 1$ and assume that
$\mathfrak{b}_{qt}\ \subseteq\ \overline{\mathfrak{a}_{pt}}\quad\text{for all }t\gg 1.$
Then, for every $1\leq i\leq d$ and every $t\gg 1$, one has
$e\!\left((\mathfrak{b}_{qt})^{[i]},\mathfrak{m}^{[d-i]}\right)\geq e\!\left((\mathfrak{a}_{pt})^{[i]},\mathfrak{m}^{[d-i]}\right).$
Consequently,
$e_{i}({\mathfrak{b}_{\bullet}})\ \geq\ \left(\frac{p}{q}\right)^{\!i}\,e_{i}({\mathfrak{a}_{\bullet}}),$
and in particular,
$\frac{q}{p}\ \geq\ \left(\frac{e_{i}({\mathfrak{a}_{\bullet}})}{e_{i}({\mathfrak{b}_{\bullet}})}\right)^{1/i}\text{ for all }1\leq i\leq d.$
Proof.
Fix $t\gg 1$ such that $\mathfrak{b}_{qt}\subseteq\overline{\mathfrak{a}_{pt}}$. By monotonicity and integral-closure invariance standard facts,
$e((\mathfrak{b}_{qt})^{[i]},\mathfrak{m}^{[d-i]})\ \geq\ e((\overline{\mathfrak{a}_{pt}})^{[i]},\mathfrak{m}^{[d-i]})=e((\mathfrak{a}_{pt})^{[i]},\mathfrak{m}^{[d-i]}),$
which proves the first displayed inequality.
The asymptotic statements is then obtained by dividing both sides of this inequality by $(qt)^{i}$ and taking the limit as $t\to\infty$. ∎
Corollary 6.5.
Let $I,J$ be $\mathfrak{m}$–primary ideals in $R$, and let $p,q\geq 1$. If $J^{\,q}\subseteq\overline{I^{\,p}}$, then for every $1\leq i\leq d$ one has
$q^{\,i}\,e(J^{[i]},\mathfrak{m}^{[d-i]})\ \geq\ p^{\,i}\,e(I^{[i]},\mathfrak{m}^{[d-i]}),$
equivalently,
$\frac{q}{p}\ \geq\ \left(\frac{e(I^{[i]},\mathfrak{m}^{[d-i]})}{e(J^{[i]},\mathfrak{m}^{[d-i]})}\right)^{1/i}.$
Proof.
This inequality is a classical consequence of Teissier’s mixed-multiplicity theory and appears in the algebraic form in Rees–Sharp; see [50, 46]. It is also a direct consequence of Proposition 6.4. ∎
An immediate consequence of Corollary 6.5, taking $i=d$, gives the following lower bound:
$\mathcal{L}_{J}(I)\ \geq\ \left(\frac{e(I)}{e(J)}\right)^{1/d}.$
6.4. Mixed Łojasiewicz ratios and Teissier-type monotonicity
The Alexandrov–Fenchel inequalities imply strong structure for certain ratios of mixed multiplicities. These ratios play the role of “mixed slopes” in the classical theory.
Proposition 6.6 (Log-concavity and monotonicity of $\Lambda_{i}$).
Let ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ be an $\mathfrak{m}$–primary graded family. Then, mixed multiplicity ratios satisfy
$\Lambda_{1}({\mathfrak{a}_{\bullet}})\ \geq\ \Lambda_{2}({\mathfrak{a}_{\bullet}})\ \geq\ \cdots\ \geq\ \Lambda_{d}({\mathfrak{a}_{\bullet}}).$
Proof.
Fix $p,t$ and apply Theorem 6.3 to the pair $(\mathfrak{a}_{pt},\mathfrak{m})$, we get
[…]
It follows that $\Lambda_{i}(\mathfrak{a}_{pt})\geq\Lambda_{i+1}(\mathfrak{a}_{pt})$. Dividing both sides of this inequality by appropriate powers of $pt$ and taking the limit along $p\to\infty$ then give $\Lambda_{1}({\mathfrak{a}_{\bullet}})\geq\Lambda_{2}({\mathfrak{a}_{\bullet}})\geq\cdots\geq\Lambda_{d}({\mathfrak{a}_{\bullet}})$ as claimed. ∎
Corollary 6.7 (Teissier monotonicity).
Let $I$ be $\mathfrak{m}$–primary. Then, the sequence $(\Lambda_{i}(I))_{i=1}^{d}$ is nonincreasing:
$\Lambda_{1}(I)\ \geq\ \Lambda_{2}(I)\ \geq\ \cdots\ \geq\ \Lambda_{d}(I).$
Equivalently, the sequence $i\longmapsto e(I^{[i]},\mathfrak{m}^{[d-i]})$ is log-concave.
Proof.
Apply Proposition 6.6 to the power family ${\mathfrak{a}_{\bullet}}=\{I^{p}\}$. This monotonicity/log-concavity was established by Teissier in [50]. ∎
6.5. Jacobian ideals, Milnor and Tjurina numbers (algebraic viewpoint)
We now explain how classical “Jacobian polygon” data can be encoded algebraically via mixed multiplicities.
Assume in this subsection that $R=\mathcal{O}_{\mathbb{C}^{d},0}$ and let $f\in\mathfrak{m}$ define an isolated singularity at the origin. Let $J(f)\subseteq R$ be the Jacobian ideal generated by $\partial f/\partial x_{1},\dots,\partial f/\partial x_{d}$.
Definition 6.8 (Milnor and Tjurina numbers).
Define the Milnor number and Tjurina number by
$\mu(f):=\dim_{\mathbb{C}}R/J(f),\quad\tau(f):=\dim_{\mathbb{C}}R/(f,J(f)).$
The next object is the algebraic substitute for Teissier’s Jacobian polygon.
Definition 6.9 (Jacobian mixed multiplicities and polygon).
For $0\leq i\leq d$ set
$m_{i}(f):=e\bigl((f)^{[d-i]},\,J(f)^{[i]}\bigr).$
The Jacobian polygon $\text{JP}(f)$ of $f$ is the lower convex hull in $\mathbb{R}^{2}$ of the points
$(0,m_{0}(f)),\ (1,m_{1}(f)),\ \dots,\ (d,m_{d}(f)).$
Remark 6.10.
In Teissier’s work, the slopes and vertices of the Jacobian polygon encode polar multiplicities and vanishing rates. In our framework, the same numerical data are captured by the mixed multiplicities $m_{i}(f)$, which are defined for any Noetherian local ring once $(f)+J(f)$ is $\mathfrak{m}$–primary.
Proposition 6.11 (Concavity of the Jacobian polygon).
Let $f$ define an isolated singularity. Then, the sequence $i\mapsto\log m_{i}(f)$ is concave, equivalently
$m_{i}(f)^{2}\ \geq\ m_{i-1}(f)\,m_{i+1}(f)\quad(1\leq i\leq d-1).$
Proof.
Apply Theorem 6.3 to the pair of ideals $((f),J(f))$, noticing that $(f)+J(f)$ is $\mathfrak{m}$–primary. ∎
Corollary 6.12.
Let $f$ define an isolated singularity, and let $m_{i}(f)$ be as in Definition 6.9. Then, the sequence of successive ratios
$\frac{m_{1}(f)}{m_{0}(f)},\ \frac{m_{2}(f)}{m_{1}(f)},\ \dots,\ \frac{m_{d}(f)}{m_{d-1}(f)}$
is nonincreasing. Equivalently, for $1\leq i\leq d-1$ one has
$\frac{m_{i}(f)}{m_{i-1}(f)}\ \geq\ \frac{m_{i+1}(f)}{m_{i}(f)}.$
Proof.
By Proposition 6.11, we have
$m_{i}(f)^{2}\ \geq\ m_{i-1}(f)\,m_{i+1}(f)\quad(1\leq i\leq d-1).$
Dividing both sides by $m_{i-1}(f)\,m_{i}(f)$ (which is positive) gives
$\frac{m_{i}(f)}{m_{i-1}(f)}\ \geq\ \frac{m_{i+1}(f)}{m_{i}(f)}.\qed$
6.6. Teissier-type bounds for the gradient Łojasiewicz exponent
We now connect $L(f,0):=\mathcal{L}^{an}_{\mathfrak{m}}(J(f))=\mathcal{L}_{\mathfrak{m}}(J(f))$ to the mixed multiplicity data. The lower bound in our next result, when $i=d$, recovers the classical estimate in terms of $e(J(f))$, while the full family of inequalities reflects finer mixed-multiplicity data of the Jacobian ideal.
Theorem 6.13 (Mixed-multiplicity lower bounds for $L(f,0)$).
Let $f\in\mathcal{O}_{\mathbb{C}^{d},0}$ define an isolated singularity. Then, for each $1\leq i\leq d$,
[…]
In particular,
$L(f,0)\geq\ \left(\frac{e(J(f))}{e(\mathfrak{m})}\right)^{1/d}.$
Proof.
For every $t\gg 1$, the rational number $\frac{\lceil t\,L(f,0)\rceil}{t}$ is admissible for $\mathcal{L}_{\mathfrak{m}}(J(f))$, and hence
$\mathfrak{m}^{\lceil t\,L(f,0)\rceil}\ \subseteq\ \overline{J(f)^{\,t}}.$
Apply Corollary 6.5 with $(I,J)=(J(f),\mathfrak{m})$ and $(p,q)=(t,\lceil tL(f,0)\rceil)$. For each $1\leq i\leq d$ we obtain
$\lceil t\,L(f,0)\rceil^{\,i}\,e(\mathfrak{m}^{[i]},\mathfrak{m}^{[d-i]})\ \geq\ t^{\,i}\,e(J(f)^{[i]},\mathfrak{m}^{[d-i]}).$
Divide by $t^{i}$ and let $t\to\infty$ to conclude
$L(f,0)^{\,i}\,e(\mathfrak{m}^{[i]},\mathfrak{m}^{[d-i]})\ \geq\ e(J(f)^{[i]},\mathfrak{m}^{[d-i]}),$
which is equivalent to the stated lower bound. ∎
Remark 6.14.
Classical analytic estimates bounding $L(f,0)$ from below by numerical invariants (Milnor numbers, polar multiplicities, etc.) can be viewed as instances of Theorem 6.13 together with identifications between those invariants and mixed multiplicities; see [1, 3].
6.7. Extension principle: from ideals to filtrations
We will restrict to the case when the filtration is asymptotically integral-closure-equivalent to powers of an $\mathfrak{m}$–primary ideal, i.e., exhibiting linearly bounded behavior as in Theorem 5.14.
Definition 6.15 (Linearly bounded).
A filtration ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}$ of $\mathfrak{m}$–primary ideals is linearly bounded if there exists an $\mathfrak{m}$–primary ideal $I$ and constants $c,C>0$ such that
$\overline{I^{\lceil cp\rceil}}\subseteq\overline{\mathfrak{a}_{p}}\subseteq\overline{I^{\lfloor Cp\rfloor}}\quad\text{for all }p\geq 1.$
Remark 6.16.
Noetherian graded families (i.e., finitely generated Rees algebras) and many Newton-type filtrations are linearly bounded.
Proposition 6.17 (Mixed multiplicities for linearly bounded filtrations).
Let $(R,\mathfrak{m})$ be a Noetherian local ring of dimension $d$. Let ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ and ${\mathfrak{b}_{\bullet}}=\{\mathfrak{b}_{q}\}_{q\geq 1}$ be filtrations of $\mathfrak{m}$–primary ideals. Assume that $\mathfrak{a}_{\bullet}$} is linearly bounded by an $\mathfrak{m}$–primary ideal $I$ with constants $c_{\mathfrak{a}},C_{\mathfrak{a}}>0$ and that $\mathfrak{b}_{\bullet}$} is linearly bounded by an $\mathfrak{m}$–primary ideal $J$ with constants $c_{\mathfrak{b}},C_{\mathfrak{b}}>0$, in the sense of Definition 6.15, i.e.,
$\overline{I^{\lceil c_{\mathfrak{a}}p\rceil}}\subseteq\overline{\mathfrak{a}_{p}}\subseteq\overline{I^{\lfloor C_{\mathfrak{a}}p\rfloor}},\quad\overline{J^{\lceil c_{\mathfrak{b}}q\rceil}}\subseteq\overline{\mathfrak{b}_{q}}\subseteq\overline{J^{\lfloor C_{\mathfrak{b}}q\rfloor}}\quad\text{for all }p,q\geq 1.$
If $\mathfrak{b}_{q}\subseteq\overline{\mathfrak{a}_{p}}$ for some integers $p,q\geq 1$, then setting
$\alpha:=\lceil c_{\mathfrak{b}}q\rceil,\quad\beta:=\lfloor C_{\mathfrak{a}}p\rfloor,$
we have, for each $1\leq i\leq d$,
$\left(\frac{\alpha}{\beta}\right)^{i}\ \geq\ \frac{e\bigl(I^{[i]},\mathfrak{m}^{[d-i]}\bigr)}{e\bigl(J^{[i]},\mathfrak{m}^{[d-i]}\bigr)}.$
Equivalently,
$\frac{\alpha}{\beta}\ \geq\ \left(\frac{e\bigl(I^{[i]},\mathfrak{m}^{[d-i]}\bigr)}{e\bigl(J^{[i]},\mathfrak{m}^{[d-i]}\bigr)}\right)^{1/i}.$
Proof.
Assume $\mathfrak{b}_{q}\subseteq\overline{\mathfrak{a}_{p}}$. Then, $\overline{\mathfrak{b}_{q}}\subseteq\overline{\mathfrak{a}_{p}}$. By the linear boundedness hypotheses,
$\overline{J^{\alpha}}\ \subseteq\ \overline{\mathfrak{b}_{q}}\ \subseteq\ \overline{\mathfrak{a}_{p}}\ \subseteq\ \overline{I^{\beta}}.$
In particular, $J^{\alpha}\subseteq\overline{I^{\beta}}$. Apply Corollary 6.5 to the containment $J^{\alpha}\subseteq\overline{I^{\beta}}$ (with $(I,J)$ there replaced by $(I,J)$ here and $(p,q)$ replaced by $(\beta,\alpha)$): for each $1\leq i\leq d$ we obtain
$\frac{\alpha}{\beta}\ \geq\ \left(\frac{e\bigl(I^{[i]},\mathfrak{m}^{[d-i]}\bigr)}{e\bigl(J^{[i]},\mathfrak{m}^{[d-i]}\bigr)}\right)^{1/i},$
and raising both sides to the ii–th power gives the stated inequality. ∎
Corollary 6.18 (Explicit slope bounds in terms of $q/p$).
In the setting of Proposition 6.17, fix $1\leq i\leq d$ and set
$r_{i}:=\left(\frac{e\!\left(I^{[i]},\mathfrak{m}^{[d-i]}\right)}{e\!\left(J^{[i]},\mathfrak{m}^{[d-i]}\right)}\right)^{\!1/i}.$
If $\mathfrak{b}_{q}\subseteq\overline{\mathfrak{a}_{p}}$ for some $p>1/C_{\mathfrak{a}}$ and $q\geq 1$, then
(6.1) $\frac{q}{p}\ \geq\ \frac{C_{\mathfrak{a}}}{c_{\mathfrak{b}}}\,r_{i}\ -\ \frac{r_{i}+1}{c_{\mathfrak{b}}\,p}.$
In particular, for any $\varepsilon>0$, if $p\geq\dfrac{r_{i}+1}{c_{\mathfrak{b}}\,\varepsilon}$ then
$\frac{q}{p}\ \geq\ \frac{C_{\mathfrak{a}}}{c_{\mathfrak{b}}}\,r_{i}-\varepsilon.$
Proof.
Let $\alpha=\lceil c_{\mathfrak{b}}q\rceil$ and $\beta=\lfloor C_{\mathfrak{a}}p\rfloor$ as in Proposition 6.17. That proposition gives
$\frac{\alpha}{\beta}\ \geq\ r_{i}.$
Using $\alpha\leq c_{\mathfrak{b}}q+1$ and $\beta\geq C_{\mathfrak{a}}p-1$ (note that $p>1/C_{\mathfrak{a}}$ guarantees $\beta\geq 0$), inequality $(\ast)$ implies
$\frac{c_{\mathfrak{b}}q+1}{C_{\mathfrak{a}}p-1}\ \geq\ r_{i},$
hence
$c_{\mathfrak{b}}q+1\ \geq\ r_{i}(C_{\mathfrak{a}}p-1)=r_{i}C_{\mathfrak{a}}p-r_{i},$
so
[…]
Dividing by $p$ gives (6.1), and the final statement follows immediately. ∎
7. Log canonical thresholds and generic hyperplane sections
In this section we relate the Lojasiewicz exponent to another fundamental valuative threshold invariant, the asymptotic log canonical threshold. Particularly, we compare the valuative descriptions of $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})$ and $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$, and show that they satisfy the optimal bound $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})\cdot\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\geq\operatorname{\mathrm{lct}}({\mathfrak{b}_{\bullet}})$ (Theorem 7.1). This inequality complements the Teissier–type estimates established in Section 6 and identifies a natural equality regime governed by a common extremal valuation. Finally, we establish a generic hyperplane–section principle for filtrations, showing that admissible containment ratios descend and that the associated sectional Łojasiewicz exponents are monotone under restriction along a general flag (Corollary 7.6).
7.1. Valuative definition of the log canonical threshold
We shall work with a regular local ring $(R,\mathfrak{m})$ of dimension $d$ (essentially of finite type over a field of characteristic 0). For an ideal $\mathfrak{a}\subseteq R$, the log canonical threshold of $\mathfrak{a}$, denoted by $\operatorname{\mathrm{lct}}(\mathfrak{a})$, is an important invariant that has been much studied in both singularity theory and birational geometry (cf. [32, 35]). Particularly, if $\mathfrak{a}=(f_{1},\dots,f_{s})$ then
$\operatorname{\mathrm{lct}}(\mathfrak{a})=\sup\{\theta>0:\big(\sum_{i}|f_{i}|^{2}\big)^{-\theta}\text{ is locally integrable}\}.$
Let $A(v)$ denote the log discrepancy of a valuation $v$ (see [5, 32, 35] for more details on log discrepancy and log canonical threshold of multiplier ideals). Then, $\operatorname{\mathrm{lct}}(\mathfrak{a})$ is also given by
$\operatorname{\mathrm{lct}}(\mathfrak{a})=\inf_{v}\frac{A(v)}{v(\mathfrak{a})},$
where $v$ ranges over all divisorial valuations of $R$ centered at $\mathfrak{m}$.
For a graded family ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ of ideals in $R$, the asymptotic log canonical threshold of $\mathfrak{a}_{\bullet}$}, denoted by $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})$, was defined and studied by Jonsson–Mustaţǎ [31]. Particularly,
$\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})=\inf_{v}\frac{A(v)}{v({\mathfrak{a}_{\bullet}})},$
where $v$ ranges over all divisorial valuations of $R$ centered at $\mathfrak{m}$. It was observed in [31] that, in general, there may not exist a divisorial valuation $v$ realizing $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})$. On the other hand, it was also shown in [31, Theorem A], that given any graded family $\mathfrak{a}_{\bullet}$}, there is a real valuation $v$ of $R$ centered at $\mathfrak{m}$ that computes $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})$, i.e., $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})=\frac{A(v)}{v({\mathfrak{a}_{\bullet}})}$.
7.2. A sharp inequality: $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})\cdot\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\geq\operatorname{\mathrm{lct}}({\mathfrak{b}_{\bullet}})$
We now prove the key “dual” inequality and isolate the rigidity mechanism. In the analytic setting, when $R=\mathcal{O}_{\mathbb{C}^{d},0}$, ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}^{p}\}_{p\geq 1}$ and ${\mathfrak{b}_{\bullet}}=\{\mathfrak{m}^{q}\}_{q\geq 1}$, this inequality was known by Bivià-Ausina–Fukui [4]. Our point is that the same inequality holds in greater generality and fits naturally into the valuative min-max framework developed in this paper. We also characterize the equality in the next section.
Theorem 7.1 (Sharp inequality).
Let $(R,\mathfrak{m})$ be regular local of dimension $d$. Let $\mathfrak{a}_{\bullet}$} and $\mathfrak{b}_{\bullet}$} be graded families of ideals in $R$. Then,
$\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})\cdot\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\geq\operatorname{\mathrm{lct}}({\mathfrak{b}_{\bullet}}).$
Particularly, $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})\cdot\mathcal{L}_{\mathfrak{m}}({\mathfrak{a}_{\bullet}})\geq d$. Equivalently,
$\Theta({\mathfrak{a}_{\bullet}}):=\frac{1}{d}\,\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})\cdot\mathcal{L}_{\mathfrak{m}}({\mathfrak{a}_{\bullet}})\ \geq\ 1.$
Proof.
By the valuative formulas and Theorem 5.6,
$\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})=\inf_{v}\frac{A(v)}{v({\mathfrak{a}_{\bullet}})}\quad\text{and}\quad\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\geq\sup_{v}\frac{v({\mathfrak{a}_{\bullet}})}{v({\mathfrak{b}_{\bullet}})},$
where $v$ ranges over divisorial valuations centered at $\mathfrak{m}$. Therefore,
$\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})\cdot\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\geq\Bigl(\inf_{v}\frac{A(v)}{v({\mathfrak{a}_{\bullet}})}\Bigr)\cdot\Bigl(\sup_{w}\frac{w({\mathfrak{a}_{\bullet}})}{w({\mathfrak{b}_{\bullet}})}\Bigr)\ \geq\ \inf_{v}\Bigl(\frac{A(v)}{v({\mathfrak{a}_{\bullet}})}\cdot\frac{v({\mathfrak{a}_{\bullet}})}{v({\mathfrak{b}_{\bullet}})}\Bigr)=\inf_{v}\frac{A(v)}{v({\mathfrak{b}_{\bullet}})}=\operatorname{\mathrm{lct}}({\mathfrak{b}_{\bullet}}).$
When ${\mathfrak{b}_{\bullet}}=\{\mathfrak{m}^{q}\}_{q\geq 1}$, we have $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})\cdot\mathcal{L}_{\mathfrak{m}}({\mathfrak{a}_{\bullet}})\geq\operatorname{\mathrm{lct}}(\mathfrak{m})$. On the other hand, since $R$ is regular of dimension $d$, one has $\operatorname{\mathrm{lct}}(\mathfrak{m})=d$; see, for instance [35, § 9]. Thus, $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})\cdot\mathcal{L}_{\mathfrak{m}}({\mathfrak{a}_{\bullet}})\geq d$. ∎
7.3. Generic hyperplane sections and the Łojasiewicz exponent
Let $(R,\mathfrak{m})$ be a Noetherian local ring with infinite residue field.
Definition 7.2 (Eventual degree–one normality).
An $\mathfrak{m}$–primary filtration ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ is said to be eventually degree–one normal if there exists an integer $p_{0}\geq 1$ such that
$\overline{\mathfrak{a}_{p}}=\overline{\mathfrak{a}_{1}^{\,p}}\quad\text{for all }p\geq p_{0}.$
Remark 7.3 (Specialization of integral closure).
Let $I$ be an ideal of height at least $2$ in a Noetherian ring $R$ which is normal, locally equidimensional, universally catenary, and such that $R_{\mathrm{red}}$ is locally analytically unramified. The following statements hold:
  1. (i)
    ([28, Theorem 2.1]) Integral closure is compatible with specialization by a generic element of $I$.
  2. (ii)
    ([26, Proposition 3.7]) For a general element $x\in I$, one has
    $\overline{I^{n}}\,R/(x)=\overline{I^{n}R/(x)}\quad\text{for all }n\gg 0.$
The following result is an application of specialization of integral closure to filtrations of ideals.
Theorem 7.4 (Stable–range commutation for $\overline{\mathfrak{a}_{pt}}$).
Assume that $R$ satisfies the hypotheses of Remark 7.3. Let $\mathfrak{a}_{\bullet}$} be an $\mathfrak{m}$–primary filtration that is eventually degree-one normal, and set $I=\mathfrak{a}_{1}$. Then, there exists a Zariski open dense subset $U\subseteq I/\mathfrak{m}I$ such that for every $x\in I$ with class in $U$, and every integer $p\geq 1$, there exists an integer $t_{0}=t_{0}(p)$ with the property that
$\overline{\mathfrak{a}_{pt}}\,R^{\prime}=\overline{\mathfrak{a}_{pt}R^{\prime}},\text{ for all }t\geq t_{0},\text{ where }R^{\prime}=R/(x).$
Proof.
Fix $p\geq 1$. By eventual degree–one normality, there exists $t_{1}(p)$ such that $\overline{\mathfrak{a}_{pt}}=\overline{I^{pt}}$ for all $t\geq t_{1}(p)$. Since $\mathfrak{a}_{\bullet}$} is graded, $I^{pt}\subseteq\mathfrak{a}_{pt}$, and so for such $t$,
$I^{pt}\subseteq\mathfrak{a}_{pt}\subseteq\overline{I^{pt}}.$
Let $x\in I$ be general as in Remark 7.3(ii) and set $R^{\prime}=R/(x)$. By [26, Proposition 3.7], there exists $n_{0}$ such that
$\overline{I^{n}}\,R^{\prime}=\overline{I^{n}R^{\prime}}\quad\text{for all }n\geq n_{0}.$
Taking $n=pt$, this equality holds for all $t\geq t_{2}(p):=\lceil n_{0}/p\rceil$.
For $t\geq\max\{t_{1}(p),t_{2}(p)\}$ we therefore have
$\overline{\mathfrak{a}_{pt}}\,R^{\prime}=\overline{I^{pt}}\,R^{\prime}=\overline{I^{pt}R^{\prime}}=\overline{\mathfrak{a}_{pt}R^{\prime}},$
where the last equality uses the inclusions $I^{pt}R^{\prime}\subseteq\mathfrak{a}_{pt}R^{\prime}\subseteq\overline{I^{pt}}\,R^{\prime}$ and the fact that integral closure depends only on the integral–closure class of the ideal. ∎
Corollary 7.5 (Admissible ratios descend).
Assume the hypotheses of Theorem 7.4. Let ${\mathfrak{b}_{\bullet}}=\{\mathfrak{b}_{q}\}_{q\geq 1}$ be an arbitrary filtration. Fix $p,q\geq 1$, and let $x\in\mathfrak{a}_{1}$ be a general element, and set $R^{\prime}=R/(x)$. If
$\mathfrak{b}_{qt}\subseteq\overline{\mathfrak{a}_{pt}}\quad\text{for }t\gg 1,$
then
$\mathfrak{b}_{qt}R^{\prime}\subseteq\overline{\mathfrak{a}_{pt}R^{\prime}}\quad\text{for }t\gg 1.$
Proof.
For $t\gg 1$, quotienting gives $\mathfrak{b}_{qt}R^{\prime}\subseteq\overline{\mathfrak{a}_{pt}}\,R^{\prime}$. For the fixed integer $p$, Theorem 7.4 yields $\overline{\mathfrak{a}_{pt}}\,R^{\prime}=\overline{\mathfrak{a}_{pt}R^{\prime}}$ for all $t\gg 1$, which proves the claim. ∎
Corollary 7.6 (Monotonicity under generic hyperplane section).
Under the same assumptions as in Corollary 7.5,
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}R^{\prime}}({\mathfrak{a}_{\bullet}}R^{\prime})\;\leq\;\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}}).$
Proof.
Every ratio $\frac{q}{p}$ admissible for $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$ remains admissible after restriction by Corollary 7.5. Taking infima gives the inequality. ∎
Iterating Corollary 7.6 along a general flag $x_{1},\dots,x_{d-k}\in\mathfrak{a}_{1}$ yields the monotonicity of the sectional Łojasiewicz exponents
$\mathcal{L}^{(d)}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\ \geq\ \mathcal{L}^{(d-1)}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\ \geq\ \cdots\ \geq\ \mathcal{L}^{(1)}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}}),$
where the sectional invariants are defined by restriction to $R/(x_{1},\dots,x_{d-k})$.
Part V Structural Consequences of the Finite-Max Principle
We develop rigidity and stratification phenomena that follow once the Łojasiewicz exponent is computed by a finite candidate set of divisorial valuations.
8. Rigidity results
The theme of this section is that equality in the structural inequalities developed in Sections 6 and 7 is highly rigid: it forces a common extremal valuation to compute several a priori different invariants simultaneously. We first formulate a general common-extremal-valuation principle (Theorem 8.1) and then analyze its consequences in increasingly structured settings, including a complete classification in the monomial case (Theorem 8.3) and extensions to Newton nondegenerate regimes. We also treat the equality case in Teissier–Rees–Sharp type inequalities, showing that proportionality/equality of the relevant mixed-multiplicity data is detected valuatively and leads to strong constraints on the underlying ideals and filtrations (e.g., Proposition 8.7).
8.1. Rigidity I: the common extremal valuation principle
We begin with the basic mechanism: equality in Theorem 7.1 forces the same divisorial valuation to compute all extremal quantities involved.
Theorem 8.1 (Common extremal valuation criterion).
With hypotheses as in Theorem 7.1, the equality
$\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})\cdot\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})=\operatorname{\mathrm{lct}}({\mathfrak{b}_{\bullet}})$
holds if and only if there exists real valuation $v$ centered at $\mathfrak{m}$ such that:
  1. (a)
    $v$ computes $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})$, i.e. $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})=A(v)/v({\mathfrak{a}_{\bullet}})$;
  2. (b)
    $v$ computes $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$, i.e. $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})=v({\mathfrak{a}_{\bullet}})/v({\mathfrak{b}_{\bullet}})$;
  3. (c)
    $v$ computes $\operatorname{\mathrm{lct}}({\mathfrak{b}_{\bullet}})$, i.e. $\operatorname{\mathrm{lct}}({\mathfrak{b}_{\bullet}})=A(v)/v({\mathfrak{b}_{\bullet}})$.
Proof.
If. If $v$ satisfies (a)–(c), then
$\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})\cdot\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})=\frac{A(v)}{v({\mathfrak{a}_{\bullet}})}\cdot\frac{v({\mathfrak{a}_{\bullet}})}{v({\mathfrak{b}_{\bullet}})}=\frac{A(v)}{v({\mathfrak{b}_{\bullet}})}=\operatorname{\mathrm{lct}}({\mathfrak{b}_{\bullet}}).$
Only if. Assume $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})\cdot\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})=\operatorname{\mathrm{lct}}({\mathfrak{b}_{\bullet}})$. Choose a real valuation $v_{0}$ centered at $\mathfrak{m}$ that computes $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})$, whose existence was established in [31, Theorem A]. That is, $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})=A(v_{0})/v_{0}({\mathfrak{a}_{\bullet}})$. By the valuative definition of $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$ in Theorem 5.6, we have
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\geq\frac{v_{0}({\mathfrak{a}_{\bullet}})}{v_{0}({\mathfrak{b}_{\bullet}})}.$
Hence,
$\operatorname{\mathrm{lct}}({\mathfrak{b}_{\bullet}})=\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})\cdot\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\ \geq\ \frac{A(v_{0})}{v_{0}({\mathfrak{a}_{\bullet}})}\cdot\frac{v_{0}({\mathfrak{a}_{\bullet}})}{v_{0}({\mathfrak{b}_{\bullet}})}=\frac{A(v_{0})}{v_{0}({\mathfrak{b}_{\bullet}})}\ \geq\ \inf_{v}\frac{A(v)}{v({\mathfrak{b}_{\bullet}})}=\operatorname{\mathrm{lct}}({\mathfrak{b}_{\bullet}}).$
Therefore, all inequalities are equalities. In particular $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})=v_{0}({\mathfrak{a}_{\bullet}})/v_{0}({\mathfrak{b}_{\bullet}})$ and $\operatorname{\mathrm{lct}}({\mathfrak{b}_{\bullet}})=A(v_{0})/v_{0}({\mathfrak{b}_{\bullet}})$. Thus, $v_{0}$ satisfies (a)–(c). ∎
8.2. Rigidity II: the monomial case
Focusing on one ideal, we now give a strong and explicit rigidity statement in the monomial case. Assume $R=\mathbb{C}[x_{1},\dots,x_{d}]_{(x_{1},\dots,x_{d})}$ and $I$ is an $\mathfrak{m}$–primary monomial ideal.
Lemma 8.2.
Let $I\subseteq R$ be an $\mathfrak{m}$–primary monomial ideal, and let $\operatorname{NP}(I)\subseteq\mathbb{R}_{\geq 0}^{d}$ denote its Newton polyhedron. Let $\mathcal{U}(I)$ be the finite set of primitive inward normal vectors $u\in\mathbb{N}^{d}$ to the compact facets of $\operatorname{NP}(I)$, and for each such $u$ let $v_{u}$ denote the associated monomial divisorial valuation. Then
$\mathcal{L}_{\mathfrak{m}}(I)\;=\;\max_{u\in\mathcal{U}(I)}\frac{v_{u}(I)}{v_{u}(\mathfrak{m})}.$
Proof.
By Theorem 5.17, $\mathcal{L}_{\mathfrak{m}}(I)$ is computed by finitely many Rees valuations of $I$.
Since $I$ is monomial, the Rees valuations of $I$ are precisely the monomial valuations obtained from the (non-redundant) bounding hyperplanes of the Newton polyhedron $\operatorname{NP}(I)$; see [30, Theorem 10.3.5]. For an $\mathfrak{m}$–primary monomial ideal, the relevant bounding hyperplanes are exactly those defining the compact facets of $\operatorname{NP}(I)$, hence they are indexed by the primitive inward normal vectors $u\in\mathcal{U}(I)$, and the corresponding valuations are the $v_{u}$. Therefore
$\mathcal{L}_{\mathfrak{m}}(I)=\max_{u\in\mathcal{U}(I)}\frac{v_{u}(I)}{v_{u}(\mathfrak{m})}.\qed$
Theorem 8.3 (Monomial $\Theta(I)\geq 1$ and equality classification).
Let $I\subseteq\mathbb{C}[x_{1},\dots,x_{d}]_{\mathfrak{m}}$ be an $\mathfrak{m}$–primary monomial ideal. Then
$\Theta(I)=\frac{1}{d}\,\operatorname{\mathrm{lct}}(I)\cdot\mathcal{L}_{\mathfrak{m}}(I)\ \geq\ 1.$
Moreover, $\Theta(I)=1$ if and only if there exists a vector $u=(t,\dots,t)\in\mathbb{Z}_{>0}^{d}$ such that the same monomial (toric) valuation $v_{u}$ computes both $\operatorname{\mathrm{lct}}(I)$ and $\mathcal{L}_{\mathfrak{m}}(I)$. Equivalently, $\operatorname{NP}(I)$ has a unique compact facet meeting the diagonal ray $\mathbb{R}_{\geq 0}(1,\dots,1)$, and its inward normal is proportional to $(1,\dots,1)$.
Proof.
Step 1: Monomial formulas. For a monomial valuation $v_{u}$ with $u\in\mathbb{Z}_{>0}^{d}$ we have
$v_{u}(x^{\alpha})={\langle u,\alpha\rangle},\quad v_{u}(\mathfrak{m})=\min_{i}u_{i},\quad A(v_{u})=\sum_{i=1}^{d}u_{i}.$
Also $v_{u}(I)=h_{I}(u)$ where $h_{I}$ is the support function of $\operatorname{NP}(I)$.
Howald’s formula ([29]) for monomial ideals gives
$\operatorname{\mathrm{lct}}(I)=\min_{u\in\mathbb{Z}_{>0}^{d}}\frac{A(v_{u})}{v_{u}(I)}=\min_{u\in\mathbb{Z}_{>0}^{d}}\frac{\sum_{i}u_{i}}{h_{I}(u)}.$
Moreover, by Lemma 8.2, we have
(8.1) $\mathcal{L}_{\mathfrak{m}}(I)=\max_{u\in\mathcal{U}}\frac{v_{u}(I)}{v_{u}(\mathfrak{m})}=\max_{u\in\mathcal{U}}\frac{h_{I}(u)}{\min_{i}u_{i}},$
where $\mathcal{U}\subseteq\mathbb{Z}_{>0}^{d}$ is the finite set of primitive inward normals to the compact facets of $\operatorname{NP}(I)$.
Step 2: Inequality and equality. The inequality $\Theta(I)\geq 1$ follows from Theorem 7.1.
If $\Theta(I)=1$, then $\operatorname{\mathrm{lct}}(I)\cdot\mathcal{L}_{\mathfrak{m}}(I)=d=\operatorname{\mathrm{lct}}(\mathfrak{m})$. By Theorem 8.1, equality forces a common valuation computing both $\operatorname{\mathrm{lct}}(I)$ and $\mathcal{L}_{\mathfrak{m}}(I)$ and also satisfying $A(v)/v(\mathfrak{m})=d$. In the monomial category, such a valuation must be $v_{u}$ for some $u\in\mathbb{Z}_{>0}^{d}$, and then
$\frac{A(v_{u})}{v_{u}(\mathfrak{m})}=\frac{\sum_{i}u_{i}}{\min_{i}u_{i}}=d\quad\Longleftrightarrow\quad u_{1}=\cdots=u_{d}.$
Hence, the common extremal valuation must be $v_{u}$ with $u$ proportional to $(1,\dots,1)$. Conversely, if there exists such a $u$ computing both invariants, then the computation in Step 2 shows $\Theta(I)=1$. The polyhedral restatement in terms of a compact facet meeting the diagonal ray $\mathbb{R}_{\geq 0}(1,\dots,1)$ is the standard interpretation of the valuation $v_{(1,\dots,1)}$. ∎
8.3. Rigidity III: extension to Newton nondegenerate settings
We remark that the monomial rigidity picture can be extended to Newton nondegenerate settings via the reduction to the monomial model. (In Section 11, particularly Definition 11.4, we will discuss the notion of Newton nondegeneracy in more details.) The rigidity results above extend to any class of ideals where:
  1. (a)
    both $\operatorname{\mathrm{lct}}(I)$ and $\mathcal{L}_{\mathfrak{m}}(I)$ depend only on $\overline{I}$; and
  2. (b)
    $\overline{I}$ is controlled by an explicit monomial model (same Newton polyhedral data).
This includes Newton nondegenerate ideals and toric models developed in later sections.
Remark 8.4 (Integral-closure reduction preserves $\Theta$).
If $\overline{I}=\overline{J}$, then $\Theta(I)=\Theta(J)$.
Proof.
Since $\overline{I}=\overline{J}$, $\operatorname{\mathrm{lct}}(I)=\operatorname{\mathrm{lct}}(J)$. Also $\mathcal{L}_{\mathfrak{m}}(I)=\mathcal{L}_{\mathfrak{m}}(J)$ since $\mathcal{L}_{\mathfrak{m}}(I)$ is defined via integral-closure containments and hence depends only on $\overline{I}$. Therefore, $\Theta(I)=\Theta(J)$. ∎
Remark 8.5.
Remark 8.4 is the key mechanism for transporting rigidity from the monomial/toric setting to more general singularity classes. For instance, in Newton nondegenerate situations where one can replace an ideal $I$ by a monomial ideal $I_{\mathrm{mon}}$ with $\overline{I}=\overline{I_{\mathrm{mon}}}$, the full inequality and equality classification of Theorem 8.3 applies to $I$ via Remark 8.4.
8.4. Rigidity IV: equality in Teissier–Rees–Sharp inequality
We record a classical but conceptually important source of rigidity: equality in Teissier’s Minkowski-type inequalities for multiplicities forces the ideals involved to be proportional in the Rees–valuative sense.
Lemma 8.6.
Let $I,J\subseteq R$ be nonzero ideals and assume $I+J$ is $\mathfrak{m}$–primary. If $\overline{I^{p}}=\overline{J^{q}}$ for some $p,q\in\mathbb{Z}_{>0}$, then for every Rees valuation $v$ of $IJ$ one has
$\frac{v(I)}{v(J)}=\frac{q}{p}.$
In particular, the ratio $v(I)/v(J)$ is constant on the finite set of Rees valuations of $IJ$.
Proof.
By the valuative criterion for integral closure, $\overline{I^{p}}=\overline{J^{q}}$ is equivalent to pv(I)=qv(J)p\,v(I)=q\,v(J) for every divisorial valuation $v$ centered at $\mathfrak{m}$. Applying this to the (finitely many) Rees valuations of $IJ$ yields the claim. ∎
Proposition 8.7 (Teissier–Rees–Sharp equality $\Rightarrow$ proportionality).
Let $(R,\mathfrak{m})$ be a quasi–unmixed Noetherian local ring of dimension $d$, and let $I,J\subseteq R$ be $\mathfrak{m}$–primary ideals. Assume that equality holds in Teissier’s Minkowski (or first mixed multiplicity) inequality, i.e.,
$e\!\left(I^{[d-1]},J^{[1]}\right)^{d}\;=\;e(I)^{d-1}\,e(J).$
Then, there exist $p,q\in\mathbb{Z}_{>0}$ such that
$\overline{I^{p}}=\overline{J^{q}}.$
Consequently, the ratio $v(I)/v(J)$ is constant (and equal to $q/p$) on the finite set of Rees valuations of $IJ$.
Proof.
This is the classical characterization of the equality case in Teissier’s Minkowski-type inequalities for multiplicities; see Teissier [50] and the detailed local-algebraic treatment of Rees–Sharp [46]. ∎
Corollary 8.8 (Rigidity of extremal ratios in the equality case).
In the setting of Proposition 8.7, suppose moreover that one is considering any valuative extremal ratio of the form
$\mathcal{L}_{J}(I):=\ \sup_{v}\frac{v(I)}{v(J)},$
where $v$ ranges over divisorial valuations centered at $\mathfrak{m}$ (or any subclass that contains the Rees valuations of $IJ$). Then, $\mathcal{L}_{J}(I)=q/p$, and every Rees valuation of $IJ$ computes $\mathcal{L}_{J}(I)$.
Proof.
By Proposition 8.7 we have $\overline{I^{p}}=\overline{J^{q}}$ for some $p,q>0$, hence pv(I)=qv(J)p\,v(I)=q\,v(J) for all divisorial $v$, so $v(I)/v(J)=q/p$ identically. The conclusion follows, and in particular every Rees valuation computes the supremum. ∎
9. Functoriality, stability, and stratification
This section develops the structural consequences of the finite-max principle in families. Assuming $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$ is computed by a finite candidate set of divisorial valuations on a fixed birational model, we study how these candidate sets and the corresponding maximizers behave under basic operations (localization, completion, integral closure, and natural local homomorphisms), establishing functoriality and upper-semicontinuity properties that make the maximizer meaningful geometric data. We then show that, under a uniform finite candidate set hypothesis, parameter spaces admit a finite stratification (equivalently, a wall–chamber decomposition) on which the identity of the computing valuation is constant and $\mathcal{L}_{{\mathfrak{b}_{\bullet}}(t)}({\mathfrak{a}_{\bullet}}(t))$ is given by a single valuative ratio on each stratum (Theorem 9.9). Finally, for one-parameter valuation-concave paths we prove that $t\mapsto 1/\mathcal{L}_{{\mathfrak{b}_{\bullet}}(t)}({\mathfrak{a}_{\bullet}}(t))$ is piecewise fractional linear and that affineness forces persistence of the controlling valuation data (Theorem 9.14).
Throughout this section, $R$ is a Noetherian local ring with maximal ideal $\mathfrak{m}$.
9.1. Finite candidate sets and the maximizer map
A major source of rigidity is finite attainment on a fixed birational model. We isolate a hypothesis that holds in many situations of interest, including: Noetherian families (finitely generated Rees algebra), toric/monomial polyhedral families (see also Section 10), Newton nondegenerate integral-closure reductions, and families arising from a fixed log smooth model.
Definition 9.1 (Finite candidate set hypothesis).
Let $\mathcal{V}=\{v_{1},\dots,v_{N}\}$ be a finite set of divisorial valuations centered at $\mathfrak{m}$. We say that a pair $({\mathfrak{a}_{\bullet}},{\mathfrak{b}_{\bullet}})$ satisfies the finite candidate set hypothesis with respect to $\mathcal{V}$ if
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})=\max_{1\leq i\leq N}\frac{v_{i}({\mathfrak{a}_{\bullet}})}{v_{i}({\mathfrak{b}_{\bullet}})}.$
Remark 9.2.
When $\mathfrak{a}_{\bullet}$} is $\mathfrak{m}$–primary and Noetherian (or sufficiently polyhedral), Section 5 gives that $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$ is computed by finitely many valuations coming from a single normalized blowup. This is exactly the situation of Definition 9.1.
9.2. Functoriality
We record functoriality properties under localization, completion and integral-closure.
Proposition 9.3 (Monotonicity under maps).
Let $\varphi:R\to S$ be a local homomorphism of Noetherian local rings with maximal ideals $\mathfrak{m}$ and $\mathfrak{n}$, respectively, such that $\varphi(\mathfrak{m})\subseteq\mathfrak{n}$. Let ${\mathfrak{a}_{\bullet}},{\mathfrak{b}_{\bullet}}$ be filtrations of $\mathfrak{m}$–primary ideals in $R$ and set
$\varphi({\mathfrak{a}_{\bullet}}):=\{\mathfrak{a}_{p}S\}_{p\geq 1},\quad\varphi({\mathfrak{b}_{\bullet}}):=\{\mathfrak{b}_{q}S\}_{q\geq 1}.$
Then,
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\ \geq\ \mathcal{L}_{\varphi({\mathfrak{b}_{\bullet}})}(\varphi({\mathfrak{a}_{\bullet}})).$
Moreover, if $\phi$ is faithfully flat and has the integral-closure contraction property, i.e., $\overline{IS}\cap R=\overline{I}$ for any ideal $I\subseteq R$, then
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\ =\ \mathcal{L}_{\varphi({\mathfrak{b}_{\bullet}})}(\varphi({\mathfrak{a}_{\bullet}})).$
Proof.
Fix $\varepsilon>0$. By definition of $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$ as an infimum of admissible rational slopes, there exist integers $p,q\geq 1$ such that
$\frac{q}{p}<\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})+\varepsilon\quad\text{and}\quad\mathfrak{b}_{qt}\subseteq\overline{\mathfrak{a}_{pt}}\ \forall\ t\gg 1.$
Applying $\varphi$ gives $\mathfrak{b}_{qt}S\subseteq(\overline{\mathfrak{a}_{pt}})S.$ Since integral dependence is preserved under ring homomorphisms, we obtain $\mathfrak{b}_{qt}S\subseteq\overline{\mathfrak{a}_{pt}S}\ \forall\ t\gg 1.$ Thus, the rational number $\frac{q}{p}$ is an admissible slope for the pair of filtrations $\varphi({\mathfrak{a}_{\bullet}})$ and $\varphi({\mathfrak{b}_{\bullet}})$. By definition of $\mathcal{L}_{\varphi({\mathfrak{b}_{\bullet}})}(\varphi({\mathfrak{a}_{\bullet}}))$, it follows that
$\mathcal{L}_{\varphi({\mathfrak{b}_{\bullet}})}(\varphi({\mathfrak{a}_{\bullet}}))\leq\frac{q}{p}<\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})+\varepsilon.$
Letting $\varepsilon\to 0$ yields $\mathcal{L}_{\varphi({\mathfrak{b}_{\bullet}})}(\varphi({\mathfrak{a}_{\bullet}}))\leq\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$ as desired.
Suppose, in addition, that $\phi$ is faithfully flat and has the integral-closure contraction property. If $\mathfrak{b}_{qt}S\subseteq\overline{\mathfrak{a}_{pt}S}$ for $t\gg 1$, then
$\mathfrak{b}_{qt}\subseteq(\mathfrak{b}_{qt}S)\cap R\subseteq\overline{\mathfrak{a}_{pt}S}\cap R=\overline{\mathfrak{a}_{pt}}\ \forall\ t\gg 1.$
That is, every admissible slope $\frac{q}{p}$ for the filtrations $\phi({\mathfrak{a}_{\bullet}})$ and $\phi({\mathfrak{b}_{\bullet}})$ is also an admissible slope for the filtrations $\mathfrak{a}_{\bullet}$} and $\mathfrak{b}_{\bullet}$}. Therefore, by a similar argument, this implies that $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\leq\mathcal{L}_{\phi({\mathfrak{b}_{\bullet}})}(\phi({\mathfrak{a}_{\bullet}}))$. Hence, we obtained the equality as claimed. ∎
Corollary 9.4 (Localization invariance for $\mathfrak{m}$–primary data).
Let $(R,\mathfrak{m})$ be a Noetherian local ring and let $\mathfrak{p}\subseteq R$ be a prime ideal. Set $R_{\mathfrak{p}}$ with maximal ideal $\mathfrak{p}R_{\mathfrak{p}}$. Let ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ and ${\mathfrak{b}_{\bullet}}=\{\mathfrak{b}_{q}\}_{q\geq 1}$ be filtrations of $\mathfrak{p}$–primary ideals in $R$. Then
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})=\mathcal{L}_{{\mathfrak{b}_{\bullet}}R_{\mathfrak{p}}}({\mathfrak{a}_{\bullet}}R_{\mathfrak{p}}).$
Proof.
This is a direct consequence of Proposition 9.3, noting that $R\rightarrow R_{\mathfrak{p}}$ is faithfully flat and has the integral-closure contraction property. ∎
Corollary 9.5 (Completion invariance).
Let $(R,\mathfrak{m})$ be an excellent Noetherian local ring and let $\widehat{R}$ be the $\mathfrak{m}$–adic completion. Assume that $R$ is analytically unramified (equivalently, $\widehat{R}$ is reduced). Let ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ and ${\mathfrak{b}_{\bullet}}=\{\mathfrak{b}_{q}\}_{q\geq 1}$ be filtrations of $\mathfrak{m}$–primary ideals in $R$. Then,
[…]
Proof.
Observe that $R\rightarrow\widehat{R}$ is faithfully flat. Since $R$ is excellent, $R\rightarrow\widehat{R}$ also has the integral-closure contraction property. Thus, the assertion is a consequence of Proposition 9.3. ∎
Proposition 9.6 (Integral-closure invariance for filtrations).
Let ${\mathfrak{a}_{\bullet}},{\mathfrak{b}_{\bullet}}$ be filtrations in $R$. Define $\overline{{\mathfrak{a}_{\bullet}}}:=\{\overline{\mathfrak{a}_{p}}\}_{p\geq 1}$ and similarly $\overline{{\mathfrak{b}_{\bullet}}}$. Then
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})=\mathcal{L}_{\overline{{\mathfrak{b}_{\bullet}}}}(\overline{{\mathfrak{a}_{\bullet}}}).$
Proof.
For every $p,q$ we have $\overline{\mathfrak{b}_{q}}\subseteq\overline{\mathfrak{a}_{p}}$ if and only if $\mathfrak{b}_{q}\subseteq\overline{\mathfrak{a}_{p}}$, since $\mathfrak{b}_{q}\subseteq\overline{\mathfrak{b}_{q}}$ and integral closure is idempotent. Hence, the set of admissible ratios $q/p$ in the containment definition is unchanged, and $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})=L_{\overline{{\mathfrak{b}_{\bullet}}}}(\overline{{\mathfrak{a}_{\bullet}}})$. ∎
Remark 9.7 (Saturation).
For non-$\mathfrak{m}$–primary ideals one may consider saturations $\mathfrak{a}_{p}^{\mathrm{sat}}$. In the present paper, our main invariants are defined for $\mathfrak{m}$–primary filtrations, for which saturation is trivial. When extending beyond $\mathfrak{m}$–primary data, one must track the change in the center of valuations; the valuative formula remains valid but the candidate set of valuations changes.
9.3. Stratification by the computing valuation
We describe the induced stratification by the computing valuation. Assume the finite candidate set hypothesis (Definition 9.1) holds uniformly for a family $\{({\mathfrak{a}_{\bullet}}(t),{\mathfrak{b}_{\bullet}}(t))\}_{t\in T}$ with the same finite set of divisorial valuations $V=\{v_{1},\dots,v_{N}\}$.
Definition 9.8 (Maximizer set for families).
For each $t\in T$ define
$\alpha_{i}(t):=v_{i}(\mathfrak{a}_{\bullet}(t)),\quad\beta_{i}(t):=v_{i}(\mathfrak{b}_{\bullet}(t))\quad(1\leq i\leq N),$
and assume $\alpha_{i}(t)\in[0,\infty)$ and $\beta_{i}(t)\in(0,\infty)$ for all $i$. (For instance, this holds if ${\mathfrak{b}_{\bullet}}(t)$ is linearly bounded from below in the sense of Remark 5.10.) Define the maximizer set
$\operatorname{Max}({\mathfrak{a}_{\bullet}}(t),{\mathfrak{b}_{\bullet}}(t)):=\left\{\,i\in\{1,\dots,N\}\;:\;\frac{\alpha_{i}(t)}{\beta_{i}(t)}=\max_{1\leq k\leq N}\frac{\alpha_{k}(t)}{\beta_{k}(t)}\right\}.$
If $\operatorname{Max}({\mathfrak{a}_{\bullet}}(t),{\mathfrak{b}_{\bullet}}(t))=\{j\}$ is a singleton, we write $\operatorname{argmax}({\mathfrak{a}_{\bullet}}(t),{\mathfrak{b}_{\bullet}}(t)):=j$.
The stratification phenomenon in Theorem 9.9 below is driven by the finite–max principle established earlier. Indeed, under the standing hypotheses, the value $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$ is computed by a valuation belonging to a fixed finite set (depending only on $\mathfrak{a}_{\bullet}$}), so changes in the value of $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$ can occur only when the identity of the maximizing valuation changes.
Theorem 9.9 (Finite wall–chamber decomposition).
Assume the setup of Definition 9.8. Assume that for each $t\in T$ the finite candidate set hypothesis gives
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}(t)}({\mathfrak{a}_{\bullet}}(t))=\max_{1\leq i\leq N}\frac{\alpha_{i}(t)}{\beta_{i}(t)},\quad\alpha_{i}(t):=v_{i}({\mathfrak{a}_{\bullet}}(t)),\ \beta_{i}(t):=v_{i}({\mathfrak{b}_{\bullet}}(t)).$
Assume moreover that each $\alpha_{i}$ and $\beta_{i}$ is continuous on $T$. Then, $T$ admits a finite partition into locally closed strata, on each of which the maximizer set $\operatorname{Max}({\mathfrak{a}_{\bullet}}(t),{\mathfrak{b}_{\bullet}}(t))$ is constant. In particular, on each stratum, the function $t\mapsto\mathcal{L}_{{\mathfrak{b}_{\bullet}}(t)}({\mathfrak{a}_{\bullet}}(t))$ is given by a single ratio $\alpha_{j}(t)/\beta_{j}(t)$.
Moreover, for each $j\in\{1,\dots,N\}$ the locus
$T_{j}^{\circ}:=\{\,t\in T:\operatorname{Max}({\mathfrak{a}_{\bullet}}(t),{\mathfrak{b}_{\bullet}}(t))=\{j\}\,\}$
is open in $T$.
Proof.
For each nonempty $J\subseteq\{1,\dots,N\}$ set
$T_{J}:=\Big(\bigcap_{i,i^{\prime}\in J}\{\,t:\alpha_{i}(t)\beta_{i^{\prime}}(t)=\alpha_{i^{\prime}}(t)\beta_{i}(t)\,\}\Big)\cap\Big(\bigcap_{i\in J,\ k\notin J}\{\,t:\alpha_{i}(t)\beta_{k}(t)>\alpha_{k}(t)\beta_{i}(t)\,\}\Big).$
By continuity, each equality locus is closed and each strict-inequality locus is open. Thus, each $T_{J}$ is locally closed.
Fix $t\in T$ and write $\rho_{i}(t):=\alpha_{i}(t)/\beta_{i}(t)$, which is well-defined since $\beta_{i}(t)\in(0,\infty)$ for all $i$ by Definition 9.8. Let $J:=\operatorname{Max}({\mathfrak{a}_{\bullet}}(t),{\mathfrak{b}_{\bullet}}(t))$. Then, $\rho_{i}(t)=\rho_{i^{\prime}}(t)$ for $i,i^{\prime}\in J$ and $\rho_{i}(t)>\rho_{k}(t)$ for $i\in J$, $k\notin J$. Multiplying by the positive denominators yields precisely the defining conditions of $T_{J}$. Therefore, $t\in T_{J}$.
Conversely, if $t\in T_{J}$, the defining equalities/inequalities imply that the maximum of the finite set $\{\rho_{1}(t),\dots,\rho_{N}(t)\}$ is attained exactly on $J$, i.e. $\operatorname{Max}({\mathfrak{a}_{\bullet}}(t),{\mathfrak{b}_{\bullet}}(t))=J$. Hence, the sets $T_{J}$ form a finite partition of $T$ by maximizer set, proving the first assertion.
Finally,
$T_{j}^{\circ}=\bigcap_{i\neq j}\{\,t:\alpha_{j}(t)\beta_{i}(t)>\alpha_{i}(t)\beta_{j}(t)\,\},$
an intersection of open sets, and so is open as claimed. ∎
Corollary 9.10 (Stability from a strict gap).
Assume the hypotheses of Theorem 9.9. Fix $t_{0}\in T$ and an index $j\in\{1,\dots,N\}$. If there exists $\delta>0$ such that
(9.1) $\frac{\alpha_{j}(t_{0})}{\beta_{j}(t_{0})}\geq\frac{\alpha_{i}(t_{0})}{\beta_{i}(t_{0})}+\delta\quad\text{for all }i\neq j,$
then there exists an open neighborhood $U$ of $t_{0}$ such that for all $t\in U$ one has
$\operatorname{Max}({\mathfrak{a}_{\bullet}}(t),{\mathfrak{b}_{\bullet}}(t))=\{j\},\text{ and so }\mathcal{L}_{{\mathfrak{b}_{\bullet}}(t)}({\mathfrak{a}_{\bullet}}(t))=\frac{\alpha_{j}(t)}{\beta_{j}(t)}.$
Equivalently, $t_{0}\in T_{j}^{\circ}$ and $T_{j}^{\circ}$ is a neighborhood of $t_{0}$.
Proof.
Since $\beta_{i}(t)\in(0,\infty)$ for all $i$ and $t$ (Definition 9.8), the functions $t\mapsto\alpha_{i}(t)/\beta_{i}(t)$ are well-defined. For each $i\neq j$, set
$g_{i}(t):=\frac{\alpha_{j}(t)}{\beta_{j}(t)}-\frac{\alpha_{i}(t)}{\beta_{i}(t)}.$
By the assumed continuity of $\alpha_{\bullet},\beta_{\bullet}$ and positivity of $\beta_{\bullet}$, each $g_{i}$ is continuous on $T$. The strict gap hypothesis (9.1) gives $g_{i}(t_{0})\geq\delta>0$ for all $i\neq j$. Hence, for each $i\neq j$ there exists an open neighborhood $U_{i}$ of $t_{0}$ such that $g_{i}(t)>0$ for all $t\in U_{i}$. Let $U:=\bigcap_{i\neq j}U_{i}$. Then, for all $t\in U$,
$\frac{\alpha_{j}(t)}{\beta_{j}(t)}>\frac{\alpha_{i}(t)}{\beta_{i}(t)}\quad\text{for all }i\neq j,$
so the unique maximizer is $j$, i.e. $\operatorname{Max}({\mathfrak{a}_{\bullet}}(t),{\mathfrak{b}_{\bullet}}(t))=\{j\}$. The displayed formula for $\mathcal{L}_{{\mathfrak{b}_{\bullet}}(t)}({\mathfrak{a}_{\bullet}}(t))$ follows immediately from the finite-candidate identity of $\mathcal{L}_{{\mathfrak{b}_{\bullet}}(t)}({\mathfrak{a}_{\bullet}}(t))$ given in Theorem 9.9. ∎
Example 9.11 (Change of maximizing valuation).
Let $R=\mathbb{C}[x,y]_{(x,y)}$ and consider the monomial ideal
$\mathfrak{a}=(x^{4},\ x^{2}y^{3},\ y^{5})\subseteq R.$
The Newton polygon $\operatorname{NP}(\mathfrak{a})$ has two compact edges: the segment joining $(4,0)$ to $(2,3)$ with primitive inward normal $u_{1}=(3,2)$, and the segment joining $(2,3)$ to $(0,5)$ with primitive inward normal $u_{2}=(1,1)$. Accordingly,
$v_{u_{1}}(\mathfrak{a})=12,\quad v_{u_{2}}(\mathfrak{a})=5.$
For $t\in[0,1]$, define a family of principal ideals
$\mathfrak{b}_{p}(t):=(x^{\lceil(1+t)p\rceil}y^{\lceil 2p\rceil}),\quad{\mathfrak{b}_{\bullet}}(t):=\{\mathfrak{b}_{p}(t)\}_{p\geq 1}.$
Then, asymptotically,
$v_{u_{1}}({\mathfrak{b}_{\bullet}}(t))=7+3t,\quad v_{u_{2}}({\mathfrak{b}_{\bullet}}(t))=3+t.$
Hence,
$\frac{v_{u_{1}}(\mathfrak{a})}{v_{u_{1}}({\mathfrak{b}_{\bullet}}(t))}=\frac{12}{7+3t},\quad\frac{v_{u_{2}}(\mathfrak{a})}{v_{u_{2}}({\mathfrak{b}_{\bullet}}(t))}=\frac{5}{3+t}.$
These two ratios are equal precisely at $t=\frac{1}{3}$. Consequently,
$\operatorname{argmax}(\mathfrak{a},{\mathfrak{b}_{\bullet}}(t))=\begin{cases}\{u_{1}\},&t<\frac{1}{3},\\[4.0pt] \{u_{1},u_{2}\},&t=\frac{1}{3},\\[4.0pt] \{u_{2}\},&t>\frac{1}{3}.\end{cases}$
In particular, the finite candidate set is fixed, but the maximizing valuation jumps as $t$ crosses the wall $t=\frac{1}{3}$, illustrating the stratification phenomenon of Theorem 9.9.
This example shows that even under finite testing hypotheses, the Łojasiewicz exponent need not be globally linear in families, and rigidity corresponds precisely to the absence of such changes in the maximizing valuation.
Before continue to discuss stability and rigidity consequences, we record a general semicontinuity property of divisorial values in families.
Proposition 9.12 (Upper semicontinuity of divisorial order in algebraic families).
Let $T$ be a Noetherian scheme, let $X$ be a Noetherian integral scheme, and set $X_{T}:=X\times T$. Let $\pi:Y\to X_{T}$ be a proper birational morphism with $Y$ normal, let $E\subseteq Y$ be a prime divisor whose image dominates $T$, and let $\mathcal{I}\subseteq\mathcal{O}_{X_{T}}$ be a coherent ideal sheaf, and for $t\in T$ let $\mathcal{I}_{t}$ be its fiber on $X\times\{t\}$. Define
$\alpha(t):=\operatorname{\mathrm{ord}}_{E}(\mathcal{I}_{t})\in\mathbb{Z}_{\geq 0}\cup\{\infty\}.$
Then, for each $m\geq 0$, the locus
$T_{\geq m}:=\{\,t\in T:\alpha(t)\geq m\,\}$
is Zariski closed in $T$. Equivalently, $t\mapsto\alpha(t)$ is upper semicontinuous on $T$.
Proof.
Set $\mathcal{J}:=\pi^{-1}\mathcal{I}\cdot\mathcal{O}_{Y}$, a coherent ideal sheaf on $Y$, and let $p:E\to T$ be the induced morphism. Since $\pi$ is proper, $p$ is proper.
Fix $m\geq 0$. Consider the coherent $\mathcal{O}_{Y}$–module
$\mathcal{Q}_{m}:=\frac{\mathcal{J}+\mathcal{O}_{Y}(-mE)}{\mathcal{O}_{Y}(-mE)}.$
Let $\mathcal{G}_{m}:=\mathcal{Q}_{m}|_{E}$. Then, $\operatorname{\mathrm{Supp}}(\mathcal{G}_{m})$ is closed in $E$, hence $p(\operatorname{\mathrm{Supp}}(\mathcal{G}_{m}))$ is closed in $T$. We claim that
$T_{\geq m}=T\setminus p(\operatorname{\mathrm{Supp}}(\mathcal{G}_{m})).$
Indeed, let $t\in T$ and let $\eta_{t}$ be the generic point of the prime divisor $E_{t}:=E\times_{T}\kappa(t)$. Since $Y$ is normal, the local ring $\mathcal{O}_{Y_{t},\eta_{t}}$ is a DVR with valuation $\operatorname{\mathrm{ord}}_{E_{t}}$. By definition of $\mathcal{Q}_{m}$, the stalk $(\mathcal{Q}_{m})_{\eta_{t}}$ vanishes if and only if
$\mathcal{J}_{\eta_{t}}\subseteq\mathcal{O}_{Y,\eta_{t}}(-mE),$
and after restricting to the fiber this is equivalent to
$(\mathcal{J}_{t})_{\eta_{t}}\subseteq\mathcal{O}_{Y_{t},\eta_{t}}(-mE_{t})\quad\Longleftrightarrow\quad\operatorname{\mathrm{ord}}_{E_{t}}(\mathcal{J}_{t})\geq m.$
Because $\mathcal{J}=\pi^{-1}\mathcal{I}\cdot\mathcal{O}_{Y}$, we have $\operatorname{\mathrm{ord}}_{E_{t}}(\mathcal{J}_{t})=\operatorname{\mathrm{ord}}_{E_{t}}(\mathcal{I}_{t})=\alpha(t)$. Thus, $(\mathcal{Q}_{m})_{\eta_{t}}=0$ if and only if $\alpha(t)\geq m$, i.e. $\eta_{t}\notin\operatorname{\mathrm{Supp}}(\mathcal{G}_{m})$. Equivalently, $t\notin p(\operatorname{\mathrm{Supp}}(\mathcal{G}_{m}))$ if and only if $\alpha(t)\geq m$, proving the claim. Hence, $T_{\geq m}$ is closed. ∎
9.4. Geodesics of filtrations and fractional linearity of $1/L$
We now turn to one-parameter families of filtrations and prove a fractional linearity statement under a maximizer stability hypothesis. This provides a rigidity mechanism: affineness of $1/L$ characterizes common extremal valuations along the path.
Definition 9.13 (Valuation-concave paths).
Let $\{({\mathfrak{a}_{\bullet}}(t),{\mathfrak{b}_{\bullet}}(t))\}_{t\in[0,1]}$ be a one-parameter family of pairs of filtrations of $\mathfrak{m}$–primary ideals. We say it is valuation-concave (or valuation-affine) if for every divisorial valuation $v$ centered at $\mathfrak{m}$, the functions
[…]
are affine on $[0,1]$.
Assume now that the finite candidate set hypothesis holds uniformly along $t\in[0,1]$ for some finite set $\mathcal{V}=\{v_{1},\dots,v_{N}\}$, and that the family is valuation-concave. Recall that $\alpha_{i}(t):=v_{i}({\mathfrak{a}_{\bullet}}(t))\text{ and }\beta_{i}(t):=v_{i}({\mathfrak{b}_{\bullet}}(t)).$ Then, $\alpha_{i},\beta_{i}$ are affine in $t$. Define
$L(t):=\mathcal{L}_{{\mathfrak{b}_{\bullet}}(t)}({\mathfrak{a}_{\bullet}}(t))=\max_{1\leq i\leq N}\frac{\alpha_{i}(t)}{\beta_{i}(t)}.$
Theorem 9.14 (Piecewise fractional linearity of $1/L$).
Assume that
  1. (1)
    for all $t\in[0,1]$ one has the finite candidate set hypothesis, as in Definition 9.1, with respect to a fixed set $\mathcal{V}=\{v_{1},\dots,v_{N}\}$;
  2. (2)
    the path $t\mapsto({\mathfrak{a}_{\bullet}}(t),{\mathfrak{b}_{\bullet}}(t))$ is valuation-concave, in the sense of Definition 9.13, and $\beta_{i}(t)>0$.
Then, there exists a finite subset $\Sigma\subset[0,1]$ such that, on every connected component $I\subset[0,1]\setminus\Sigma$, one can choose an index $j=j(I)\in\{1,\dots,N\}$ with
$\frac{1}{L(t)}=\frac{\beta_{j}(t)}{\alpha_{j}(t)}\quad\text{for all }t\in I.$
In particular, $t\mapsto 1/L(t)$ is fractional linear (hence $C^{\infty}$) on each such $I$, and its second derivative has constant sign on $I$. Moreover, on a given component $I$ the restriction $t\mapsto 1/L(t)$ is affine if and only if the affine functions $\alpha_{j}$ and $\beta_{j}$ are proportional on $I$.
Proof.
For $1\leq i<j\leq N$ set
$h_{ij}(t):=\alpha_{i}(t)\beta_{j}(t)-\alpha_{j}(t)\beta_{i}(t).$
By valuation-concavity, each $\alpha_{\ell},\beta_{\ell}$ is affine in $t$, hence each $h_{ij}$ is a polynomial of degree at most $2$. Define
$\Sigma:=\{0,1\}\ \cup\ \bigcup_{\begin{subarray}{c}1\leq i<j\leq N\\ h_{ij}\not\equiv 0\end{subarray}}\{\,t\in[0,1]:h_{ij}(t)=0\,\}.$
Then, $\Sigma$ is finite, since each nonzero polynomial $h_{ij}$ has finitely many zeros (or is identically zero, in which case the pair $(i,j)$ does not contribute to “walls” and can be ignored) and there are only finitely many pairs $(i,j)$.
Let $I\subset[0,1]\setminus\Sigma$ be a connected component. Fix $i<j$. If $h_{ij}\equiv 0$, then $\alpha_{i}(t)\beta_{j}(t)=\alpha_{j}(t)\beta_{i}(t)$ for all $t$, hence $\alpha_{i}(t)/\beta_{i}(t)=\alpha_{j}(t)/\beta_{j}(t)$ for all $t$ (using $\beta_{i}(t)\beta_{j}(t)>0$). If $h_{ij}\not\equiv 0$, then by definition of $\Sigma$ we have $h_{ij}(t)\neq 0$ for all $t\in I$, and since $h_{ij}$ is continuous it has constant sign on $I$. Using again $\beta_{i}(t)\beta_{j}(t)>0$, this implies that the comparison
$\frac{\alpha_{i}(t)}{\beta_{i}(t)}\ \gtreqless\ \frac{\alpha_{j}(t)}{\beta_{j}(t)}\quad\Longleftrightarrow\quad h_{ij}(t)\ \gtreqless\ 0$
is independent of $t\in I$. Therefore, the total ordering (with ties) of the finite set of numbers $\{\alpha_{1}(t)/\beta_{1}(t),\dots,\alpha_{N}(t)/\beta_{N}(t)\}$ does not change with $t\in I$. Hence, the maximizer set $\operatorname{Max}({\mathfrak{a}_{\bullet}}(t),{\mathfrak{b}_{\bullet}}(t))$ is constant on $I$. Choose any index $j=j(I)$ in this constant maximizer set. Then, for all $t\in I$,
$L(t)=\max_{1\leq i\leq N}\frac{\alpha_{i}(t)}{\beta_{i}(t)}=\frac{\alpha_{j}(t)}{\beta_{j}(t)}.$
Since ${\mathfrak{a}_{\bullet}}(t)$ consists of $\mathfrak{m}$–primary ideals and $v_{j}$ is centered at $\mathfrak{m}$, we have $\alpha_{j}(t)=v_{j}({\mathfrak{a}_{\bullet}}(t))>0$ on $[0,1]$. Write $\alpha_{j}(t)=a_{0}+a_{1}t\text{ and }\beta_{j}(t)=b_{0}+b_{1}t.$ Then, on $I$,
$\frac{1}{L(t)}=\frac{b_{0}+b_{1}t}{a_{0}+a_{1}t},$
a ratio of affine functions with positive denominator, so fractional linear and $C^{\infty}$ on $I$. Differentiating twice gives
$\frac{d^{2}}{dt^{2}}\left(\frac{b_{0}+b_{1}t}{a_{0}+a_{1}t}\right)=-\frac{2a_{1}(b_{1}a_{0}-b_{0}a_{1})}{(a_{0}+a_{1}t)^{3}},$
whose sign is constant on $I$ because the denominator is positive and the numerator is constant. Finally, $\beta_{j}/\alpha_{j}$ is affine on $I$ if and only if $b_{1}a_{0}-b_{0}a_{1}=0$, i.e. if and only if $\alpha_{j}$ and $\beta_{j}$ are proportional affine functions on $I$. ∎
Remark 9.15 (A cleaner sufficient condition).
If $\alpha_{j}(t)$ is constant in $t$ (or more generally $a_{1}=0$), then $\beta_{j}/\alpha_{j}$ is affine and hence concave. This occurs in many geometric situations where $\mathfrak{a}_{\bullet}$} is fixed and only $\mathfrak{b}_{\bullet}$} varies.
Corollary 9.16 (Geodesic rigidity from global affineness).
Assume the hypotheses of Theorem 9.14. If the function $t\mapsto 1/L(t)$ is affine on $[0,1]$, then there exists an index $j\in\{1,\dots,N\}$ such that
$L(t)=\frac{\alpha_{j}(t)}{\beta_{j}(t)}\quad\text{for all }t\in[0,1].$
Equivalently, $j\in\operatorname{Max}({\mathfrak{a}_{\bullet}}(t),{\mathfrak{b}_{\bullet}}(t))$ for every $t\in[0,1]$.
Proof.
By Theorem 9.14, there exists a finite set $\Sigma\subset[0,1]$ such that on each connected component $I\subset[0,1]\setminus\Sigma$ one can choose an index $j=j(I)$ with
$\frac{1}{L(t)}=\frac{\beta_{j}(t)}{\alpha_{j}(t)}\quad\text{for all }t\in I.$
Fix one such component $I$ and let $j=j(I)$. By assumption, $1/L(t)$ is affine on $[0,1]$, say
$\frac{1}{L(t)}=c_{0}+c_{1}t\quad(t\in[0,1]).$
Thus, on $I$ we have
$\beta_{j}(t)=\bigl(c_{0}+c_{1}t\bigr)\alpha_{j}(t).$
Define $F(t):=\beta_{j}(t)-\bigl(c_{0}+c_{1}t\bigr)\alpha_{j}(t).$ Since $\alpha_{j}$ and $\beta_{j}$ are affine functions of $t$ (valuation–concavity), the function $F(t)$ is a polynomial of degree at most $2$. As $F$ vanishes on the nonempty open interval $I$, it follows that $F\equiv 0$ on $[0,1]$. Hence,
$\frac{\beta_{j}(t)}{\alpha_{j}(t)}=c_{0}+c_{1}t\quad\text{for all }t\in[0,1].$
Since $\alpha_{j}(t)>0$ for all $t$ (because ${\mathfrak{a}_{\bullet}}(t)$ consists of $\mathfrak{m}$–primary ideals and $v_{j}$ is centered at $\mathfrak{m}$), we may invert to obtain
$L(t)=\frac{\alpha_{j}(t)}{\beta_{j}(t)}\quad\text{for all }t\in[0,1].$
Finally, this identity implies $j\in\operatorname{Max}({\mathfrak{a}_{\bullet}}(t),{\mathfrak{b}_{\bullet}}(t))$ for every $t$ by the definition of $\operatorname{Max}$. ∎
Part VI Polyhedral Models and Explicit Computation
In this part, we specialize the general valuative framework to toric and Newton nondegenerate settings, where the Łojasiewicz exponent admits explicit polyhedral descriptions. We will also give a better understanding of Łojasiewicz exponent at infinity.
10. Toric and monomial models: polyhedra, valuations, and formulas
This section provides the explicit polyhedral model underlying finite–max computations in toric and monomial settings. We recall how integral closure of monomial ideals is encoded by Newton polyhedra and how toric (monomial) divisorial valuations correspond to linear functionals/support functions, so that valuative containment thresholds become linear optimization problems (Theorem 10.8). From this perspective we derive Newton/facet formulas for the algebraic Łojasiewicz exponent, showing that $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$ is computed by finitely many toric valuations determined by facet data (Theorems 10.11 and 10.18). In particular, the finite candidate set hypothesis becomes completely concrete: the maximizing valuation comes from a facet normal, yielding uniform finite-max formulas and effective computation (Corollary 10.16).
10.1. Affine semigroup rings and monomial ideals
For basic terminology in toric geometry, we refer to [11, 18]. Let $N\cong\mathbb{Z}^{n}$ be a lattice and let $M:=\operatorname{Hom}_{\mathbb{Z}}(N,\mathbb{Z})$ be the dual lattice. Fix a finitely generated (not necessarily saturated) subsemigroup $S\subseteq M$ such that the rational cone
$\sigma^{\vee}:=\mathbb{R}_{\geq 0}\cdot S\ \subseteq M_{\mathbb{R}}:=M\otimes_{\mathbb{Z}}\mathbb{R}$
is a pointed rational polyhedral cone of full dimension $n$ (equivalently, $\sigma^{\vee}$ contains no nontrivial linear subspace). Let $\sigma\subseteq N_{\mathbb{R}}$ be the dual cone:
$\sigma:=\{u\in N_{\mathbb{R}}:\ {\langle u,x\rangle}\geq 0\ \text{for all }x\in\sigma^{\vee}\}.$
Consider the affine semigroup ring $k[S]$ and its maximal monomial ideal
$\mathfrak{m}_{S}:=(\chi^{m}:m\in S\setminus\{0\}).$
We work in the local ring $R:=k[S]_{\mathfrak{m}_{S}},$ which is an excellent Noetherian local domain, not necessarily normal. This is the local ring of the affine toric variety $X_{\sigma}=\operatorname{Spec}(k[S])$ at its distinguished torus-fixed point.
An ideal $I\subseteq R$ is monomial if it is generated by monomials $\chi^{m}$, $m\in S$. Equivalently, it is the extension to $R$ of a monomial ideal of $k[S]$.
Remark 10.1.
In $R=k[x_{1},\dots,x_{n}]_{(x_{1},\dots,x_{n})}$ one recovers the usual monomial ideals by taking $S=\mathbb{N}^{n}$ and $\chi^{m}=x^{m}$.
10.2. Newton polyhedra in the toric setting
For a monomial ideal $I\subseteq R$, define its exponent set
$E(I):=\{m\in S:\ \chi^{m}\in I\}\subseteq S\subseteq M.$
Definition 10.2 (Newton polyhedron).
The Newton polyhedron of a monomial ideal $I\subseteq R$ is
$\operatorname{NP}(I):=\operatorname{conv}(E(I))+\sigma^{\vee}\ \subseteq M_{\mathbb{R}}.$
Particularly, if $I$ is generated by monomials $\chi^{m_{1}},\dots,\chi^{m_{r}}$, then $\operatorname{NP}(I)=\operatorname{conv}(m_{1},\dots,m_{r})+\sigma^{\vee}$ is a closed rational polyhedron.
Lemma 10.3 (Basic properties).
Let $I,J\subseteq R$ be monomial ideals. Then:
  1. (a)
    $\operatorname{NP}(I)$ is a closed convex set satisfying $\operatorname{NP}(I)+\sigma^{\vee}=\operatorname{NP}(I)$.
  2. (b)
    $\operatorname{NP}(IJ)=\operatorname{NP}(I)+\operatorname{NP}(J)$ (Minkowski sum).
  3. (c)
    $\operatorname{NP}(I^{p})=p\cdot\operatorname{NP}(I)$ for $p\geq 1$.
Proof.
All statements are standard for Newton polyhedra of monomial ideals; see, e.g., [48, 51]. For completeness: (b) follows from $E(IJ)=E(I)+E(J)$ and convexity, and (c) follows by iterating (b). ∎
10.3. Toric (monomial) valuations and support functions
Each vector $u\in\sigma\cap N$ defines a monomial valuation on $k[S]$ by
$v_{u}(\chi^{m}):={\langle u,m\rangle}\quad(m\in S),$
and we extend it to $R=k[S]_{\mathfrak{m}_{S}}$.
Definition 10.4 (Toric valuation).
For $u\in\sigma\cap N$, let $v_{u}$ denote the valuation on $R$ determined by $v_{u}(\chi^{m})={\langle u,m\rangle}$. We call $v_{u}$ a toric valuation. It is centered at $\mathfrak{m}_{S}$ if $u$ lies in the relative interior $\sigma^{\circ}$, equivalently $v_{u}(\mathfrak{m}_{S})>0$.
Definition 10.5 (Support function).
Let $C\subseteq M_{\mathbb{R}}$ be a nonempty closed convex set with $C+\sigma^{\vee}=C$. For $u\in\sigma$ define the support function
$h_{C}(u):=\inf\{{\langle u,x\rangle}:x\in C\}\in\mathbb{R}_{\geq 0}.$
For a monomial ideal $I$ we write $h_{I}:=h_{\operatorname{NP}(I)}$.
Lemma 10.6 (Valuations and support functions).
Let $I\subseteq R$ be a monomial ideal and let $u\in\sigma\cap N$. Then
$v_{u}(I)=h_{I}(u).$
Proof.
Since $I$ is generated by monomials, $v_{u}(I)=\min\{\langle u,m\rangle:m\in E(I)\}$. Taking convex hull and adding $\sigma^{\vee}$ does not change the minimum because $\langle u,\cdot\rangle$ is linear and nonnegative on $\sigma^{\vee}$. Thus, $v_{u}(I)=h_{\operatorname{NP}(I)}(u)$; see, for instance, [11]. ∎
Definition 10.7 (Normalization by a fixed monomial ideal).
Fix an $\mathfrak{m}_{S}$-primary monomial ideal $\mathfrak{q}\subseteq R$. For $u\in\sigma^{\circ}\cap N$, set
$\widehat{u}:=\frac{u}{h_{\mathfrak{q}}(u)}\in\sigma^{\circ}\cap N_{\mathbb{R}},\quad\text{so that}\quad v_{\widehat{u}}(\mathfrak{q})=1.$
(Since $\mathfrak{q}$ is $\mathfrak{m}_{S}$–primary, $h_{\mathfrak{q}}(u)>0$ for all $u\in\sigma^{\circ}\cap N$.)
10.4. Integral closure and Newton polyhedra in toric rings
The next result is standard and is the backbone of toric computations: for monomial ideals, integral closure is detected by the Newton polyhedron.
Theorem 10.8 (Integral closure of monomial ideals).
Let $R=k[S]_{\mathfrak{m}_{S}}$ as above and let $I\subseteq R$ be a monomial ideal. Then, $\overline{I}$ is monomial and satisfies
$\chi^{m}\in\overline{I}\quad\Longleftrightarrow\quad m\in\operatorname{NP}(I)\cap S.$
Equivalently, $\operatorname{NP}(\overline{I})=\operatorname{NP}(I)$.
Proof.
We give a direct valuative proof that is self-contained in our context.
Let $m\in S$. As in the proof of Lemma 5.2, $\chi^{m}\in\overline{I}$ if and only if, for any prime divisor $E$ appearing on the normalized blowup of $I$,
$\operatorname{\mathrm{ord}}_{E}(\chi^{m})\geq\operatorname{\mathrm{ord}}_{E}(I).$
Since $I$ is monomial, the blowup of $I$ is torus-invariant, and its normalization is a normal toric modification; in particular, its exceptional prime divisors $E$ centered at $\mathfrak{m}_{S}$ are torus-invariant and correspond to rays $\rho$ of the associated fan with primitive generators $u_{\rho}\in\sigma^{\circ}\cap N$ (see, for example, [11]). The corresponding divisorial valuation satisfies
$\operatorname{\mathrm{ord}}_{E}(\chi^{m})=\langle u_{\rho},m\rangle\text{ and }\operatorname{\mathrm{ord}}_{E}(I)=v_{u_{\rho}}(I).$
By Lemma 10.6, $v_{u_{\rho}}(I)=h_{I}(u_{\rho})$, the support function of the Newton polyhedron $\operatorname{NP}(I)$. Hence, $\chi^{m}\in\overline{I}$ if and only if
[…]
Finally, $\operatorname{NP}(I)$ is a rational polyhedron, so it is the intersection of its supporting halfspaces defined by the primitive inward normals of its compact faces. Therefore, the above inequalities are equivalent to $m\in\operatorname{NP}(I)\cap S$. ∎
10.5. Toric formula for the algebraic Łojasiewicz exponent
We now specialize our general invariant to monomial ideals (and then to monomial filtrations) in the toric local ring $R=k[S]_{\mathfrak{m}_{S}}$.
Lemma 10.9 (Toric valuative formula for monomial ideals).
Let $\mathfrak{a},\mathfrak{b}\subseteq R$ be monomial ideals with $\mathfrak{a}$ being $\mathfrak{m}_{S}$-primary. Then
$\mathcal{L}_{\mathfrak{b}}(\mathfrak{a})=\max_{u}\frac{v_{u}(\mathfrak{a})}{v_{u}(\mathfrak{b})}=\max_{u}\frac{h_{\operatorname{NP}(\mathfrak{a})}(u)}{h_{\operatorname{NP}(\mathfrak{b})}(u)},$
where the max is taken over finitely many $u\in\sigma^{\circ}\cap N$.
Proof.
By Theorem 5.17, when $\mathfrak{a}$ is $\mathfrak{m}_{S}$-primary, $\mathcal{L}_{\mathfrak{b}}(\mathfrak{a})$ is computed by finitely many Rees valuations of $\mathfrak{a}$. For the monomial ideal $\mathfrak{a}$, the normalized blowup of $\mathfrak{a}$ is toric and its prime divisors are torus-invariant. Hence, its Rees valuations are toric valuations $v_{u}$ for finitely many primitive lattice vectors $u\in\sigma^{\circ}\cap N$, lying on rays of a fan refining $\sigma$. Therefore, the maximum over finitely many Rees valuations equals the maximum over finitely many toric valuations $v_{u}$. Finally, Lemma 10.6 gives $v_{u}(\cdot)=h_{\operatorname{NP}(\cdot)}(u)$ on monomial ideals. ∎
10.6. Facet normals and finite candidate sets
We next package the maximization problem in purely polyhedral terms.
A compact facet of $\operatorname{NP}(\mathfrak{a})$ means a facet not parallel to $\sigma^{\vee}$ (i.e., a facet whose supporting hyperplane intersects $\sigma^{\vee}$ only at the origin). These are precisely facets $F$ whose inward normal vector $u_{F}$ lies in $\sigma^{\circ}\cap N$ and, thus, gives a valuation centered at $\mathfrak{m}_{S}$.
Definition 10.10 (Facet normals).
Let $\mathfrak{a}$ be an $\mathfrak{m}_{S}$-primary monomial ideal. Let $\mathcal{N}(\mathfrak{a})$ denote the finite set of primitive inward normals $u_{F}\in\sigma\cap N$ to the compact facets $F$ of $\operatorname{NP}(\mathfrak{a})$.
Theorem 10.11 (Facet finite–max formula).
Let $\mathfrak{a}\subseteq R$ be an $\mathfrak{m}_{S}$-primary monomial ideal and let $\mathfrak{b}\subseteq R$ be any monomial ideal. Then
$\mathcal{L}_{\mathfrak{b}}(\mathfrak{a})=\max_{u\in\mathcal{N}(\mathfrak{a})}\frac{h_{\operatorname{NP}(\mathfrak{a})}(u)}{h_{\operatorname{NP}(\mathfrak{b})}(u)}.$
In particular, $\mathcal{L}_{\mathfrak{b}}(\mathfrak{a})$ is attained by a toric divisorial valuation $v_{u}$ with $u\in\mathcal{N}(\mathfrak{a})$.
Proof.
By Theorem 5.17, $\mathcal{L}_{\mathfrak{b}}(\mathfrak{a})$ is the maximum of $v(\mathfrak{a})/v(\mathfrak{b})$ over the Rees valuations of $\mathfrak{a}$. For a monomial ideal $\mathfrak{a}$, the Rees valuations correspond exactly to the torus-invariant prime divisors on the normalized blowup of $\mathfrak{a}$. Polyhedrally, these correspond to the compact facets of $\operatorname{NP}(\mathfrak{a})$, and their valuations are the toric valuations $v_{u_{F}}$, where $u_{F}$ is the primitive inward normal to the facet $F$. (Equivalently, $\operatorname{NP}(\mathfrak{a})$ determines a normal fan; the rays corresponding to compact facets yield precisely these divisors.) Thus, the maximum is achieved among the finite set $\{v_{u}:u\in\mathcal{N}(\mathfrak{a})\}$. Finally, apply Lemma 10.6. ∎
Note that, in Theorem 10.11, only compact facets contribute since $\mathfrak{a}$ being $\mathfrak{m}_{S}$–primary implies $h_{\operatorname{NP}(\mathfrak{a})}(u)>0$ precisely along inward normals to compact faces.
10.7. Monomial filtrations in the toric setting
We now extend the preceding discussion from powers of ideals to monomial filtrations.
Definition 10.12 (Monomial filtration and Newton polyhedra).
A filtration ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ in $R$ is monomial if each $\mathfrak{a}_{p}$ is a monomial ideal. For such a family define
$\operatorname{NP}({\mathfrak{a}_{\bullet}}):=\bigcup_{p\geq 1}\frac{1}{p}\,\operatorname{NP}(\mathfrak{a}_{p})\ \subseteq M_{\mathbb{R}}.$
Remark 10.13.
When $\mathfrak{a}_{p}=\overline{\mathfrak{a}^{p}}$, one has $\operatorname{NP}(\mathfrak{a}_{p})=p\,\operatorname{NP}(\mathfrak{a})$ and hence $\operatorname{NP}({\mathfrak{a}_{\bullet}})=\operatorname{NP}(\mathfrak{a})$. For more general families $\operatorname{NP}({\mathfrak{a}_{\bullet}})$ is a convex region encoding asymptotic integral closures.
Lemma 10.14 (Support function for a monomial filtration).
Let $\mathfrak{a}_{\bullet}$} be a monomial graded family and $u\in\sigma^{\circ}\cap N$. Then
$v_{u}({\mathfrak{a}_{\bullet}})=\inf_{p\geq 1}\frac{v_{u}(\mathfrak{a}_{p})}{p}=\inf_{p\geq 1}\frac{h_{\operatorname{NP}(\mathfrak{a}_{p})}(u)}{p}=\inf_{x\in\operatorname{NP}({\mathfrak{a}_{\bullet}})}{\langle u,x\rangle}.$
Proof.
The first equality is Remark 5.1. The second is Lemma 10.6 applied to each $\mathfrak{a}_{p}$. For the last equality, note that $\operatorname{NP}({\mathfrak{a}_{\bullet}})$ is the union of sets $\frac{1}{p}\operatorname{NP}(\mathfrak{a}_{p})$, and
$\inf_{x\in\frac{1}{p}\operatorname{NP}(\mathfrak{a}_{p})}{\langle u,x\rangle}=\frac{1}{p}\inf_{y\in\operatorname{NP}(\mathfrak{a}_{p})}{\langle u,y\rangle}=\frac{h_{\operatorname{NP}(\mathfrak{a}_{p})}(u)}{p}.$
Taking infima over $p$ gives the claim. ∎
Lemma 10.15 (Finite–max on a finite fan).
Let $\sigma\subseteq N_{\mathbb{R}}$ be a strongly convex rational polyhedral cone and set $\sigma^{\circ}$ for its interior. Let $\Sigma$ be a finite rational polyhedral fan supported on $\sigma$ with rays $\rho_{1},\dots,\rho_{M}$ and primitive generators $u_{1},\dots,u_{M}\in N$. Let $\alpha,\beta:\sigma^{\circ}\to\mathbb{R}_{>0}$ be functions such that:
  1. (1)
    $\alpha$ and $\beta$ extend to continuous functions on $\sigma^{\circ}$ and are positive on $\sigma^{\circ}$;
  2. (2)
    for every cone $\tau\in\Sigma$, the restrictions $\alpha|_{\tau^{\circ}}$ and $\beta|_{\tau^{\circ}}$ are linear forms in $u$.
Then,
$\sup_{u\in\sigma^{\circ}\cap N}\frac{\alpha(u)}{\beta(u)}\;=\;\max_{1\leq j\leq M}\frac{\alpha(u_{j})}{\beta(u_{j})}.$
In particular, the supremum is a maximum attained at a primitive ray generator of $\Sigma$.
Proof.
Since $\beta(u)>0$ on $\sigma^{\circ}$, we may normalize on the compact slice
$K:=\{u\in\sigma^{\circ}:\beta(u)=1\}.$
On each cone $\tau\in\Sigma$, the set $K\cap\tau$ is a compact polytope and $\alpha/\beta=\alpha$ on $K$ is linear on $\tau$ by hypothesis. Hence, the maximum of $\alpha$ on $K\cap\tau$ is attained at a vertex of that polytope, which lies on an extreme ray of $\tau$. Taking the maximum over finitely many cones yields that the global supremum on $K$ is attained at one of the finitely many ray generators $u_{1},\dots,u_{M}$ of $\Sigma$, giving the claimed formula. ∎
Corollary 10.16 (Polyhedral Newton data $\Rightarrow$ finite toric candidate set).
Assume that the toric valuation functions
$u\longmapsto v_{u}(\mathfrak{a}_{\bullet}),\quad u\longmapsto v_{u}(\mathfrak{b}_{\bullet})$
are piecewise-linear on a common finite fan $\Sigma$ supported on $\sigma$ (for example, when the relevant Newton data is rational polyhedral and one takes $\Sigma$ a common refinement of the associated normal fans). Then, the toric supremum is a finite maximum:
$\sup_{u\in\sigma^{\circ}\cap N}\frac{v_{u}(\mathfrak{a}_{\bullet})}{v_{u}(\mathfrak{b}_{\bullet})}\;=\;\max_{u\in\mathrm{Ray}(\Sigma)}\frac{v_{u}(\mathfrak{a}_{\bullet})}{v_{u}(\mathfrak{b}_{\bullet})}.$
Proof.
Apply Lemma 10.15 with $\alpha(u)=v_{u}(a_{\bullet})$ and $\beta(u)=v_{u}(b_{\bullet})$. ∎
Remark 10.17.
If $f$ is Newton nondegenerate and $\operatorname{NP}(f)$ is fixed, one typically chooses a regular fan $\Sigma$ refining the normal fan of $\operatorname{NP}(f)$. On such a fixed toric model, the relevant toric valuation data attached to $f$ (and, in the standard setup, to $J(f)$ as well) is determined by linear forms on the cones of $\Sigma$. In that situation Corollary 10.16 gives a finite candidate set that is independent of parameters as long as the Newton polyhedron is unchanged.
Theorem 10.18 (Toric valuative formula for monomial filtrations).
Let $\mathfrak{a}_{\bullet}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ and $\mathfrak{b}_{\bullet}=\{\mathfrak{b}_{q}\}_{q\geq 1}$ be graded families of monomial ideals in $R$.
(1) (Universal toric lower bound) One has
$\mathcal{L}_{\mathfrak{b}_{\bullet}}(\mathfrak{a}_{\bullet})\ \geq\ \sup_{u\in\sigma^{\circ}\cap N}\frac{v_{u}(\mathfrak{a}_{\bullet})}{v_{u}(\mathfrak{b}_{\bullet})}=\sup_{u\in\sigma^{\circ}\cap N}\frac{\inf_{y\in\operatorname{NP}(\mathfrak{a}_{\bullet})}\langle u,y\rangle}{\inf_{x\in\operatorname{NP}(\mathfrak{b}_{\bullet})}\langle u,x\rangle}.$
(2) (Finite–max among toric valuations) Suppose that the functions $u\mapsto v_{u}(\mathfrak{a}_{\bullet})$ and $u\mapsto v_{u}(\mathfrak{b}_{\bullet})$ are piecewise-linear on a common finite rational fan $\Sigma$ supported on $\sigma$. Then
$\sup_{u\in\sigma^{\circ}\cap N}\frac{v_{u}(\mathfrak{a}_{\bullet})}{v_{u}(\mathfrak{b}_{\bullet})}=\max_{u\in\mathrm{Ray}(\Sigma)}\frac{v_{u}(\mathfrak{a}_{\bullet})}{v_{u}(\mathfrak{b}_{\bullet})}.$
(3) (Equality with $\mathcal{L}_{\mathfrak{b}_{\bullet}}(\mathfrak{a}_{\bullet})$ under toric testing) Assume in addition that for all $p,q$ the containment $\mathfrak{b}_{q}\subseteq\overline{\mathfrak{a}_{p}}$ is equivalent to the system of inequalities
$v_{u}(\mathfrak{b}_{q})\geq v_{u}(\mathfrak{a}_{p})\quad\text{for all }u\in\mathrm{Ray}(\Sigma).$
Then,
$\mathcal{L}_{\mathfrak{b}_{\bullet}}(\mathfrak{a}_{\bullet})=\max_{u\in\mathrm{Ray}(\Sigma)}\frac{v_{u}(\mathfrak{a}_{\bullet})}{v_{u}(\mathfrak{b}_{\bullet})}.$
Proof.
(1) By Theorem 5.6 we have the universal bound $\mathcal{L}_{\mathfrak{b}_{\bullet}}(\mathfrak{a}_{\bullet})\geq\sup_{v}\frac{v(\mathfrak{a}_{\bullet})}{v(\mathfrak{b}_{\bullet})}$ over divisorial valuations $v$ centered at $\mathfrak{m}_{S}$. Restricting to toric divisorial valuations $v=v_{u}$ gives the first inequality. The expression in terms of $\operatorname{NP}(\mathfrak{a}_{\bullet})$ and $\operatorname{NP}(\mathfrak{b}_{\bullet})$ follows from Lemma 10.14.
(2) This is exactly Lemma 10.15 applied to the functions $u\mapsto v_{u}(\mathfrak{a}_{\bullet})$ and $u\mapsto v_{u}(\mathfrak{b}_{\bullet})$.
(3) Under the stated testing equivalence, the finite testing hypothesis (H1) of Theorem 5.8 holds with $V=\{v_{u}:u\in\mathrm{Ray}(\Sigma)\}$, hence Theorem 5.8 yields the desired equality. ∎
Remark 10.19 (Mechanism behind finite testing in the toric case).
The finite–max formulas obtained in this section are not a consequence of toric geometry per se, but of the underlying polyhedral structure. In the monomial setting, Newton polyhedra encode integral–closure containments by linear inequalities, toric valuations correspond to linear functionals on these polyhedra, and support functions are piecewise linear with respect to a finite fan. As a result, the valuative optimization problem defining the Lojasiewicz exponent reduces to a finite maximization over extremal directions, namely the primitive inward normals of compact facets. This mechanism provides the prototypical example of the general “finite testing” principle developed earlier in the paper.
Remark 10.20 (Limits of the toric mechanism).
The explicit finite testing principle established here relies crucially on the existence of a polyhedral model controlling all relevant integral–closure containments on a fixed toric modification. Outside the toric or monomial setting, such a uniform polyhedral structure need not exist: although valuative formulas remain valid in great generality, the set of divisorial valuations that compute the Lojasiewicz exponent may depend on the chosen principalization and need not admit an a priori finite reduction. Section 12 investigates to what extent Newton–type hypotheses recover finite testing and explains why additional verifiable assumptions are necessary beyond the toric case.
11. Łojasiewicz exponent at infinity and Newton nondegeneracy
This section focuses on the Łojasiewicz exponent at infinity for polynomial mappings. Particularly, we will consider: (i) a single fixed birational model on which the relevant divisor ideals become invertible, yielding finite-minimum formulas for Łojasiewicz exponent at infinity (Theorems 11.1 and 11.2); and (ii) in Newton nondegenerate settings, a toric model that makes the minimizing divisors explicit, recovering classical Newton-type formulas (Theorem 11.10 and Corollary 11.11).
11.1. Finiteness of the algebraic Łojasiewicz exponent at infinity
Even though, in Definition 4.8 (and Theorem 4.9), the local Łojasiewicz exponent at infinity of a polynomial mapping $F$, $\mathcal{L}^{\text{an}}_{\infty,\xi}(F)$, is defined in the local ring $\mathcal{O}_{\overline{\Gamma}_{F},\xi}$, there exists a uniform description after passing to a single normal proper birational model on which the divisor ideals $\mathscr{I}_{X}$ and $\mathscr{I}_{Y}$ become invertible. On such a model, integral–closure containments are equivalent to finitely many divisorial inequalities, so the local and global exponents at infinity admit a finite minima formulation similar to what we have seen for the exponent at an isolated critical point (Theorem 5.17).
Notations. We keep the notation from Section 4.3. In particular, $\overline{\Gamma}_{F}\subseteq\mathbb{P}^{n}_{\mathbb{C}}\times\mathbb{P}^{m}_{\mathbb{C}}$ denotes the normalization of the Zariski closure of the graph of $F$, with projections $\pi,\rho$ and divisor ideals $\mathscr{I}_{X}=\mathcal{I}_{D_{X}}$ and $\mathscr{I}_{Y}=\mathcal{I}_{D_{Y}}$.
Since $\overline{\Gamma}_{F}$ is projective over $\mathbb{C}$, there exists a proper birational morphism
(11.1) $\mu:Z\longrightarrow\overline{\Gamma}_{F}$
with $Z$ normal such that both $\mathscr{I}_{X}\mathcal{O}_{Z}$ and $\mathscr{I}_{Y}\mathcal{O}_{Z}$ are invertible (for instance, take a principalization of $\mathscr{I}_{X}\cdot\mathscr{I}_{Y}$). Write
(11.2) $\mathscr{I}_{X}\mathcal{O}_{Z}=\mathcal{O}_{Z}(-A)\ \text{ and }\ \mathscr{I}_{Y}\mathcal{O}_{Z}=\mathcal{O}_{Z}(-B),$
where $A$ and $B$ are effective Cartier divisors on $Z$. For any prime divisor $E\subseteq Z$, set
$\operatorname{\mathrm{ord}}_{E}(\mathscr{I}_{X}):=\text{coeff}_{E}(A)\quad\text{and}\quad\operatorname{\mathrm{ord}}_{E}(\mathscr{I}_{Y}):=\text{coeff}_{E}(B).$
Thus, $\operatorname{\mathrm{ord}}_{E}(\mathscr{I}_{X})>0$ if and only if $E\subseteq\operatorname{\mathrm{Supp}}(A)$.
Local exponents on the fixed model. Fix $\xi\in D_{X}$ and, as before, set $R_{\xi}:=\mathcal{O}_{\overline{\Gamma}_{F},\xi}$, $\mathscr{I}_{X,\xi}:=\mathscr{I}_{X}\cdot R_{\xi}$, and $\mathscr{I}_{Y,\xi}:=\mathscr{I}_{Y}\cdot R_{\xi}$. Base-changing $\mu$ to $\operatorname{Spec}(R_{\xi})$ yields a proper birational morphism
$\mu_{\xi}:Z_{\xi}:=Z\times_{\overline{\Gamma}_{F}}\operatorname{Spec}(R_{\xi})\longrightarrow\operatorname{Spec}(R_{\xi}),$
with $Z_{\xi}$ normal, such that $\mathscr{I}_{X,\xi}\mathcal{O}_{Z_{\xi}}$ and $\mathscr{I}_{Y,\xi}\mathcal{O}_{Z_{\xi}}$ are invertible. Denote by $A_{\xi}$ and $B_{\xi}$ the pullbacks of $A$ and $B$, so that
$\mathscr{I}_{X,\xi}\mathcal{O}_{Z_{\xi}}=\mathcal{O}_{Z_{\xi}}(-A_{\xi})\text{ and }\mathscr{I}_{Y,\xi}\mathcal{O}_{Z_{\xi}}=\mathcal{O}_{Z_{\xi}}(-B_{\xi}).$
For our purposes, the only prime divisors on $Z_{\xi}$ that matter are those lying over codimension–one points of $Z$ whose image contains $\xi$; equivalently, whenever $E\subseteq Z$ is a prime divisor with $\xi\in\mu(E)$, its pullback to $Z_{\xi}$ has codimension–one points, and testing integral-closure containments on $R_{\xi}$ is reduced to testing the corresponding divisorial orders along such $E$.
Theorem 11.1 (Local finite–minimum formula).
Fix $\xi\in D_{X}$. With notation as above, one has
(11.3) $\mathcal{L}^{\text{an}}_{\infty,\xi}(F)=\min\left\{\frac{\operatorname{\mathrm{ord}}_{E}(\mathscr{I}_{Y})}{\operatorname{\mathrm{ord}}_{E}(\mathscr{I}_{X})}\ \middle|\ E\in\operatorname{\mathrm{Supp}}(A)\text{ and }\xi\in\mu(E)\right\}\ \in\ \mathbb{R}_{\geq 0}\cup\{\infty\}.$
Proof.
For simplicity of notation, set $r_{\xi}=\mathcal{L}_{\mathscr{I}_{Y,\xi}}(\mathscr{I}_{X,\xi})$, i.e.,
$r_{\xi}=\inf\left\{\frac{q}{p}\in\mathbb{Q}_{>0}\ \middle|\ \mathscr{I}_{Y,\xi}^{\,qt}\subseteq\overline{\mathscr{I}_{X,\xi}^{\,pt}}\ \text{ for all }t\gg 1\right\}.$
By Theorem 4.9, we have $\mathcal{L}^{\text{an}}_{\infty,\xi}(F)=1/r_{\xi}$ (with the conventions $\inf\varnothing=\infty$, $1/\infty=0$, and $1/0=\infty$).
Fix $\frac{q}{p}\in\mathbb{Q}_{>0}$ and $t\geq 1$. Since $\mathscr{I}_{X,\xi}\mathcal{O}_{Z_{\xi}}$ and $\mathscr{I}_{Y,\xi}\mathcal{O}_{Z_{\xi}}$ are invertible, their powers are integrally closed on $Z_{\xi}$. Hence, by the fixed–model integral–closure criterion (see Lemma 5.4), the containment
$\mathscr{I}_{Y,\xi}^{\,qt}\subseteq\overline{\mathscr{I}_{X,\xi}^{\,pt}}\quad\text{in }R_{\xi}$
is equivalent to the divisor inequality
$qt\,B_{\xi}\ \geq\ pt\,A_{\xi}\quad\text{on }Z_{\xi},$
and therefore to the collection of inequalities
(11.4) $qt\cdot\operatorname{\mathrm{ord}}_{E}(\mathscr{I}_{Y})\ \geq\ pt\cdot\operatorname{\mathrm{ord}}_{E}(\mathscr{I}_{X})\quad\text{for every prime divisor }E\subseteq Z\text{ with }\xi\in\mu(E).$
If $E\not\subseteq\operatorname{\mathrm{Supp}}(A)$ then $\operatorname{\mathrm{ord}}_{E}(\mathscr{I}_{X})=0$ and the inequalities (11.4) are vacuously true, i.e., such a prime divisor $E$ imposes no restriction on $q/p$. If $E\in\operatorname{\mathrm{Supp}}(A)$ and $\xi\in\mu(E)$, then $\operatorname{\mathrm{ord}}_{E}(\mathscr{I}_{X})>0$. For such $E$, the inequalities (11.4) are equivalent to the ratio inequalities
$\frac{q}{p}\ \geq\ \frac{\operatorname{\mathrm{ord}}_{E}(\mathscr{I}_{X})}{\operatorname{\mathrm{ord}}_{E}(\mathscr{I}_{Y})}\quad\text{whenever }\operatorname{\mathrm{ord}}_{E}(\mathscr{I}_{Y})>0,$
and if $\operatorname{\mathrm{ord}}_{E}(\mathscr{I}_{Y})=0$ then (11.4) fails for every $\frac{q}{p}>0$.
It follows that
$r_{\xi}=\sup\left\{\frac{\operatorname{\mathrm{ord}}_{E}(\mathscr{I}_{X})}{\operatorname{\mathrm{ord}}_{E}(\mathscr{I}_{Y})}\ \middle|\ E\in\operatorname{\mathrm{Supp}}(A)\text{ and }\xi\in\mu(E)\right\}\ \in\ [0,\infty],$
where the supremum is $\infty$ if some such $E$ satisfies $\operatorname{\mathrm{ord}}_{E}(\mathscr{I}_{Y})=0$. Taking reciprocals (with the stated conventions) yields (11.3). ∎
Global finiteness and a finite minimum. Recall that
$\operatorname{\mathrm{Supp}}(A)=\{\,E\subseteq Z\text{ prime divisor}\mid\operatorname{\mathrm{ord}}_{E}(\mathscr{I}_{X})>0\,\}.$
Theorem 11.2 (Global finite–minimum formula).
With notation as above, one has
(11.5) $\mathcal{L}^{\text{an}}_{\infty}(F)=\min_{E\in\operatorname{\mathrm{Supp}}(A)}\ \frac{\operatorname{\mathrm{ord}}_{E}(\mathscr{I}_{Y})}{\operatorname{\mathrm{ord}}_{E}(\mathscr{I}_{X})}\ \in\ \mathbb{R}_{\geq 0}\cup\{\infty\}.$
In particular,
$\mathcal{L}^{\text{an}}_{\infty}(F)>0\quad\Longleftrightarrow\quad\operatorname{\mathrm{ord}}_{E}(\mathscr{I}_{Y})>0\ \text{ for every }E\in\operatorname{\mathrm{Supp}}(A).$
Proof.
By Definition 4.8,
$\mathcal{L}^{\text{an}}_{\infty}(F)=\inf_{\xi\in D_{X}}\ \mathcal{L}^{\text{an}}_{\infty,\xi}(F).$
Applying Theorem 11.1 gives
$\mathcal{L}^{\text{an}}_{\infty}(F)=\inf_{\xi\in D_{X}}\ \min\left\{\frac{\operatorname{\mathrm{ord}}_{E}(\mathscr{I}_{Y})}{\operatorname{\mathrm{ord}}_{E}(\mathscr{I}_{X})}\ \middle|\ E\in\operatorname{\mathrm{Supp}}(A)\text{ and }\xi\in\mu(E)\right\}.$
Fix $E\in\operatorname{\mathrm{Supp}}(A)$. Since $\operatorname{\mathrm{ord}}_{E}(\mathscr{I}_{X})>0$, the image $\mu(E)$ is contained in $D_{X}$, so there exists $\xi\in D_{X}$ with $\xi\in\mu(E)$. Therefore, each $E\in\operatorname{\mathrm{Supp}}(A)$ occurs in the inner minimum for some $\xi$. Since $\operatorname{\mathrm{Supp}}(A)$ is finite, the outer infimum equals the minimum over $E\in\operatorname{\mathrm{Supp}}(A)$, proving (11.5). The final equivalence is immediate. ∎
Positivity and attainment at infinity. We now specialize the general finite–max and Rees valuations machinery developed in Section 5 to Łojasiewicz exponents at infinity.
Fix $\xi\in D_{X}\cap D_{Y}$. By the valuative criterion for integral closure (see Lemma 5.2), for fixed $p,q,t\geq 1$ the containment $\mathscr{I}_{Y,\xi}^{\,qt}\subseteq\overline{\mathscr{I}_{X,\xi}^{\,pt}}$ holds if and only if
[…]
Consequently, the optimal exponent is characterized by
(11.6) $\mathcal{L}^{\text{an}}_{\infty,\xi}(F)=\inf_{v}\ \frac{v(\mathscr{I}_{Y,\xi})}{v(\mathscr{I}_{X,\xi})},$
where the infimum ranges over all divisorial valuations $v$ of $R_{\xi}$ with $v(\mathscr{I}_{X,\xi})>0$. In fact, as pointed out in Lemma 5.2, it is enough to let $v$ range over finitely many Rees valuations of $\mathscr{I}_{X,\xi}$.
Theorem 11.3 (Positivity and attainment).
Fix $\xi\in D_{X}\cap D_{Y}$.
  1. (1)
    $\mathcal{L}^{\text{an}}_{\infty,\xi}(F)>0$ if and only if there is no divisorial valuation $v$ of $R_{\xi}$ such that
    $v(\mathscr{I}_{X,\xi})>0\quad\text{and}\quad v(\mathscr{I}_{Y,\xi})=0.$
  2. (2)
    If $\mathcal{L}^{\text{an}}_{\infty,\xi}(F)>0$, then the infimum in (11.6) is a minimum.
  3. (3)
    More precisely, if $\mathcal{L}^{\text{an}}_{\infty,\xi}(F)>0$, then $\mathcal{L}^{\text{an}}_{\infty,\xi}(F)$ is attained by one of the finitely many Rees valuations of the ideal $\mathscr{I}_{X,\xi}\subseteq R_{\xi}$.
Proof.
(1) If there exists a divisorial valuation $v$ with $v(\mathscr{I}_{X,\xi})>0$ and $v(\mathscr{I}_{Y,\xi})=0$, then the ratio $v(\mathscr{I}_{Y,\xi})/v(\mathscr{I}_{X,\xi})$ equals 0, so (11.6) gives $\mathcal{L}^{\text{an}}_{\infty,\xi}(F)=0$. Conversely, if no such valuation exists, then every divisorial valuation with $v(\mathscr{I}_{X,\xi})>0$ also satisfies $v(\mathscr{I}_{Y,\xi})>0$, so all ratios in (11.6) are positive, hence $\mathcal{L}^{\text{an}}_{\infty,\xi}(F)>0$.
(2)–(3) Assume $\mathcal{L}^{\text{an}}_{\infty,\xi}(F)>0$. Taking reciprocals in (11.6) gives
$\frac{1}{\mathcal{L}^{\text{an}}_{\infty,\xi}(F)}=\sup_{v}\ \frac{v(\mathscr{I}_{X,\xi})}{v(\mathscr{I}_{Y,\xi})},$
where the supremum ranges over divisorial valuations with $v(\mathscr{I}_{Y,\xi})>0$. By (1) this supremum is finite. Applying the finite–maximum theorem for ideal pairs (Theorem 5.17) to the pair $(\mathscr{I}_{X,\xi},\mathscr{I}_{Y,\xi})$ in the Noetherian local domain $R_{\xi}$, the supremum is attained by one of the finitely many Rees valuations of the ideal $\mathscr{I}_{X,\xi}$. Taking reciprocals again shows that the infimum in (11.6) is a minimum attained by the same finite set of valuations. ∎
11.2. Newton nondegeneracy and toric compactification
We recall the notion of Newton nondegeneracy for polynomial mappings (or ideals) following Saia [47]. This is slightly different than Kouchnirenko’s [34]). Particularly, a function $f\in k[x_{1},\dots,x_{n}]$ is Newton nondegenerate in Kouchnirenko’s sense if and only if the ideal generated by $x_{i}\frac{\partial f}{\partial x_{i}}$ is Newton nondegenerate in Saia’s sense.
Definition 11.4 (Newton polyhedron and nondegeneracy).
Write $F=(f_{1},\dots,f_{r})$ with $f_{j}\in k[x_{1},\dots,x_{n}]$. For $f=\sum_{\alpha\in\mathbb{N}^{n}}c_{\alpha}x^{\alpha}$ with $c_{\alpha}\neq 0$, set
$\operatorname{\mathrm{Supp}}(f):=\{\alpha\in\mathbb{N}^{n}:c_{\alpha}\neq 0\}.$
Define the Newton polyhedron of $F$ (or equivalently, of the ideal generated by $F$) by
$\Gamma(F):=\mathrm{conv}\Bigl(\bigcup_{j=1}^{r}\operatorname{\mathrm{Supp}}(f_{j})\cup\{0\}\Bigr)+\mathbb{R}_{\geq 0}^{n}\ \subseteq\ \mathbb{R}^{n},$
and for $w\in\mathbb{R}_{\geq 0}^{n}$ define
$m_{w}(f):=\min_{\alpha\in\operatorname{\mathrm{Supp}}(f)}\langle w,\alpha\rangle,\quad f_{w}:=\sum_{\alpha\in\operatorname{\mathrm{Supp}}(f)\,:\,\langle w,\alpha\rangle=m_{w}(f)}c_{\alpha}x^{\alpha}.$
We say that $F$ is Newton nondegenerate at infinity if for every $w\in\mathbb{R}_{>0}^{n}$ the system
$(f_{1})_{w}=\cdots=(f_{r})_{w}=0$
has no solution in $(k^{*})^{n}$.
We will use a toric compactification adapted to the Newton polyhedron at infinity following Khovanskii [33]. Let $\Sigma_{\Gamma}$ denote the normal fan of $\Gamma(F)$ and let $\Sigma_{\mathbb{P}^{n}}$ denote the standard fan of $\mathbb{P}^{n}$. Choose a smooth complete fan $\Sigma$ refining both $\Sigma_{\Gamma}$ and $\Sigma_{\mathbb{P}^{n}}$ (such a refinement exists by toric resolution of singularities; see [33, 42], [11, Chapter 11] and [18, Chapter 2]). Then, the associated toric variety $X_{\Sigma}$ is a smooth projective toric compactification of the torus $T=(k^{*})^{n}$, and the refinement $\Sigma\succ\Sigma_{\mathbb{P}^{n}}$ induces a proper toric morphism $\pi_{\Sigma}:X_{\Sigma}\longrightarrow\mathbb{P}^{n}$ extending the standard open embedding $T\hookrightarrow\mathbb{P}^{n}$ (see [11, Chapter 3] and [18, Chapters 1–2]). Moreover, since $\Sigma$ also refines the normal fan of $\Gamma(F)$, the toric boundary divisors of $X_{\Sigma}$ correspond to weight vectors governing the face systems (initial forms) of $F$, which control the asymptotic behavior at infinity as in Khovanskii’s compactification.
Notation 11.5 (Khovanskii compactification).
Let $\Sigma$ be a smooth complete fan refining both the normal fan of $\Gamma(F)$ and the fan of $\mathbb{P}^{n}$, and let
$\pi_{\Sigma}:X_{\Sigma}\to\mathbb{P}^{n}$
be the induced toric morphism. For each cone $\sigma\in\Sigma$, let $O(\sigma)\subseteq X_{\Sigma}$ denote the corresponding torus orbit. Let $\Gamma_{F}^{\Sigma}\subseteq X_{\Sigma}\times\mathbb{P}^{r}$ denote the normalization of the closure of the graph of $F$. Write
$\pi_{\Sigma}:\Gamma_{F}^{\Sigma}\to X_{\Sigma}\text{ and }\rho_{\Sigma}:\Gamma_{F}^{\Sigma}\to\mathbb{P}^{r}$
for the two natural projections.
Lemma 11.6 (Values along divisors at infinity).
Assume that $F:\mathbb{A}^{n}_{k}\to\mathbb{A}^{r}_{k}$ is Newton nondegenerate at infinity in the sense of Definition 11.4. Let $\xi\in D_{X}\cap D_{Y}$ and set $R_{\xi}:=\mathcal{O}_{\overline{\Gamma}_{F},\xi}$. Then, for every divisorial valuation $v$ of $R_{\xi}$,
$v(\mathscr{I}_{X,\xi})>0\quad\Longrightarrow\quad v(\mathscr{I}_{Y,\xi})>0.$
Equivalently, there is no divisorial valuation $v$ with $v(\mathscr{I}_{X,\xi})>0$ and $v(\mathscr{I}_{Y,\xi})=0$.
Proof.
We work on a toric compactification attached to $\Gamma(F)$ and then descend to divisorial valuations of $R_{\xi}$.
Step 1: toric boundary data and face systems. Let $\Sigma$ and $\pi_{\Sigma}:X_{\Sigma}\to\mathbb{P}^{n}$ be as above (Khovanskii’s compactification). Write $D_{\operatorname{Tor}}:=X_{\Sigma}\setminus(k^{*})^{n}$ for the toric boundary. Each torus–invariant prime divisor $E\subseteq D_{\operatorname{Tor}}$ corresponds to a primitive ray vector $w_{E}\in\mathbb{Z}_{\geq 0}^{n}\setminus\{0\}$. For every Laurent monomial $x^{\alpha}$ one has $\operatorname{\mathrm{ord}}_{E}(x^{\alpha})=\langle w_{E},\alpha\rangle$, and for every polynomial $f$ one has $\operatorname{\mathrm{ord}}_{E}(f)=m_{w_{E}}(f)$ with initial form $f_{w_{E}}$ on the orbit $O(E):=O(w_{E})$.
Step 2: boundary intersections of the graph and Newton nondegeneracy. Let $\Gamma_{F}^{\Sigma}\subseteq X_{\Sigma}\times\mathbb{P}^{r}$ be as before, and let $\mu_{\Sigma}:\Gamma_{F}^{\Sigma}\to\overline{\Gamma}_{F}$ be the induced proper birational morphism.
Fix a torus–invariant divisor $E\subseteq D_{\operatorname{Tor}}$ and consider the locally closed stratum $O(E)\subseteq X_{\Sigma}$. Work on the target chart $\{W_{0}\neq 0\}\subseteq\mathbb{P}^{r}$, with affine coordinates $y_{j}:=W_{j}/W_{0}$. Restricting to $O(E)\times\mathbb{A}^{r}$, we work in an affine toric chart $U_{\sigma}=\operatorname{Spec}k[\sigma^{\vee}\cap M]$ containing $O(E)$, where $\sigma\in\Sigma$ is a cone whose rays include $w_{E}$. In the corresponding toric coordinates $u_{1},\dots,u_{n}$, each coordinate $x_{i}$ is a monomial in the $u_{j}$, and each polynomial $f_{j}(x)$ admits a decomposition of the form
$f_{j}(x)=u^{m_{j}}\Bigl((f_{j})_{w_{E}}(u)+\sum_{\langle w_{E},\alpha\rangle>m_{j}}c_{j,\alpha}u^{\alpha}\Bigr),$
where $m_{j}=\min_{\alpha\in\Gamma(f_{j})}\langle w_{E},\alpha\rangle$, and $(f_{j})_{w_{E}}$ is the face polynomial consisting of those terms achieving this minimum. Consequently, the strict transform of the graph in $U_{\sigma}\times\mathbb{A}^{r}$ meets $O(E)\times\mathbb{A}^{r}$ if and only if the face system
$(f_{1})_{w_{E}}=\cdots=(f_{r})_{w_{E}}=0$
admits a solution in the torus $O(E)\cong(k^{*})^{n}$. This is the standard initial-form criterion for strict transforms in toric modifications; see, for example, [33, 42]. In particular,
(11.7) $\Gamma_{F}^{\Sigma}\cap\bigl(O(E)\times\{W_{0}\neq 0\}\bigr)=\varnothing\Leftrightarrow(f_{1})_{w_{E}}=\cdots=(f_{r})_{w_{E}}=0\ \text{has no solution in }(k^{*})^{n}.$
Since $F$ is Newton nondegenerate at infinity, the right-hand condition holds for every $w_{E}\in\mathbb{R}_{>0}^{n}$, and so (11.7) implies: for every boundary divisor $E$ with $w_{E}\in\mathbb{R}_{>0}^{n}$,
$\Gamma_{F}^{\Sigma}\cap\bigl(O(E)\times\{W_{0}\neq 0\}\bigr)=\varnothing,$
i.e. every point of $\Gamma_{F}^{\Sigma}$ lying over $O(E)$ maps into the hyperplane at infinity $H_{Y}=\{W_{0}=0\}$. Equivalently, $\rho_{\Sigma}^{-1}(H_{Y})$ contains $\mu_{\Sigma}^{-1}(\pi_{\Sigma}^{-1}(H_{X}))$ along all such boundary strata.
Step 3: descend to divisorial valuations over $\xi$. Let $v$ be a divisorial valuation of the local domain $R_{\xi}=\mathcal{O}_{\overline{\Gamma}_{F},\xi}$. Choose a normal modification $\nu:W\to\overline{\Gamma}_{F}$ and a prime divisor $E^{\prime}\subseteq W$ with center $\xi$ such that $v=\operatorname{\mathrm{ord}}_{E^{\prime}}$ on $R_{\xi}$. Take a common normal refinement $W^{\prime}\to W$ and $W^{\prime}\to\Gamma_{F}^{\Sigma}$ dominating both models, and let $\widetilde{E}^{\prime}\subseteq W^{\prime}$ be the strict transform of $E^{\prime}$. Then
[…]
Assume $v(\mathscr{I}_{X,\xi})>0$. This means that $\widetilde{E}^{\prime}$ lies over $D_{X}=\pi^{-1}(H_{X})$, so its image in $X_{\Sigma}$ is contained in the toric boundary lying over $H_{X}$. Hence, $\widetilde{E}^{\prime}$ dominates (after possibly replacing $\Sigma$ by a further refinement) a torus–invariant boundary divisor $E\subseteq D_{\operatorname{Tor}}$ with ray vector $w_{E}\in\mathbb{R}_{>0}^{n}$. Indeed, divisors corresponding to rays lying on coordinate faces of the cone $\mathbb{R}_{\geq 0}^{n}$ map into the coordinate boundary of $\mathbb{A}^{n}_{k}\subseteq\mathbb{P}^{n}$ rather than into the hyperplane at infinity $H_{X}=\{X_{0}=0\}$. Since $\widetilde{E}^{\prime}$ lies over $D_{X}=\pi^{-1}(H_{X})$, it must dominate a torus-invariant divisor corresponding to a ray in the relative interior $\mathbb{R}_{>0}^{n}$ of the cone. This is a standard property of toric compactifications attached to Newton polyhedra; see [33, 42]. By Step 2, $\Gamma_{F}^{\Sigma}$ over $O(E)$ maps into $H_{Y}$, so $\widetilde{E}^{\prime}$ is contained in the pullback of $D_{Y}=\rho^{-1}(H_{Y})$. Equivalently, $\operatorname{\mathrm{ord}}_{\widetilde{E}^{\prime}}(\mathscr{I}_{Y}\mathcal{O}_{W^{\prime}})>0$, i.e. $v(\mathscr{I}_{Y,\xi})>0$.
This proves $v(\mathscr{I}_{X,\xi})>0\Rightarrow v(\mathscr{I}_{Y,\xi})>0$. The second statement is immediate. ∎
Corollary 11.7.
Assume that $F$ is Newton nondegenerate at infinity. Then, for every $\xi\in D_{X}\cap D_{Y}$ one has
$\mathcal{L}^{\text{an}}_{\infty,\xi}(F)>0.$
Proof.
Fix $\xi\in D_{X}\cap D_{Y}$. By Lemma 11.6, there is no divisorial valuation $v$ of $R_{\xi}$ with $v(\mathscr{I}_{X,\xi})>0$ and $v(\mathscr{I}_{Y,\xi})=0$. The conclusion follows immediately from Theorem 11.3(1). ∎
Remark 11.8.
At this stage, Theorem 11.3 guarantees finiteness and attainment by a finite set of Rees valuations of $\mathscr{I}_{X,\xi}$. In the Newton nondegenerate regime, we can do better: we identify the relevant divisors with explicit toric boundary divisors arising from a toric compactification adapted to $\Gamma(F)$, leading to concrete Newton polyhedral formulas.
11.3. Newton nondegenerate at infinity and Łojasiewicz exponent
We will show how classical formulas for Łojasiewicz exponents at infinity in the Newton nondegenerate case are recovered within our algebraic framework. Let $F:\mathbb{A}^{n}_{k}\to\mathbb{A}^{r}_{k}$ denote a polynomial mapping, that is Newton nondegeneracy at infinity as in Definition 11.4.
Proposition 11.9.
Assume that $F$ is Newton nondegenerate at infinity. Then, there exists a toric compactification $X_{\Sigma}$ adapted to $\Gamma(F)$ and a proper birational morphism
$\mu:Z\longrightarrow\overline{\Gamma}_{F}$
with $Z$ smooth such that:
  1. (1)
    the reduced divisor $(D_{X})_{\text{red}}+(D_{Y})_{\text{red}}$ pulls back to an SNC divisor
    $E_{1}+\cdots+E_{N}\quad\text{on }Z;$
  2. (2)
    for every $\xi\in D_{X}\cap D_{Y}$, after base change to $\operatorname{Spec}(R_{\xi})$ the ideals $\mathscr{I}_{X,\xi}$ and $\mathscr{I}_{Y,\xi}$ pull back to invertible ideals supported on the strict transforms of those $E_{i}$ meeting $\mu^{-1}(\xi)$.
Proof.
Choose a toric compactification $X_{\Sigma}$ adapted to $\Gamma(F)$. Since $X_{\Sigma}$ is proper over $\mathbb{P}^{n}$, the normalization $\Gamma_{F}^{\Sigma}$ of the graph closure in $X_{\Sigma}\times\mathbb{P}^{r}$ is proper over $\overline{\Gamma}_{F}$, and so there is a proper birational morphism $\Gamma_{F}^{\Sigma}\to\overline{\Gamma}_{F}$. Now take a log resolution of the reduced divisor
$\bigl(\mu_{\Sigma}^{-1}(D_{X})\bigr)_{\text{red}}+\bigl(\mu_{\Sigma}^{-1}(D_{Y})\bigr)_{\text{red}}$
on $\Gamma_{F}^{\Sigma}$; composing this resolution with $\Gamma_{F}^{\Sigma}\to\overline{\Gamma}_{F}$ produces a proper birational morphism $\mu:Z\to\overline{\Gamma}_{F}$ with $Z$ smooth, such that the reduced pullback $(D_{X})_{\text{red}}+(D_{Y})_{\text{red}}$ has SNC support. Denote its irreducible components by $E_{1},\dots,E_{N}$, proving (1).
For (2), fix $\xi\in D_{X}\cap D_{Y}$ and base change $Z\to\overline{\Gamma}_{F}$ to $\operatorname{Spec}(R_{\xi})$. On the smooth scheme $Z_{\xi}:=Z\times_{\overline{\Gamma}_{F}}\operatorname{Spec}(R_{\xi})$, the pullbacks of $\mathscr{I}_{X,\xi}$ and $\mathscr{I}_{Y,\xi}$ define Cartier divisors supported on the pullback of $(D_{X})_{\text{red}}+(D_{Y})_{\text{red}}$ and, hence, on the strict transforms of those $E_{i}$ meeting $\mu^{-1}(\xi)$. In particular, both ideals become invertible on $Z_{\xi}$ with support contained in those components, as claimed. ∎
Theorem 11.10.
Assume that $F$ is Newton nondegenerate at infinity. Then, for every $\xi\in D_{X}\cap D_{Y}$ one has
$0<\mathcal{L}^{\text{an}}_{\infty,\xi}(F)<\infty.$
Moreover, on the fixed model $Z$ of Proposition 11.9, one has
$\mathcal{L}^{\text{an}}_{\infty,\xi}(F)=\min\left\{\frac{\operatorname{\mathrm{ord}}_{E_{i}}(\mathscr{I}_{Y})}{\operatorname{\mathrm{ord}}_{E_{i}}(\mathscr{I}_{X})}\ \middle|\ 1\leq i\leq N,\ \xi\in\mu(E_{i})\right\}.$
Proof.
Fix $\xi\in D_{X}\cap D_{Y}$. By Proposition 11.9, on the base change of $Z$ to $\operatorname{Spec}(R_{\xi})$ both $\mathscr{I}_{X,\xi}$ and $\mathscr{I}_{Y,\xi}$ pull back to invertible ideals supported on the strict transforms of $E_{1},\dots,E_{N}$ meeting $\mu^{-1}(\xi)$. Therefore, the local finite–minimum formula Theorem 11.1 applies on this fixed model and yields
$\mathcal{L}^{\text{an}}_{\infty,\xi}(F)=\min\left\{\frac{\operatorname{\mathrm{ord}}_{E}(\mathscr{I}_{Y})}{\operatorname{\mathrm{ord}}_{E}(\mathscr{I}_{X})}\ \middle|\ E\subseteq Z\text{ prime divisor with }\operatorname{\mathrm{ord}}_{E}(\mathscr{I}_{X})>0,\ \xi\in\mu(E)\right\}.$
Since every such $E$ is among the components $E_{i}$, this gives the stated formula. Positivity follows from Lemma 11.6: if $\operatorname{\mathrm{ord}}_{E_{i}}(\mathscr{I}_{X})>0$ and $\xi\in\mu(E_{i})$, then the corresponding divisorial valuation on $R_{\xi}$ satisfies $v(\mathscr{I}_{X,\xi})>0$, and so $v(\mathscr{I}_{Y,\xi})>0$, i.e. $\operatorname{\mathrm{ord}}_{E_{i}}(\mathscr{I}_{Y})>0$. Thus, every ratio in the above minimum is a positive real number. Hence, $0<\mathcal{L}^{\text{an}}_{\infty,\xi}(F)<\infty$ as claimed. ∎
Corollary 11.11 (Complex Newton–nondegenerate polynomial mappings).
Let $F:\mathbb{C}^{n}\to\mathbb{C}^{m}$ be a polynomial mapping that is Newton nondegenerate at infinity in the sense of Definition 11.4. Then, the Łojasiewicz exponent at infinity $\mathcal{L}^{\text{an}}_{\infty}(F)$ is finite and is computed by finitely many divisorial valuations coming from a toric compactification adapted to the Newton polyhedron $\Gamma(F)$.
Proof.
By Newton nondegeneracy at infinity, Lemma 11.6 shows that no divisorial valuation $v$ of $\mathcal{O}_{\overline{\Gamma}_{F},\xi}$ satisfies $v(\mathscr{I}_{X,\xi})>0$ and $v(\mathscr{I}_{Y,\xi})=0$. Therefore, $\mathcal{L}^{\text{an}}_{\infty,\xi}(F)>0$ for every $\xi\in D_{X}\cap D_{Y}$ by Theorem 11.3(1). The finite–minimum formula of Theorem 11.10 then expresses $\mathcal{L}^{\text{an}}_{\infty,\xi}(F)$, and hence $\mathcal{L}^{\text{an}}_{\infty}(F)$, as a minimum over finitely many divisorial valuations on a fixed toric model associated to $\Gamma(F)$. ∎
When $n=2$, the toric divisorial valuations appearing in Corollary 11.11 correspond bijectively to the edges of the Newton polygon at infinity.
Corollary 11.12 (Plane polynomial mappings).
Let $F:\mathbb{C}^{2}\to\mathbb{C}^{2}$ be a polynomial mapping that is Newton nondegenerate at infinity. Then, $\mathcal{L}^{\text{an}}_{\infty}(F)$ is finite and can be computed explicitly from the Newton polygon at infinity $\Gamma(F)$ by a finite set of rational numbers determined by its edges.
Proof.
By Corollary 11.11, $\mathcal{L}^{\text{an}}_{\infty}(F)$ is given by a finite minimum over divisorial valuations arising from a toric compactification adapted to $\Gamma(F)$. In dimension two, these divisorial valuations are exactly the toric valuations associated to the edges of the Newton polygon. By results in Section 10, the relevant orders $\operatorname{\mathrm{ord}}_{E}(\mathscr{I}_{X})$ and $\operatorname{\mathrm{ord}}_{E}(\mathscr{I}_{Y})$ are computed by the support functions of $\Gamma(F)$ along those edges. Substituting into the finite–minimum formula of Theorem 11.10 yields the stated Newton–polygon expression. ∎
Corollaries 11.11 and 11.12 recover classical results on Łojasiewicz exponent at infinity in the Newton nondegenerate case; see, for instance, [44].
Remark 11.13.
Over $\mathbb{R}$, the finiteness of Łojasiewicz exponent at infinity is more subtle. This is mainly because we do not have the equivalence in Theorem 4.9 nor a single birational model which realizes all relevant valuations. Nevertheless, similar results in Newton nondegenerate situations, where divisors at infinity are torus-invariant and relevant valuations are toric valuations associated to finitely many rays/facets (supporting hyperplanes of the corresponding Newton polyhedra), were obtained under various additional hypotheses (see, for example, [24, 41, 43]).
Part VII Necessity and Limits of the Hypotheses
In this part, we examine verification, uniformity, and sharpness issues for the finite–max principle and for the Newton–type formulas arising from the valuative framework.
12. Uniformity, verification, and sharpness
This section addresses when the finite testing hypotheses from Sections 5 and 10 can be verified uniformly in families, and why these hypotheses are close to optimal. On the positive side, in the Newton nondegenerate regime we prove the existence of a single toric modification simultaneously principalizing finitely many ideals, producing a uniform finite divisorial testing set across an entire collection of data (Theorem 12.1); as consequences we obtain automatic finite testing for broad Newton-controlled families (Theorem 12.2) and a genuine polyhedral wall–chamber structure governing the variation of $\mathcal{L}$ (Theorem 12.3). On the sharpness side, we construct monomial examples where successive principalizations introduce new relevant facets, showing that outside verifiable finite-testing regimes one cannot expect any a priori finite reduction of the valuative optimization problem (Theorem 12.4).
12.1. Newton polyhedra for ideals
[…]
12.2. Uniform verification on a fixed toric model
We now explain how, in Newton nondegenerate situations, the hypotheses of the log-resolution criterion in Section 5 (and its integral-closure reduction Lemma 5.4) are automatically satisfied for a large class of filtrations.
Theorem 12.1 (Toric model simultaneously principalizing finitely many Newton nondegenerate ideals).
Let $(R,\mathfrak{m})$ be an excellent equicharacteristic regular local ring with chosen parameters $x=(x_{1},\dots,x_{n})$. Let $I^{(1)},\dots,I^{(r)}\subseteq R$ be ideals such that each $I^{(j)}$ is Newton nondegenerate with respect to $x$ (i.e., the relevant face ideals define smooth complete intersections on the torus). Then, there exists a proper birational morphism $\pi:Y\to\operatorname{Spec}R$ such that:
  1. (1)
    $Y$ is regular and the exceptional divisor $E=\sum_{i=1}^{N}E_{i}$ has simple normal crossings.
  2. (2)
    For each $j=1,\dots,r$, the total transform of $I^{(j)}$ is an invertible ideal:
    $I^{(j)}\mathcal{O}_{Y}\ =\ \mathcal{O}_{Y}(-F_{j}),\quad F_{j}=\sum_{i=1}^{N}a_{ij}E_{i},\ \ a_{ij}\in\mathbb{N}.$
  3. (3)
    The morphism $\pi$ can be chosen toric with respect to the chosen parameters $x$, coming from a regular subdivision of a fan refining the Newton fans of $\Gamma(I^{(j)})$.
Proof.
We retain the notation and terminology of [22] for Newton nondegeneracy with respect to a chosen regular system of parameters.
[…]
Step 2: construction of a common toric modification. For each $j$, let $\Gamma(I^{(j)})\subseteq\mathbb{R}_{\geq 0}^{n}$ be the Newton polyhedron with respect to the chosen parameters $x=(x_{1},\dots,x_{n})$, and let $\Sigma^{(j)}$ denote its Newton fan. Since we have finitely many such fans, there exists a common rational polyhedral refinement $\Sigma^{\prime}$ refining all $\Sigma^{(j)}$. Choose a regular subdivision $\Sigma$ of $\Sigma^{\prime}$, and let
$\pi=\pi_{\Sigma}:Y=X_{\Sigma}\longrightarrow\operatorname{Spec}R$
be the associated toric morphism.
Because $\Sigma$ is regular, $Y$ is regular and its toric boundary divisor $E=\sum_{i=1}^{N}E_{i}$ is a simple normal crossings divisor. This proves (1). Property (3) holds by construction.
Step 3: principalization of each $I^{(j)}$. Fix $j\in\{1,\dots,r\}$.
Since $\Sigma$ refines the Newton fan of $\Gamma(I^{(j)})$, Khovanskii’s toroidal resolution theorem (see [33, 42]) for Newton nondegenerate systems applies. Specifically, for Newton nondegenerate ideals with respect to the parameters $x$, the toric modification associated to a regular subdivision of the Newton fan gives a good resolution and principalization. In particular, $I^{(j)}\mathcal{O}_{Y}$ is an invertible ideal sheaf supported on the toric boundary.
Step 4: computation of the coefficients $a_{ij}$. Let $E_{1},\dots,E_{N}$ be the torus-invariant prime divisors of $Y$ lying over the closed point, and let $w_{i}\in\mathbb{Z}_{\geq 0}^{n}$ be the primitive ray vector corresponding to $E_{i}$. Then, the order of vanishing of a polynomial $f$ along $E_{i}$ is given by (see [33, 42]):
$\operatorname{\mathrm{ord}}_{E_{i}}(f)=\min_{\alpha\in\Gamma(f)}\langle w_{i},\alpha\rangle.$
Applying this to generators of $I^{(j)}$ yields
$\operatorname{\mathrm{ord}}_{E_{i}}\bigl(I^{(j)}\mathcal{O}_{Y}\bigr)=\min_{\alpha\in\Gamma(I^{(j)})}\langle w_{i},\alpha\rangle=:a_{ij}\in\mathbb{N}.$
Since $I^{(j)}\mathcal{O}_{Y}$ is invertible and determined by these divisorial orders, we obtain
$I^{(j)}\mathcal{O}_{Y}=\mathcal{O}_{Y}\!\left(-\sum_{i=1}^{N}a_{ij}E_{i}\right).$
This proves (2) for each $j$. ∎
Theorem 12.2 (Automatic finite testing for Newton controlled families).
Assume the setup of Theorem 12.1, and fix real numbers $\lambda_{1},\dots,\lambda_{r}\geq 0$. Define a graded family of ideals
$\mathfrak{a}_{t}\ :=\ \overline{\prod_{j=1}^{r}\big(I^{(j)}\big)^{\lceil\lambda_{j}t\rceil}}\quad(t\in\mathbb{R}_{\geq 0}),$
where $\overline{(\cdot)}$ denotes integral closure. Then, on the same model $\pi:Y\to\operatorname{Spec}R$ of Theorem 12.1, one has for every $t$:
$\mathfrak{a}_{t}\ =\ \pi_{*}\mathcal{O}_{Y}\Big(-\sum_{i=1}^{N}\Big\lfloor t\cdot\sum_{j=1}^{r}\lambda_{j}a_{ij}\Big\rfloor\,E_{i}\Big).$
In particular, for any other family $\mathfrak{b}_{\bullet}$} of the same form (possibly with different coefficients), the containment relations $\mathfrak{b}_{q}\subseteq\mathfrak{a}_{p}$ are determined by the finitely many divisorial valuations $\{\operatorname{\mathrm{ord}}_{E_{i}}\}_{i=1}^{N}$, i.e., the hypotheses of the finite-testing criterion in Theorem 5.8 are satisfied with testing set $\{\operatorname{\mathrm{ord}}_{E_{i}}\}$.
Proof.
Let $\pi:Y\to\operatorname{Spec}R$ be as in Theorem 12.1. As constructed, $Y$ is regular, so $Y$ is normal. Write $I^{(j)}\mathcal{O}_{Y}=\mathcal{O}_{Y}(-F_{j})$ with $F_{j}=\sum_{i}a_{ij}E_{i}$. Set
[…]
Step 1: compute integral closures on a fixed principalization. Since $I^{(j)}\mathcal{O}_{Y}$ is invertible, so is $\prod_{j}(I^{(j)})^{\lceil\lambda_{j}t\rceil}\mathcal{O}_{Y}$, and in fact
$\Big(\prod_{j=1}^{r}(I^{(j)})^{\lceil\lambda_{j}t\rceil}\Big)\mathcal{O}_{Y}\ =\ \mathcal{O}_{Y}\big(-D(t)\big).$
By the fixed-model integral-closure reduction (Lemma 5.4), the integral closure of $\prod_{j}(I^{(j)})^{\lceil\lambda_{j}t\rceil}$ equals the pushforward of the corresponding invertible sheaf:
$\mathfrak{a}_{t}\ =\ \overline{\prod_{j=1}^{r}(I^{(j)})^{\lceil\lambda_{j}t\rceil}}\ =\ \pi_{*}\mathcal{O}_{Y}\big(-D(t)\big).$
Step 2: replace $\lceil\lambda_{j}t\rceil$ by linear bounds in $t$. Write $D(t)=\sum_{i}d_{i}(t)E_{i}$ as above. Since $d_{i}(t)$ is an integer-valued, piecewise linear function in $t$ with slopes $\sum_{j}\lambda_{j}a_{ij}$, we have
$d_{i}(t)\ =\ \Big\lfloor t\cdot\sum_{j=1}^{r}\lambda_{j}a_{ij}\Big\rfloor\ +\ \epsilon_{i}(t),\quad\epsilon_{i}(t)\in\{0,1,\dots,r\}.$
Absorbing the bounded error into flooring yields the displayed formula; this is the same bookkeeping already used in Section 5 when passing between integer- and real-indexed families.
Step 3: finite testing. For any $p,q$, the condition $\mathfrak{b}_{q}\subseteq\mathfrak{a}_{p}$ is equivalent to the inequality of divisors on $Y$: $D_{\mathfrak{b}_{\bullet}}(q)\ \geq\ D_{\mathfrak{a}_{\bullet}}(p)$, i.e., coefficientwise inequalities along $E_{i}$. Equivalently, $\operatorname{\mathrm{ord}}_{E_{i}}(\mathfrak{b}_{q})\geq\operatorname{\mathrm{ord}}_{E_{i}}(\mathfrak{a}_{p})$ for all $i$. This is precisely the finite-testing hypothesis in Theorem 5.8 with testing set $\{\operatorname{\mathrm{ord}}_{E_{i}}\}$. ∎
12.3. Uniformity in families and wall–chamber decomposition
We now push the previous verification to a genuine structural statement in parameter families.
Theorem 12.3 (Polyhedral wall–chamber structure).
Assume the setup of Theorem 12.2, and let $S\subseteq\mathbb{R}^{m}$ be a parameter set. Suppose that for each $s\in S$ we have two families
${\mathfrak{a}_{\bullet}}(s)=\{\mathfrak{a}(s)_{t}\}_{t\geq 0},\quad{\mathfrak{b}_{\bullet}}(s)=\{\mathfrak{b}(s)_{t}\}_{t\geq 0},$
each of the form in Theorem 12.2, with coefficients $\lambda_{j}=\lambda_{j}(s)$ depending piecewise linearly on $s$. Assume moreover that the Newton polyhedra $\Gamma(I^{(j)}_{s})$ are independent of $s$ and the nondegeneracy is uniform, so that a single toric model $\pi:Y\to\operatorname{Spec}R$ principalizes all $I^{(j)}_{s}$ simultaneously.
Then, there exists a finite polyhedral decomposition $S=\bigcup_{\alpha}C_{\alpha}$ such that on each chamber $C_{\alpha}$ there is an index $i(\alpha)\in\{1,\dots,N\}$ with
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}(s)}\big({\mathfrak{a}_{\bullet}}(s)\big)\ =\ \frac{\operatorname{\mathrm{ord}}_{E_{i(\alpha)}}\big({\mathfrak{a}_{\bullet}}(s)\big)}{\operatorname{\mathrm{ord}}_{E_{i(\alpha)}}\big({\mathfrak{b}_{\bullet}}(s)\big)}\quad\text{for all }s\in C_{\alpha}.$
In particular, $s\mapsto\mathcal{L}_{{\mathfrak{b}_{\bullet}}(s)}({\mathfrak{a}_{\bullet}}(s))$ is piecewise rational-linear and the maximizing divisor is locally constant on chambers.
Proof.
Fix the common model $\pi:Y\to\operatorname{Spec}R$ as in Theorem 12.2 and its exceptional prime divisors $E_{1},\dots,E_{N}$. Write $E=\sum_{i=1}^{N}E_{i}$.
Step 1: linearity of divisor coefficients. By Theorem 12.2, for each $s$ we can write
$\mathfrak{a}(s)_{t}=\pi_{*}\mathcal{O}_{Y}\big(-D^{(s)}_{\mathfrak{a}}(t)\big),\quad\mathfrak{b}(s)_{t}=\pi_{*}\mathcal{O}_{Y}\big(-D^{(s)}_{\mathfrak{b}}(t)\big),$
where
$D^{(s)}_{\mathfrak{a}}(t)=\sum_{i=1}^{N}d^{(s)}_{\mathfrak{a},i}(t)\,E_{i},\quad D^{(s)}_{\mathfrak{b}}(t)=\sum_{i=1}^{N}d^{(s)}_{\mathfrak{b},i}(t)\,E_{i}$
and each coefficient $d^{(s)}_{\mathfrak{a},i}(t)$ (resp. $d^{(s)}_{\mathfrak{b},i}(t)$) is an integer-valued piecewise linear function whose slope in $t$ is a piecewise linear function of $s$. Hence, for each fixed $i$ the asymptotic slopes
$\mu_{\mathfrak{a},i}(s):=\lim_{t\to\infty}\frac{1}{t}\operatorname{\mathrm{ord}}_{E_{i}}\big(\mathfrak{a}(s)_{t}\big),\quad\mu_{\mathfrak{b},i}(s):=\lim_{t\to\infty}\frac{1}{t}\operatorname{\mathrm{ord}}_{E_{i}}\big(\mathfrak{b}(s)_{t}\big)$
exist and are piecewise linear functions of $s$.
Step 2: finite-max formula on the fixed testing set. By the finite-testing criterion, Theorem 5.8, with testing set $\{\operatorname{\mathrm{ord}}_{E_{i}}\}$ (we can restrict to those with $\mu_{\mathfrak{b},i}(s)>0$),
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}(s)}({\mathfrak{a}_{\bullet}}(s))\ =\ \max_{1\leq i\leq N}\ \frac{\mu_{\mathfrak{a},i}(s)}{\mu_{\mathfrak{b},i}(s)}.$
For each ordered pair $(i,j)$, the inequality
$\frac{\mu_{\mathfrak{a},i}(s)}{\mu_{\mathfrak{b},i}(s)}\ \geq\ \frac{\mu_{\mathfrak{a},j}(s)}{\mu_{\mathfrak{b},j}(s)}$
is equivalent to the linear inequality
$\mu_{\mathfrak{a},i}(s)\mu_{\mathfrak{b},j}(s)-\mu_{\mathfrak{a},j}(s)\mu_{\mathfrak{b},i}(s)\ \geq\ 0.$
By Theorem 12.2, for each fixed exceptional divisor $E_{i}$ the functions
$\mu_{\mathfrak{a},i}(t)=\operatorname{\mathrm{ord}}_{E_{i}}(\mathfrak{a}_{t})\quad\text{and}\quad\mu_{\mathfrak{b},i}(s)=\operatorname{\mathrm{ord}}_{E_{i}}(\mathfrak{b}_{s})$
are piecewise affine linear functions of $t$ and $s$, respectively. More precisely, there exist real constants $c_{\mathfrak{a},i},c_{\mathfrak{b},i}$ and bounded error terms $\varepsilon_{\mathfrak{a},i}(t),\varepsilon_{\mathfrak{b},i}(s)$ such that
$\mu_{\mathfrak{a},i}(t)=c_{\mathfrak{a},i}\,t+\varepsilon_{\mathfrak{a},i}(t),\quad\mu_{\mathfrak{b},i}(s)=c_{\mathfrak{b},i}\,s+\varepsilon_{\mathfrak{b},i}(s),$
where $\varepsilon_{\mathfrak{a},i}(t)$ (resp. $\varepsilon_{\mathfrak{b},i}(s)$) is locally constant away from a finite set of jump points determined by the ceiling functions $\lceil\lambda_{j}t\rceil$ (resp. $\lceil\lambda_{j}s\rceil$). Consequently, on each region of the (s,t)(s,t)–plane where these ceiling functions are fixed, the functions $\mu_{\mathfrak{a},i}(t)$ and $\mu_{\mathfrak{b},i}(s)$ are affine linear. Hence, refining $S$ by the finite arrangement of the linearity regions of all $\mu_{\mathfrak{a},i},\mu_{\mathfrak{b},i}$ yields a finite polyhedral decomposition on which all comparisons are given by linear inequalities.
Step 3: chambers and constancy of maximizers. Intersecting the regions defined by the finitely many pairwise comparison inequalities for the index attaining the maximum produces the claimed chambers $C_{\alpha}$, and on each chamber the maximizing index $i(\alpha)$ is constant. ∎
12.4. Sharpness: why one needs verifiable hypotheses
We close with an explicit sharpness statement showing that without a uniform principalization mechanism, no a priori fixed finite testing set can govern all containments in a large class.
Theorem 12.4 (Sharpness via monomial ideals with new facets).
[…]
$J_{N}\ :=\ \overline{(x^{N},\ y^{N^{2}})}\ \subseteq R.$
Let $\mathcal{V}$ be any finite set of real-valued rank-one valuations on $R$ centered at $\mathfrak{m}$. Then, there exist integers $N\geq 2$ and $p,q\geq 1$ such that:
  1. (1)
    For every $v\in\mathcal{V}$, one has $v(\mathfrak{m}^{q})\geq v(J_{N}^{p})$.
  2. (2)
    Nevertheless, $\mathfrak{m}^{q}\not\subseteq J_{N}^{p}=\overline{J_{N}^{p}}$.
In particular, no single finite set of valuations can serve as a universal testing set for all containments of the form $\mathfrak{m}^{q}\subseteq\overline{J_{N}^{p}}$ as $N$ varies.
Proof.
Fix a finite set $\mathcal{V}$ as in the statement. For each $v\in\mathcal{V}$, set $a_{v}:=v(x)>0$ and $b_{v}:=v(y)>0$.
Step 1: choose $N$ so that $v(J_{N})=Na_{v}$ for all $v\in\mathcal{V}$. For $J_{N}=\overline{(x^{N},y^{N^{2}})}$, the valuative criterion for integral closure gives
$v(J_{N})\ =\ \min\{Na_{v},\ N^{2}b_{v}\}.$
Choose an integer $N\geq 2$ such that
$N\ >\ 1+\max_{v\in\mathcal{V}}\frac{a_{v}}{b_{v}}.$
Then, $Nb_{v}>a_{v}$ for all $v\in\mathcal{V}$, hence $N^{2}b_{v}>Na_{v}$ and, therefore,
$v(J_{N})=Na_{v}\quad\text{for all }v\in\mathcal{V}.$
Step 2: choose $q$ so that $v(\mathfrak{m}^{q})\geq v(J_{N}^{p})$ for all $v\in\mathcal{V}$. Set
$M\ :=\ \max_{v\in\mathcal{V}}\left\lceil\frac{a_{v}}{\min\{a_{v},b_{v}\}}\right\rceil.$
Then, $M\geq 1$, and by the choice of $N$ we have $\frac{a_{v}}{b_{v}}<N-1$ for all $v$, and so
$\frac{a_{v}}{\min\{a_{v},b_{v}\}}\ \leq\ \max\left\{1,\frac{a_{v}}{b_{v}}\right\}\ <\ N-1,$
so in particular $M\leq N-1$.
Now choose any integer $p\geq 1$ and set q:=pNM.q\ :=\ pNM. Since $v(\mathfrak{m})=\min\{a_{v},b_{v}\}$ and $v(J_{N})=Na_{v}$ for all $v\in\mathcal{V}$, we obtain
$v(\mathfrak{m}^{q})\ =\ q\,\min\{a_{v},b_{v}\}\ \geq\ pN\cdot\frac{a_{v}}{\min\{a_{v},b_{v}\}}\cdot\min\{a_{v},b_{v}\}\ =\ pNa_{v}\ =\ v(J_{N}^{p})$
for every $v\in\mathcal{V}$. This proves (1).
Step 3: explicit witness that $\mathfrak{m}^{q}\not\subseteq J_{N}^{p}$. We claim that $y^{q}\in\mathfrak{m}^{q}$ but $y^{q}\notin J_{N}^{p}$. The first inclusion is clear.
For the second, note that $J_{N}$ is a monomial integrally closed ideal, and membership in $J_{N}^{p}$ is determined by its Newton polyhedron. The Newton polyhedron of $J_{N}$ is given by
$\frac{u}{N}+\frac{v}{N^{2}}\ \geq\ 1,$
hence that of $J_{N}^{p}$ is
$\frac{u}{N}+\frac{v}{N^{2}}\ \geq\ p.$
For $y^{q}$, the exponent vector is $(u,v)=(0,q)$, so
$\frac{u}{N}+\frac{v}{N^{2}}\ =\ \frac{q}{N^{2}}\ =\ \frac{pNM}{N^{2}}\ =\ p\cdot\frac{M}{N}.$
Since $M\leq N-1$, we have $\frac{M}{N}<1$, and so $\frac{q}{N^{2}}<p$. Therefore, $y^{q}\notin J_{N}^{p}$. Consequently, $\mathfrak{m}^{q}\not\subseteq J_{N}^{p}$, proving (2).
This shows that no fixed finite set of valuations can test all containments of the form $\mathfrak{m}^{q}\subseteq J_{N}^{p}=\overline{J_{N}^{p}}$ uniformly as $N$ varies. ∎
Part VIII Outlook and the Appendix
In this final part, we collect open questions and directions suggested by the valuative framework developed in this paper.
13. Concluding remarks and open problems
This paper develops a valuative framework for understanding the Łojasiewicz exponent, with a focus on finite–max mechanisms. At the same time, Section 12 shows that finite testing cannot be expected in full generality: there exist natural filtrations for which no finite set of valuations suffices to verify containment. This highlights a fundamental dichotomy between situations where the normalized Rees algebra provides complete control and those where the valuative supremum remains intrinsically infinite.
Finite testing and birational models. A central question is to identify intrinsic conditions guaranteeing finite verification.
Question 13.1.
Which properties of a graded family or filtration ensure that integral–closure containments can be verified on a fixed birational model? In particular, can finite testing be detected purely from the structure of the normalized Rees algebra?
Answering this would clarify when the Łojasiewicz exponent is governed by finitely many divisorial valuations and when more subtle valuation-theoretic phenomena occur.
Extremal valuations and geometric structure. When finite testing holds, the valuations computing the exponent play a distinguished role.
Question 13.2.
What geometric or intrinsic properties characterize valuations that compute the Łojasiewicz exponent? For instance, are such valuations uniquely determined by the normalized Rees algebra, or do they satisfy stability or minimality properties within the valuation space?
Understanding these extremal valuations would further clarify the stratification and rigidity phenomena described in Part V.
Variation in families. Part V discusses rigidity and stratification results for the Łojasiewicz exponents of families.
Question 13.3.
More generally, under what conditions is the Łojasiewicz exponent constructible or semicontinuous in algebraic or analytic families? Can weaker hypotheses than finite testing still ensure partial rigidity or controlled variation?
Such results would contribute to a broader algebraic understanding of equisingularity and asymptotic invariants in families.
Role of non–divisorial valuations. While divisorial valuations suffice in many situations, Section 12 shows that more general valuations may be needed in principle.
Question 13.4.
Are there natural geometric or algebraic settings in which non–divisorial valuations necessarily compute the Łojasiewicz exponent? Conversely, what conditions ensure that divisorial valuations always suffice?
Clarifying the role of non–divisorial valuations would help delineate the limits of birational and Rees–algebraic methods.
These questions suggest that the valuative and Rees–algebraic framework developed here provides not only computational tools, but also a structural foundation for studying asymptotic containment invariants. We expect that further development of this perspective will lead to new insights into singularity theory, equisingularity, and related asymptotic invariants.
Appendix
A.1 Detailed proof of Lemma 4.4, following [36]
Proof of Lemma 4.4.
Set $I=\mathfrak{a}^{p}$. We use the following one–function statement, which is a specialization of [36, Théorème 7.2] (equivalence of conditions (1),(5),(6)):
(LT–S) For $h\in R$,
$h\in\overline{I}\quad\Longleftrightarrow\quad\exists U\ni 0,\exists C>0:\ |h(x)|\leq C\|f(x)\|^{p}\text{ on }U.$
For completeness, we reproduce the proof of (LT–S). Let $\pi:Y\to U_{0}$ be the normalized blow–up of $I$ over a sufficiently small neighborhood $U_{0}$ of 0. Then, $Y$ is normal, $\pi$ is proper, and $I\mathcal{O}_{Y}$ is invertible. Hence, for every $y\in\pi^{-1}(0)$ there exist a neighborhood $V\ni y$ and a holomorphic function $s$ on $V$ such that
$I\mathcal{O}_{Y}(V)=(s).$
Fix such $V$ and choose $V^{\prime}\subseteq V$.
Step 1: Two–sided comparison. Since $(f_{i}\circ\pi)^{p}\in(s)$, there exist holomorphic functions $u_{i}$ on $V$ such that
$(f_{i}\circ\pi)^{p}=su_{i}.$
Let $A=\max_{i}\sup_{V^{\prime}}|u_{i}|$. Then, for $z\in V^{\prime}$,
$\|f(\pi(z))\|^{p}\leq A|s(z)|.$
Since $I=\mathfrak{a}^{p}$ is generated by finitely many degree – $p$ monomials $F_{\nu}=f_{1}^{e_{1,\nu}}\cdots f_{r}^{e_{r,\nu}}$, their pullbacks generate $(s)$. Hence, there exist holomorphic $v_{\nu}$ on $V$ such that
$s=\sum_{\nu}v_{\nu}(F_{\nu}\circ\pi).$
Let $B=\sum_{\nu}\sup_{V^{\prime}}|v_{\nu}|$. Then, for $z\in V^{\prime}$,
$|s(z)|\leq B\|f(\pi(z))\|^{p}.$
Thus, on $V^{\prime}$,
$A^{-1}\|f\circ\pi\|^{p}\leq|s|\leq B\|f\circ\pi\|^{p}.$
Step 2: $(LT\text{--}S)$, first direction. Assume $h\in\overline{I}$. By [36, Théorème 7.2], there exists a neighborhood $\tilde{W}$ of $\pi^{-1}(0)$ such that
$h\circ\pi\in I\mathcal{O}_{Y}\quad\text{on }\tilde{W}$.
On $V\subseteq\tilde{W}$, write $h\circ\pi=sH$ with $H$ holomorphic. Let $M=\sup_{V^{\prime}}|H|$. Then, for $z\in V^{\prime}$,
$|h(\pi(z))|\leq M|s(z)|\leq MB\|f(\pi(z))\|^{p}.$
Cover $\pi^{-1}(0)$ by finitely many such $V^{\prime}_{k}$ and let $C^{\prime}$ be the maximum of the corresponding constants. There exists a neighborhood $W$ of $\pi^{-1}(0)$ such that
$|h\circ\pi|\leq C^{\prime}\|f\circ\pi\|^{p}\quad\text{on }W.$
Since $\pi$ is proper, there exists a neighborhood $U\ni 0$ with $\pi^{-1}(U)\subseteq W$. For $x\in U$,
$|h(x)|\leq C^{\prime}\|f(x)\|^{p}.$
Step 3: $(LT\text{--}S)$, converse direction. Assume $|h(x)|\leq C\|f(x)\|^{p}$ on a neighborhood $U$ of 0. Pulling back,
$|h\circ\pi|\leq C\|f\circ\pi\|^{p}.$
On $V^{\prime}\cap\pi^{-1}(U)$, using $\|f\circ\pi\|^{p}\leq A|s|$,
$|h\circ\pi|\leq CA|s|.$
Hence, $(h\circ\pi)/s$ is holomorphic and bounded on $(V^{\prime}\cap\pi^{-1}(U))\setminus\{s=0\}$ and extends holomorphically across $\{s=0\}$. Thus, $h\circ\pi\in(s)=I\mathcal{O}_{Y}$ near $\pi^{-1}(0)$. By Lejeune-Jalabert–Teissier, Théorème 7.2, $h\in\overline{I}$.
Step 4: Deduction of the lemma.
$(1)\Rightarrow(2)$. If $\mathfrak{b}^{q}\subseteq\overline{\mathfrak{a}^{p}}$, then for each $j$, $g_{j}^{q}\in\overline{\mathfrak{a}^{p}}$. Applying $(LT\text{--}S)$ to $h=g_{j}^{q}$ yields
$\|g_{j}(x)\|^{q}\leq C_{j}\|f(x)\|^{p}$
on some neighborhood $U_{j}$. Let $U=\cap_{j}U_{j}$ and $C=\max_{j}C_{j}$. Then, for $x\in U$,
$\|g(x)\|^{q}=\max_{j}|g_{j}(x)|^{q}\leq C\|f(x)\|^{p}.$
$(2)\Rightarrow(1)$. Assume $\|g(x)\|^{q}\leq C\|f(x)\|^{p}$ on a neighborhood $U$. For any qq–tuple $(j_{1},\dots,j_{q})$,
$|g_{j_{1}}(x)\cdots g_{j_{q}}(x)|\leq\|g(x)\|^{q}\leq C\|f(x)\|^{p}.$
Each monomial $g_{j_{1}}\cdots g_{j_{q}}$ is a generator of $\mathfrak{b}^{q}$. Applying $(LT\text{--}S)$ to each such monomial gives $\mathfrak{b}^{q}\subseteq\overline{\mathfrak{a}^{p}}$. ∎

### v2

@@
 2020 Mathematics Subject Classification:
-13H15, 32B10, 32S05, 14B05, 58K30
+14B05, 32B10, 32S05, 13H15, 58K30
 Part I Motivations, Goals and Statement of Main Results
@@
 Identifying when this finite–max phenomenon occurs and exploiting structural consequences, including rigidity, stability, stratification and wall-chamber behaviors, is the main objective of this paper.
+Our approach also resonates with the valuation-theoretic framework for multiplier ideals, plurisubharmonic singularities, and interpolation developed by Boucksom–Favre–Jonsson [12], Favre–Jonsson [19, 20, 21], and more recently by Bao–Guan–Mi–Yuan [1, 2, 3].
+While our setting is algebraic and centers on ideal containment, a common underlying theme is that singularity exponents can be described as optimization problems over suitable spaces of valuations.
+In this sense, the present algebraic theory of Łojasiewicz exponents runs parallel to the valuative descriptions of multiplier ideals, analytic singularities, and interpolation phenomena studied in the above works.
 2. Goals and statement of main results
@@
 Proof.
-This is the Lejeune-Jalabert–Teissier analytic criterion for integral closure; see [36, Théorème 7.2].
+This is the Lejeune-Jalabert–Teissier analytic criterion for integral closure; see [40, Théorème 7.2].
 For the reader’s convenience and to fix normalization conventions, we include a self-contained argument in the appendix.
@@
 $|u_{Y}|\leq C|u_{X}|^{p/q}\text{ on }(U_{\xi}\cap\Gamma_{F})\setminus D_{X}.$
-Equivalently, $|u_{Y}|^{q}\leq C^{q}|u_{X}|^{p}$ holds in a punctured neighborhood of $\xi$. By the Lejeune-Jalabert–Teissier criterion, Lemma 4.4, this is equivalent to the containment
-$(u_{Y})^{q}\subseteq\overline{(u_{X})^{p}}\text{ in }\mathcal{O}_{\overline{\Gamma}_{F},\xi}.$
-Since $D_{X}$ and $D_{Y}$ are Cartier divisors, we have $\mathscr{I}_{X,\xi}=(u_{X})$ and $\mathscr{I}_{Y,\xi}=(u_{Y})$. Thus, the above containment is equivalent to
+Equivalently, $|u_{Y}|^{q}\leq C^{q}|u_{X}|^{p}$ holds in a punctured neighborhood of $\xi$.
+By [40, Théorèm 2.1], this is the case if and only if $(u_{Y})^{q}\subseteq\overline{(u_{X})^{p}}$.
+Since $D_{X}$ and $D_{Y}$ are Cartier divisors, we have $\mathscr{I}_{X,\xi}=(u_{X})$ and $\mathscr{I}_{Y,\xi}=(u_{Y})$.
+Thus, the above containment is equivalent to
 $\mathscr{I}_{Y,\xi}^{\,q}\subseteq\overline{\mathscr{I}_{X,\xi}^{\,p}}\text{ in }\mathcal{O}_{\overline{\Gamma}_{F},\xi}.$
@@
 Proof.
-This result is classical and goes back to Rees [45] and Lipman [38]; see also Huneke–Swanson [30, Chapters 6 and 10] for a modern account. ∎
+This result is classical and goes back to Rees [49] and Lipman [42]; see also Huneke–Swanson [34, Chapters 6 and 10] for a modern account. ∎
 Remark 5.3 (Analytic arc criterion).
@@
 $\mathfrak{a}\mathcal{O}_{Y}=\mathcal{O}_{Y}(-A),\quad\mathfrak{b}\mathcal{O}_{Y}=\mathcal{O}_{Y}(-B),$
-with $A,B$ effective Cartier divisors supported on the exceptional locus.
+with $A,B$ effective Cartier divisors on $Y$.
 Then, for all $p,q\geq 1$, the following are equivalent:
@@
     $\operatorname{\mathrm{ord}}_{E}(\mathfrak{b}^{q})\geq\operatorname{\mathrm{ord}}_{E}(\mathfrak{a}^{p})$ for every prime divisor $E$ on $Y$.
-In particular, integral–closure containments are completely determined by the finitely many divisorial valuations $\{\operatorname{\mathrm{ord}}_{E}\}$ arising from the exceptional divisors of $\pi$.
+In particular, integral–closure containments are completely determined by the finitely many divisorial valuations $\{\operatorname{\mathrm{ord}}_{E}\}$ arising from prime divisors in the supports of $A$ and $B$.
 Proof.
@@
 Thus, we have the following standard equality (cf.
-[30, Chapters 10 and 19]):
+[34, Chapters 10 and 19]):
 $\overline{\mathfrak{a}^{p}}=\pi_{*}(\mathfrak{a}^{p}\mathcal{O}_{Y}).$
@@
 Theorem 5.9 (Finite Rees valuations for a Noetherian filtration).
-Let $(R,\mathfrak{m})$ be an excellent Noetherian local domain, and let ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ be a filtration of ideals in $R$.
+Let $(R,\mathfrak{m})$ be an excellent Noetherian local domain, and let ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ be a filtration of $\mathfrak{m}$-primary ideals in $R$.
 Assume that the Rees algebra $\mathcal{R}({\mathfrak{a}_{\bullet}}):=\bigoplus_{p\geq 0}\mathfrak{a}_{p}T^{p}$ is finitely generated as an RR–algebra.
@@
 $\mathcal{V}({\mathfrak{a}_{\bullet}})=\{v_{1},\dots,v_{r}\},$
-arising from prime divisors on $\overline{Y}$ dominating $\operatorname{Spec}R$, such that for every filtration ${\mathfrak{b}_{\bullet}}=\{\mathfrak{b}_{q}\}_{q\geq 1}$ of ideals in $R$ satisfying $v_{i}({\mathfrak{b}_{\bullet}})>0$ for all $1\leq i\leq r$, one has
+arising from prime divisors on $\overline{Y}$ lying over $\mathfrak{m}$, such that for every filtration ${\mathfrak{b}_{\bullet}}=\{\mathfrak{b}_{q}\}_{q\geq 1}$ of ideals in $R$ satisfying $v_{i}({\mathfrak{b}_{\bullet}})>0$ for all $1\leq i\leq r$, one has
 $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})=\max_{1\leq i\leq r}\frac{v_{i}({\mathfrak{a}_{\bullet}})}{v_{i}({\mathfrak{b}_{\bullet}})}.$
@@
 Corollary 5.11 (Finite testing set from finite generation).
-Let $(R,\mathfrak{m})$ be an excellent Noetherian local domain, and let ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ be a filtration whose Rees algebra $\mathcal{R}({\mathfrak{a}_{\bullet}})$ is finitely generated.
+Let $(R,\mathfrak{m})$ be an excellent Noetherian local domain, and let ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ be a filtration of $\mathfrak{m}$-primary ideals whose Rees algebra $\mathcal{R}({\mathfrak{a}_{\bullet}})$ is finitely generated.
 Then, there exists a finite set of divisorial valuations $\mathcal{V}$ with the following property:
@@
 Proof.
-By Theorem 5.9, there exists a finite set of divisorial valuations $\mathcal{V}({\mathfrak{a}_{\bullet}})$ computing $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$ for all filtrations $\mathfrak{b}_{\bullet}$}.
+By the proof of Theorem 5.9, there exists a finite set of divisorial valuations $\mathcal{V}({\mathfrak{a}_{\bullet}})$ computing $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$ for all filtrations $\mathfrak{b}_{\bullet}$}.
 The valuative criterion for integral closure then implies that the same set of valuations detects all containments $\mathfrak{b}_{q}\subseteq\overline{\mathfrak{a}_{p}}$, proving the claim.
@@
 Proof.
-By the discussion following [9, Definition 1.3] (see also [31, Proposition 5.9] and the discussion that followed), the normalized space $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$ is compact in the topology of pointwise convergence on $R$.
-Note that, by [31, Lemma 4.1], $v\mapsto v(\mathfrak{m})$ and $v\mapsto v(\mathfrak{q})$ are continuous on $\operatorname{Val}_{\mathfrak{m}}$. Define
+By the discussion following [12, Definition 1.3] (see also [35, Proposition 5.9] and the discussion that followed), the normalized space $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$ is compact in the topology of pointwise convergence on $R$.
+Note that, by [35, Lemma 4.1], $v\mapsto v(\mathfrak{m})$ and $v\mapsto v(\mathfrak{q})$ are continuous on $\operatorname{Val}_{\mathfrak{m}}$. Define
 $\Phi:$ $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})\longrightarrow\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m}),\quad\Phi(v):=\frac{1}{v(\mathfrak{m})}\,v,\text{ and }$
@@
 Proof.
-We work on the normalized slice $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})=\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})$, so $v(\mathfrak{m})=1$ for all $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$.
-We shall also use notations and terminology from Favre–Jonsson [16] monograph.
-Step 1: a non-divisorial end and divisorial approximants.
-By Favre–Jonsson’s classification of the valuative tree of $R=\mathbb{C}[[x,y]]$ in [16], there exist non-divisorial ends (for instance curve valuations or infinitely singular valuations).
-Fix such an end valuation $\nu\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$.
-Then, $\nu$ is not divisorial, and there exists an increasing sequence of divisorial valuations
-$\nu_{1}\prec\nu_{2}\prec\cdots\prec\nu_{i}\prec\cdots\prec\nu$
-converging to $\nu$ in $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$, where $\prec$ denotes the tree order.
-We use the fact that the skewness function $\alpha(\cdot)$ on the valuative tree is strictly increasing along any nontrivial segment (see [16]).
-Step 2: realizing $\alpha(v\wedge\nu_{i})$ as $v(J_{i})$. For valuations $v,\mu\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$, Favre–Jonsson [16] define the intersection product by
-$v\cdot\mu:=\alpha(v\wedge\mu),$
-and for fixed $\mu$ the associated extremal tree potential is $g_{\mu}(v):=v\cdot\mu=\alpha(v\wedge\mu)$.
-If $\mu$ is divisorial, then it was proved (see the proof of [17, Prop.
-5.15]; although this was written for the ring $\mathbb{C}\{x,y\}$ of convergent formal power series, the argument applies verbatim to the formal power series ring $\mathbb{C}[[x,y]]$) that there exist irreducible elements $\phi_{\mu,1},\phi_{\mu,2}\in\mathfrak{m}$ such that
-$g_{\mu}(v)=\min\{v(\phi_{\mu,1}),v(\phi_{\mu,2})\}\quad\text{for all }v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m}).$
-Applying this to $\mu=\nu_{i}$, we may choose irreducibles $\phi_{i,1},\phi_{i,2}\in\mathfrak{m}$ with
-$\alpha(v\wedge\nu_{i})=\min\{v(\phi_{i,1}),v(\phi_{i,2})\}\quad\forall\,v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m}).$
-Set
-$J_{i}:=(\phi_{i,1},\phi_{i,2})\subseteq R.$
-Since $\phi_{i,1},\phi_{i,2}$ are distinct irreducible germs through the origin, the ideal $J_{i}$ has height $2$ in the regular local ring $R=\mathbb{C}[[x,y]]$ and is therefore $\mathfrak{m}$–primary.
-For all $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$ we have the exact identity
-(5.4) $v(J_{i})=\alpha(v\wedge\nu_{i}).$
-Set $A_{i}:=\alpha(\nu_{i})\in(0,\infty)$.
-Step 3: normalized coefficients and definition of the filtrations. Set
-$c_{i}:=\frac{2^{-i}}{A_{i}}.$
-Since $\alpha(\operatorname{\mathrm{ord}}_{\mathfrak{m}})=1$ on the normalized slice and $\alpha$ is increasing, we have $A_{i}\geq 1$ and so $c_{i}\leq 2^{-i}$.
-Using (5.4), for all $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$,
-$0\leq c_{i}\,v(J_{i})=c_{i}\,\alpha(v\wedge\nu_{i})\leq c_{i}\,\alpha(\nu_{i})=2^{-i}.$
-Define ${\mathfrak{b}_{\bullet}}:=\mathfrak{m}^{\bullet}$, i.e., $\mathfrak{b}_{p}=\mathfrak{m}^{p}$. Define ${\mathfrak{a}_{\bullet}}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ by
-(5.5) $\mathfrak{a}_{p}:=\mathfrak{m}^{p}\cdot\prod_{i\geq 1}J_{i}^{\lfloor p\,c_{i}\rfloor}.$
-Since $c_{i}\leq 2^{-i}$, we have $\lfloor p\,c_{i}\rfloor=0$ for all $i\gg 1$, so the product is finite for each $p$.
-In particular, each $\mathfrak{a}_{p}$ is $\mathfrak{m}$–primary.
-Observe further that, for all $p,p^{\prime}\geq 1$ and all $i$, $\lfloor(p+p^{\prime})c_{i}\rfloor\geq\lfloor pc_{i}\rfloor+\lfloor p^{\prime}c_{i}\rfloor.$ Therefore,
-$\mathfrak{a}_{p}\mathfrak{a}_{p^{\prime}}=\mathfrak{m}^{p+p^{\prime}}\cdot\prod_{i}J_{i}^{\lfloor pc_{i}\rfloor+\lfloor p^{\prime}c_{i}\rfloor}\subseteq\mathfrak{m}^{p+p^{\prime}}\cdot\prod_{i}J_{i}^{\lfloor(p+p^{\prime})c_{i}\rfloor}=\mathfrak{a}_{p+p^{\prime}}.$
-Moreover, $\mathfrak{a}_{p+1}\subseteq\mathfrak{a}_{p}$, so $\mathfrak{a}_{\bullet}$} is a filtration of ideals.
-Step 4: computation of $v({\mathfrak{a}_{\bullet}})$. Fix $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$. By definition, we have
-$v(\mathfrak{a}_{p})=v(\mathfrak{m}^{p})+\sum_{i\geq 1}\lfloor p\,c_{i}\rfloor\,v(J_{i})=p\,v(\mathfrak{m})+\sum_{i\geq 1}\lfloor p\,c_{i}\rfloor\,v(J_{i}).$
-Dividing by $p$,
-$\frac{v(\mathfrak{a}_{p})}{p}=v(\mathfrak{m})+\sum_{i\geq 1}\frac{\lfloor p\,c_{i}\rfloor}{p}\,v(J_{i}).$
-For each $i$, $\frac{\lfloor p\,c_{i}\rfloor}{p}\to c_{i}$ as $p\to\infty$, and
-$0\leq\frac{\lfloor p\,c_{i}\rfloor}{p}\,v(J_{i})\leq c_{i}\,v(J_{i})\leq 2^{-i}.$
-Since $\sum_{i\geq 1}2^{-i}<\infty$, dominated convergence applies and yields
-$v({\mathfrak{a}_{\bullet}})=\lim_{p\to\infty}\frac{v(\mathfrak{a}_{p})}{p}=v(\mathfrak{m})+\sum_{i\geq 1}c_{i}\,v(J_{i})=v(\mathfrak{m})+\sum_{i\geq 1}c_{i}\,\alpha(v\wedge\nu_{i}).$
-Since ${\mathfrak{b}_{\bullet}}=\mathfrak{m}^{\bullet}$ and $v(\mathfrak{m})=1$ on $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$, we have $v({\mathfrak{a}_{\bullet}})/v({\mathfrak{b}_{\bullet}})=v({\mathfrak{a}_{\bullet}})$ on this slice.
-Step 5: unique maximizer and non-attainment by divisorial valuations. Because $\nu\wedge\nu_{i}=\nu_{i}$ for all $i$,
-$\nu({\mathfrak{a}_{\bullet}})=1+\sum_{i\geq 1}c_{i}\,\alpha(\nu_{i})=1+\sum_{i\geq 1}2^{-i}=2.$
-Let $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$ with $v\neq\nu$.
-Since $\nu$ is an end and $\nu_{i}\uparrow\nu$, there exists $i_{0}$ such that $v\not\succeq\nu_{i_{0}}$; otherwise $v\succeq\nu_{i}$ for all $i$ and so $v\succeq\nu$, forcing $v=\nu$.
-For this $i_{0}$ we have $v\wedge\nu_{i_{0}}\prec\nu_{i_{0}}$, and strict monotonicity of $\alpha$ along segments implies
-$\alpha(v\wedge\nu_{i_{0}})<\alpha(\nu_{i_{0}}).$
-For all $i$, $\alpha(v\wedge\nu_{i})\leq\alpha(\nu_{i})$, with strict inequality at $i=i_{0}$. Therefore,
-$v({\mathfrak{a}_{\bullet}})=1+\sum_{i\geq 1}c_{i}\,\alpha(v\wedge\nu_{i})<1+\sum_{i\geq 1}c_{i}\,\alpha(\nu_{i})=2=\nu({\mathfrak{a}_{\bullet}}).$
-Thus, $\nu$ is the unique maximizer of $v({\mathfrak{a}_{\bullet}})$ on $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$ and, hence, of $v({\mathfrak{a}_{\bullet}})/v({\mathfrak{b}_{\bullet}})$.
-Since $\nu$ is not divisorial, no divisorial valuation attains the supremum.
-Step 6: linear boundedness.
-Since ${\mathfrak{b}_{\bullet}}=\mathfrak{m}^{\bullet}$, it is linearly bounded with respect to $\mathfrak{q}=\mathfrak{m}$.
-For $\mathfrak{a}_{\bullet}$}, the definition (5.5) gives $\mathfrak{m}^{p}\subseteq\mathfrak{a}_{p}$ for all $p$.
-Using $J_{i}\subseteq\mathfrak{m}$ and $\sum_{i}c_{i}\leq 1$, we obtain
-$\mathfrak{a}_{p}\subseteq\mathfrak{m}^{p}\cdot\prod_{i}\mathfrak{m}^{\lfloor p\,c_{i}\rfloor}=\mathfrak{m}^{p+\sum_{i}\lfloor p\,c_{i}\rfloor}\subseteq\mathfrak{m}^{p+\lfloor p\sum_{i}c_{i}\rfloor}\subseteq\mathfrak{m}^{2p}.$
-Hence, $\mathfrak{a}_{\bullet}$} is linearly bounded with respect to $\mathfrak{q}=\mathfrak{m}$, and the hypotheses of Theorem 5.14 are satisfied. This completes the proof. ∎
+We work on the normalized slice $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})=\operatorname{Val}_{\mathfrak{m}}(\mathfrak{q})$ so that $v(\mathfrak{m})=1$ for all $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$.
+We use the notation and terminology of Favre–Jonsson monograph [19].
+[…]
+For $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$ set $\beta(v):=\alpha(v\wedge\nu)$.
+Since $\nu_{i}$ lies on the segment $\operatorname{\mathrm{ord}}_{\mathfrak{m}},\nu]$, the skewness parameterization of segments ([19, Thm.
+3.26]) gives
+$\alpha(v\wedge\nu_{i})=\min\{\beta(v),A_{i}\}.$
+Step 2: simple complete ideals attached to the $\nu_{i}$.
+For each $i$, let $b_{i}:=b(\nu_{i})$ be the generic multiplicity of the divisorial valuation $\nu_{i}$.
+Consider the positive atomic measure $\rho_{i}:=b_{i}\nu_{i}$ and define
+$J_{i}:=I_{\rho_{i}}=\{\phi\in R\mid g_{\phi}\geq g_{\rho_{i}}\},$
+where $I_{\rho}$ is as in [19, (8.3)]. By [19, Thm. 8.12], $J_{i}$ is an integrally closed ideal and its tree measure satisfies
+$\rho_{J_{i}}=\rho_{i}.$
+Together with [19, Prop.
+8.9] and the fact that the maps $I\mapsto\rho_{I}$ and $\rho\mapsto I_{\rho}$ are inverse semigroup isomorphisms between integrally closed ideals and measures in $\mathcal{M}_{\mathcal{I}}^{+}$ ([19, Thm.
+8.12]), this also implies that $g_{J_{i}}=g_{\rho_{i}}$ as the corresponding positive tree potential.
+In particular, by [19, Thm.
+7.48 and (8.1)], for all $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$,
+$v(J_{i})=g_{J_{i}}(v)=g_{\rho_{i}}(v)=\int\alpha(\mu\wedge v)\,d\rho_{i}(\mu)=b_{i}\alpha(v\wedge\nu_{i}).$
+Furthermore, by the unique factorization of $J_{i}$, as described in [19, Thm.
+8.12], and the fact that $\rho_{J_{i}}=b_{i}\nu_{i}$, it follows that this factorization of $J_{i}$ has only one divisorial atom and no curve atoms.
+Therefore, $J_{i}$ is the simple complete ideal attached to the divisorial valuation $\nu_{i}$ and, in particular, is $\mathfrak{m}$-primary.
+Particularly, there exists $\ell_{i}\geq 1$ such that
+$\mathfrak{m}^{\ell_{i}}\subseteq J_{i}.$
+Step 3: integer weights. Choose integers $n_{i}\geq 1$ inductively so that $D_{i}:=\lfloor n_{i}b_{i}A_{i}\rfloor$ is strictly increasing and
+$c_{i}:=\frac{n_{i}b_{i}A_{i}}{D_{i}}\longrightarrow 1.$
+Set $s_{i}(v):=v(J_{i}^{n_{i}})/D_{i}$. Using the previous formula for $v(J_{i})$,
+$s_{i}(v)=\frac{n_{i}b_{i}}{D_{i}}\,\alpha(v\wedge\nu_{i})=\frac{n_{i}b_{i}}{D_{i}}\,\min\{\beta(v),A_{i}\}=c_{i}\frac{\min\{\beta(v),A_{i}\}}{A_{i}}.$
+We now prove that $s_{i}(v)\to g(v):=\frac{\beta(v)}{A}=\frac{\alpha(v\wedge\nu)}{A}$ uniformly on $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$.
+Indeed, for any $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$, we have
+$\left|s_{i}(v)-\frac{\beta(v)}{A}\right|\leq|c_{i}-1|\frac{\min\{\beta(v),A_{i}\}}{A_{i}}+\left|\frac{\min\{\beta(v),A_{i}\}}{A_{i}}-\frac{\beta(v)}{A}\right|.$
+Since $\frac{\min\{\beta(v),A_{i}\}}{A_{i}}\leq 1$ and $0\leq\frac{\min\{\beta(v),A_{i}\}}{A_{i}}-\frac{\beta(v)}{A}\leq 1-\frac{A_{i}}{A},$ we obtain
+$\sup_{v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})}\left|s_{i}(v)-g(v)\right|\leq|c_{i}-1|+\left(1-\frac{A_{i}}{A}\right).$
+Because $c_{i}\to 1$ and $A_{i}\uparrow A$, the right-hand side tends to 0, so $s_{i}(v)\to g(v)$ uniformly on $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$.
+Step 4: the auxiliary filtration $\mathfrak{c}_{\bullet}$. For $p\geq 1$ define
+$\mathfrak{c}_{p}:=\sum_{\sum_{i}m_{i}D_{i}\geq p}\prod_{i}J_{i}^{n_{i}m_{i}},$
+the sum taken over finitely supported tuples $(m_{i})$. It can be seen that $\mathfrak{c}_{\bullet}$ is a filtration of ideals.
+Set ${\mathfrak{b}_{\bullet}}:=\mathfrak{m}^{\bullet}$ and $\mathfrak{a}_{p}:=\mathfrak{m}^{p}\mathfrak{c}_{p}$. Then, $\mathfrak{a}_{\bullet}$} is also a filtration of ideals.
+Step 5: computation of $v(\mathfrak{c}_{\bullet})$. For any monomial summand $M=\prod_{i}J_{i}^{n_{i}m_{i}}$ with $\sum_{i}m_{i}D_{i}\geq p$,
+$v(M)=\sum_{i}m_{i}D_{i}\,s_{i}(v)\geq p\,\inf_{i}s_{i}(v).$
+Conversely $J_{i}^{n_{i}\lceil p/D_{i}\rceil}$ occurs among the summands of $\mathfrak{c}_{p}$, so $v(\mathfrak{c}_{p})\leq\lceil p/D_{i}\rceil v(J_{i}^{n_{i}})$. We shall show that
+(5.4) $s_{i}(v)\geq g(v)\text{ for all }i\text{ and all }v.$
+Indeed, observe that $s_{i}(v)\geq\frac{\min\{\beta(v),A_{i}\}}{A_{i}}.$ If $\beta(v)\leq A_{i}$, then
+$\frac{\min\{\beta(v),A_{i}\}}{A_{i}}=\frac{\beta(v)}{A_{i}}\geq\frac{\beta(v)}{A}=g(v),$
+whereas if $\beta(v)\geq A_{i}$, then
+$\frac{\min\{\beta(v),A_{i}\}}{A_{i}}=1\geq\frac{\beta(v)}{A}=g(v).$
+This proves (5.4).
+Next, as shown above, $s_{i}(v)\longrightarrow g(v)$ uniformly on $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$. Therefore, together with (5.4), we have
+$\inf_{i}s_{i}(v)=g(v)\text{ for all }v.$
+Now, it can be seen that $v(\mathfrak{c}_{p})\geq pg(v)$.
+On the other hand, for each fixed $i$, $J_{i}^{n_{i}\lceil p/D_{i}\rceil}\subseteq\mathfrak{c}_{p},$ so $v(\mathfrak{c}_{p})\leq\left\lceil\frac{p}{D_{i}}\right\rceil v(J_{i}^{n_{i}}),$ and therefore,
+$\frac{v(\mathfrak{c}_{p})}{p}\leq s_{i}(v)+\frac{n_{i}b_{i}A_{i}}{p}.$
+We shall prove
+$\lim_{p\to\infty}\frac{v(\mathfrak{c}_{p})}{p}=g(v)$
+uniformly on $\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$. Let $\varepsilon>0$. By the uniform convergence of $s_{i}$ to $g$ proved before, choose $i_{0}$ such that
+$\sup_{v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})}|s_{i_{0}}(v)-g(v)|<\varepsilon/2.$
+Since $v(J_{i_{0}}^{n_{i_{0}}})\leq n_{i_{0}}b_{i_{0}}A_{i_{0}}$ for all $v$, choose $p_{0}$ such that
+$\frac{n_{i_{0}}b_{i_{0}}A_{i_{0}}}{p}<\varepsilon/2\text{ for all }p\geq p_{0}.$
+For such $p$, the lower bound above gives
+$\frac{v(\mathfrak{c}_{p})}{p}\geq g(v),$
+while the upper bound with $i=i_{0}$ gives
+$\frac{v(\mathfrak{c}_{p})}{p}\leq s_{i_{0}}(v)+\frac{n_{i_{0}}b_{i_{0}}A_{i_{0}}}{p}\leq g(v)+\varepsilon$
+for all $v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})$. Thus,
+$\sup_{v\in\operatorname{Val}_{\mathfrak{m}}(\mathfrak{m})}\left|\frac{v(\mathfrak{c}_{p})}{p}-g(v)\right|\leq\varepsilon$
+for all $p\geq p_{0}$, establishing the desired uniform convergence.
+Step 6: valuation of $\mathfrak{a}_{\bullet}$}. Since $\mathfrak{a}_{p}=\mathfrak{m}^{p}\mathfrak{c}_{p}$ and $v(\mathfrak{m})=1$,
+$\frac{v(\mathfrak{a}_{p})}{p}\to 1+g(v)=1+\frac{\alpha(v\wedge\nu)}{A}.$
+Hence, $v({\mathfrak{a}_{\bullet}})=1+\alpha(v\wedge\nu)/A$ and $v({\mathfrak{b}_{\bullet}})=1$.
+Step 7: maximizer.
+At $v=\nu$ this ratio equals $2$.
+If $v\neq\nu$, then $v\wedge\nu\prec\nu$ and by strict monotonicity of skewness along segments ([19, Thm.
+3.26]), $\alpha(v\wedge\nu)<A$.
+It follows that $v({\mathfrak{a}_{\bullet}})/v({\mathfrak{b}_{\bullet}})<2=\nu({\mathfrak{a}_{\bullet}})/\nu({\mathfrak{b}_{\bullet}})$.
+Thus, $\nu$ uniquely maximizes $v({\mathfrak{a}_{\bullet}})/v({\mathfrak{b}_{\bullet}})$, and since $\nu$ is not divisorial, no divisorial valuation attains the supremum.
+Step 8: linear boundedness.
+Because $\mathfrak{c}_{p}\subseteq R$, one has $\mathfrak{a}_{p}\subseteq\mathfrak{m}^{p}$.
+Also $J_{1}^{n_{1}\lceil p/D_{1}\rceil}\subseteq\mathfrak{c}_{p}$ and $\mathfrak{m}^{\ell_{1}}\subseteq J_{1}$, so $\mathfrak{m}^{p+n_{1}\ell_{1}\lceil p/D_{1}\rceil}\subseteq\mathfrak{a}_{p}$.
+Thus $\mathfrak{m}^{\lceil cp\rceil}\subseteq\mathfrak{a}_{p}\subseteq\mathfrak{m}^{p}$ for some $c>0$, and $\mathfrak{a}_{\bullet}$} is linearly bounded with respect to $\mathfrak{q}=\mathfrak{m}$.
+∎
 5.6. Finite maximizers in the ideal case
@@
 Specializing to ${\mathfrak{b}_{\bullet}}=\{\mathfrak{b}^{q}\}_{q\geq 1}$, noticing that $\mathfrak{b}\subseteq\mathfrak{m}$ so $\mathfrak{b}_{\bullet}$} is linearly bounded from below (as in Remark 5.10), yields the desired finite–maximum formula.
-Finally, the valuations coming from the codimension–one components of the exceptional locus of the normalized blowup $\operatorname{Proj}\overline{\mathcal{R}(\mathfrak{a})}\to\operatorname{Spec}R$ are precisely the classical Rees valuations of $\mathfrak{a}$; see [30, Chapter 10].
+Finally, the valuations coming from the codimension–one components of the exceptional locus of the normalized blowup $\operatorname{Proj}\overline{\mathcal{R}(\mathfrak{a})}\to\operatorname{Spec}R$ are precisely the classical Rees valuations of $\mathfrak{a}$; see [34, Chapter 10].
 Therefore, the maximum above may be taken over the Rees valuations of $\mathfrak{a}$, as claimed.
@@
 For $1\leq i\leq d$, set
-$e_{i}({\mathfrak{a}_{\bullet}}^{[i]},{\mathfrak{b}_{\bullet}}^{[d-i]})=\limsup_{p\rightarrow\infty}\frac{e(\mathfrak{a}_{p}^{[i]},\mathfrak{b}_{p}^{[d-i]})}{n^{d}}.$
+$e_{i}({\mathfrak{a}_{\bullet}}^{[i]},{\mathfrak{b}_{\bullet}}^{[d-i]})=\limsup_{p\rightarrow\infty}\frac{e(\mathfrak{a}_{p}^{[i]},\mathfrak{b}_{p}^{[d-i]})}{p^{d}}.$
 It was shown by Cid-Ruiz–Montaño [13] that, under the condition that $\dim\text{Nil}(\widehat{R})<d$, $e_{i}({\mathfrak{a}_{\bullet}}^{[i]},{\mathfrak{b}_{\bullet}}^{[d-i]})$ exists as an actual limit.
@@
 […]
-It follows that $\Lambda_{i}(\mathfrak{a}_{pt})\geq\Lambda_{i+1}(\mathfrak{a}_{pt})$.
-Dividing both sides of this inequality by appropriate powers of $pt$ and taking the limit along $p\to\infty$ then give $\Lambda_{1}({\mathfrak{a}_{\bullet}})\geq\Lambda_{2}({\mathfrak{a}_{\bullet}})\geq\cdots\geq\Lambda_{d}({\mathfrak{a}_{\bullet}})$ as claimed.
+It follows that Dividing both sides of this inequality by appropriate powers of $pt$ and taking the limit along $p\to\infty$ then give $\Lambda_{1}({\mathfrak{a}_{\bullet}})\geq\Lambda_{2}({\mathfrak{a}_{\bullet}})\geq\cdots\geq\Lambda_{d}({\mathfrak{a}_{\bullet}})$ as claimed.
 ∎
@@
 It was observed in [35] that, in general, there may not exist a divisorial valuation $v$ realizing $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})$.
-On the other hand, it was also shown in [31, Theorem A], that given any graded family $\mathfrak{a}_{\bullet}$}, there is a real valuation $v$ of $R$ centered at $\mathfrak{m}$ that computes $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})$, i.e., $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})=\frac{A(v)}{v({\mathfrak{a}_{\bullet}})}$.
+On the other hand, it was also shown in [35, Theorem A], that given any graded family $\mathfrak{a}_{\bullet}$}, there is a real valuation $v$ of $R$ centered at $\mathfrak{m}$ that computes $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})$, i.e., $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})=\frac{A(v)}{v({\mathfrak{a}_{\bullet}})}$.
 7.2.
@@
 When ${\mathfrak{b}_{\bullet}}=\{\mathfrak{m}^{q}\}_{q\geq 1}$, we have $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})\cdot\mathcal{L}_{\mathfrak{m}}({\mathfrak{a}_{\bullet}})\geq\operatorname{\mathrm{lct}}(\mathfrak{m})$.
-On the other hand, since $R$ is regular of dimension $d$, one has $\operatorname{\mathrm{lct}}(\mathfrak{m})=d$; see, for instance [35, § 9].
+On the other hand, since $R$ is regular of dimension $d$, one has $\operatorname{\mathrm{lct}}(\mathfrak{m})=d$; see, for instance [39, § 9].
 Thus, $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})\cdot\mathcal{L}_{\mathfrak{m}}({\mathfrak{a}_{\bullet}})\geq d$.
@@
   1. (i)
-    ([28, Theorem 2.1]) Integral closure is compatible with specialization by a generic element of $I$.
+    ([32, Theorem 2.1]) Integral closure is compatible with specialization by a generic element of $I$.
   2. (ii)
-    ([26, Proposition 3.7]) For a general element $x\in I$, one has
+    ([30, Proposition 3.7]) For a general element $x\in I$, one has
     $\overline{I^{n}}\,R/(x)=\overline{I^{n}R/(x)}\quad\text{for all }n\gg 0.$
@@
 Theorem 7.4 (Stable–range commutation for $\overline{\mathfrak{a}_{pt}}$).
-Assume that $R$ satisfies the hypotheses of Remark 7.3.
+Assume that $R$ satisfies the hypotheses of Remark 7.3 and $\dim R\geq 2$.
 Let $\mathfrak{a}_{\bullet}$} be an $\mathfrak{m}$–primary filtration that is eventually degree-one normal, and set $I=\mathfrak{a}_{1}$.
@@
 $I^{pt}\subseteq\mathfrak{a}_{pt}\subseteq\overline{I^{pt}}.$
-Let $x\in I$ be general as in Remark 7.3(ii) and set $R^{\prime}=R/(x)$. By [26, Proposition 3.7], there exists $n_{0}$ such that
+Let $x\in I$ be general as in Remark 7.3(ii) and set $R^{\prime}=R/(x)$. By [30, Proposition 3.7], there exists $n_{0}$ such that
 $\overline{I^{n}}\,R^{\prime}=\overline{I^{n}R^{\prime}}\quad\text{for all }n\geq n_{0}.$
@@
 Assume $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})\cdot\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})=\operatorname{\mathrm{lct}}({\mathfrak{b}_{\bullet}})$.
-Choose a real valuation $v_{0}$ centered at $\mathfrak{m}$ that computes $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})$, whose existence was established in [31, Theorem A].
+Choose a real valuation $v_{0}$ centered at $\mathfrak{m}$ that computes $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})$, whose existence was established in [35, Theorem A].
 That is, $\operatorname{\mathrm{lct}}({\mathfrak{a}_{\bullet}})=A(v_{0})/v_{0}({\mathfrak{a}_{\bullet}})$.
@@
 By Theorem 5.17, $\mathcal{L}_{\mathfrak{m}}(I)$ is computed by finitely many Rees valuations of $I$.
-Since $I$ is monomial, the Rees valuations of $I$ are precisely the monomial valuations obtained from the (non-redundant) bounding hyperplanes of the Newton polyhedron $\operatorname{NP}(I)$; see [30, Theorem 10.3.5].
+Since $I$ is monomial, the Rees valuations of $I$ are precisely the monomial valuations obtained from the (non-redundant) bounding hyperplanes of the Newton polyhedron $\operatorname{NP}(I)$; see [34, Theorem 10.3.5].
 For an $\mathfrak{m}$–primary monomial ideal, the relevant bounding hyperplanes are exactly those defining the compact facets of $\operatorname{NP}(I)$, hence they are indexed by the primitive inward normal vectors $u\in\mathcal{U}(I)$, and the corresponding valuations are the $v_{u}$.
@@
 Step 2: Inequality and equality. The inequality $\Theta(I)\geq 1$ follows from Theorem 7.1.
-If $\Theta(I)=1$, then $\operatorname{\mathrm{lct}}(I)\cdot\mathcal{L}_{\mathfrak{m}}(I)=d=\operatorname{\mathrm{lct}}(\mathfrak{m})$.
-By Theorem 8.1, equality forces a common valuation computing both $\operatorname{\mathrm{lct}}(I)$ and $\mathcal{L}_{\mathfrak{m}}(I)$ and also satisfying $A(v)/v(\mathfrak{m})=d$.
-In the monomial category, such a valuation must be $v_{u}$ for some $u\in\mathbb{Z}_{>0}^{d}$, and then
+Consider the equality $\Theta(I)=1$, i.e., $\operatorname{\mathrm{lct}}(I)\cdot\mathcal{L}_{\mathfrak{m}}(I)=d=\operatorname{\mathrm{lct}}(\mathfrak{m})$.
+Let $v_{u}$, for some $u\in\mathbb{Z}^{d}_{>0}$, be a monomial valuation that computes $\operatorname{\mathrm{lct}}(I)$ as in Step 1.
+Then, since $\mathcal{L}_{\mathfrak{m}}(I)\geq\frac{v_{u}(I)}{v_{u}(\mathfrak{m})}$ and $\frac{A(v_{u})}{v_{u}(\mathfrak{m})}\geq\operatorname{\mathrm{lct}}(\mathfrak{m})$, the equality $\Theta(I)=1$ forces $v_{u}$ to compute both $\mathcal{L}_{\mathfrak{m}}(I)$ and $\operatorname{\mathrm{lct}}(\mathfrak{m})$.
+Thus,
 $\frac{A(v_{u})}{v_{u}(\mathfrak{m})}=\frac{\sum_{i}u_{i}}{\min_{i}u_{i}}=d\quad\Longleftrightarrow\quad u_{1}=\cdots=u_{d}.$
@@
 $\frac{v_{u_{1}}(\mathfrak{a})}{v_{u_{1}}({\mathfrak{b}_{\bullet}}(t))}=\frac{12}{7+3t},\quad\frac{v_{u_{2}}(\mathfrak{a})}{v_{u_{2}}({\mathfrak{b}_{\bullet}}(t))}=\frac{5}{3+t}.$
-These two ratios are equal precisely at $t=\frac{1}{3}$. Consequently,
+These two ratios are equal precisely at $t=1\frac{1}{3}$. Consequently,
 $\operatorname{argmax}(\mathfrak{a},{\mathfrak{b}_{\bullet}}(t))=\begin{cases}\{u_{1}\},&t<\frac{1}{3},\\[4.0pt] \{u_{1},u_{2}\},&t=\frac{1}{3},\\[4.0pt] \{u_{2}\},&t>\frac{1}{3}.\end{cases}$
@@
 Let $T$ be a Noetherian scheme, let $X$ be a Noetherian integral scheme, and set $X_{T}:=X\times T$.
-Let $\pi:Y\to X_{T}$ be a proper birational morphism with $Y$ normal, let $E\subseteq Y$ be a prime divisor whose image dominates $T$, and let $\mathcal{I}\subseteq\mathcal{O}_{X_{T}}$ be a coherent ideal sheaf, and for $t\in T$ let $\mathcal{I}_{t}$ be its fiber on $X\times\{t\}$.
+Let $\pi:Y\to X_{T}$ be a proper birational morphism with $Y$ normal and $Y\to T$ a flat family, and let $E\subseteq Y$ be an effective Cartier divisor that dominates $T$.
+Let $\mathcal{I}\subseteq\mathcal{O}_{X_{T}}$ be a coherent ideal sheaf, and for $t\in T$ let $\mathcal{I}_{t}$ be its fiber on $X\times\{t\}$.
 Define
@@
 In particular, $t\mapsto 1/L(t)$ is fractional linear (hence $C^{\infty}$) on each such $I$, and its second derivative has constant sign on $I$.
-Moreover, on a given component $I$ the restriction $t\mapsto 1/L(t)$ is affine if and only if the affine functions $\alpha_{j}$ and $\beta_{j}$ are proportional on $I$.
+Moreover, on a given component $I$ the restriction $t\mapsto 1/L(t)$ is affine if and only if either $\alpha_{j}$ is constant on $I$ or the affine functions $\alpha_{j}$ and $\beta_{j}$ are proportional on $I$.
 Proof.
@@
 whose sign is constant on $I$ because the denominator is positive and the numerator is constant.
-Finally, $\beta_{j}/\alpha_{j}$ is affine on $I$ if and only if $b_{1}a_{0}-b_{0}a_{1}=0$, i.e.
-if and only if $\alpha_{j}$ and $\beta_{j}$ are proportional affine functions on $I$.
+Finally, $\beta_{j}/\alpha_{j}$ is affine on $I$ if and only if $a_{1}(b_{1}a_{0}-b_{0}a_{1})=0$, i.e.
+if and only if either $\alpha_{j}$ is constant or $\alpha_{j}$ and $\beta_{j}$ are proportional affine functions on $I$.
 ∎
@@
 […]
-Finally, $\operatorname{NP}(I)$ is a rational polyhedron, so it is the intersection of its supporting halfspaces defined by the primitive inward normals of its compact faces.
+Finally, $\operatorname{NP}(I)$ is a rational polyhedron, so it is the intersection of its supporting halfspaces defined by the primitive inward normals of its compact facets.
 Therefore, the above inequalities are equivalent to $m\in\operatorname{NP}(I)\cap S$.
@@
 ∎
-Note that, in Theorem 10.11, only compact facets contribute since $\mathfrak{a}$ being $\mathfrak{m}_{S}$–primary implies $h_{\operatorname{NP}(\mathfrak{a})}(u)>0$ precisely along inward normals to compact faces.
 10.7. Monomial filtrations in the toric setting
@@
 Lemma 10.15 (Finite–max on a finite fan).
-Let $\sigma\subseteq N_{\mathbb{R}}$ be a strongly convex rational polyhedral cone and set $\sigma^{\circ}$ for its interior.
-Let $\Sigma$ be a finite rational polyhedral fan supported on $\sigma$ with rays $\rho_{1},\dots,\rho_{M}$ and primitive generators $u_{1},\dots,u_{M}\in N$.
-Let $\alpha,\beta:\sigma^{\circ}\to\mathbb{R}_{>0}$ be functions such that:
+Let $\alpha,\beta:\sigma\to\mathbb{R}_{\geq 0}$ be continuous functions that are positively homogeneous of degree $1$. Assume that:
   1. (1)
-    $\alpha$ and $\beta$ extend to continuous functions on $\sigma^{\circ}$ and are positive on $\sigma^{\circ}$;
+    $\beta(u)>0$ for all $u\in\sigma^{\circ}$;
   2. (2)
-    for every cone $\tau\in\Sigma$, the restrictions $\alpha|_{\tau^{\circ}}$ and $\beta|_{\tau^{\circ}}$ are linear forms in $u$.
-Then,
-$\sup_{u\in\sigma^{\circ}\cap N}\frac{\alpha(u)}{\beta(u)}\;=\;\max_{1\leq j\leq M}\frac{\alpha(u_{j})}{\beta(u_{j})}.$
-In particular, the supremum is a maximum attained at a primitive ray generator of $\Sigma$.
+    there exists a finite rational polyhedral fan $\Sigma$ supported on $\sigma$ such that $\alpha$ and $\beta$ are linear on every cone $\tau\in\Sigma$.
+Let $\rho_{1},\dots,\rho_{M}$ be the rays of $\Sigma$ and $u_{1},\dots,u_{M}\in N$ their primitive generators. Then,
+$\sup_{u\in\sigma^{\circ}\cap N}\frac{\alpha(u)}{\beta(u)}=\sup_{u\in\sigma^{\circ}}\frac{\alpha(u)}{\beta(u)}=\max_{\{j:\beta(u_{j})>0\}}\frac{\alpha(u_{j})}{\beta(u_{j})}.$
 Proof.
-Since $\beta(u)>0$ on $\sigma^{\circ}$, we may normalize on the compact slice
-$K:=\{u\in\sigma^{\circ}:\beta(u)=1\}.$
-On each cone $\tau\in\Sigma$, the set $K\cap\tau$ is a compact polytope and $\alpha/\beta=\alpha$ on $K$ is linear on $\tau$ by hypothesis.
-Hence, the maximum of $\alpha$ on $K\cap\tau$ is attained at a vertex of that polytope, which lies on an extreme ray of $\tau$.
-Taking the maximum over finitely many cones yields that the global supremum on $K$ is attained at one of the finitely many ray generators $u_{1},\dots,u_{M}$ of $\Sigma$, giving the claimed formula.
+Fix a cone $\tau\in\Sigma$, and let $w_{1},\dots,w_{s}$ be the primitive generators of its rays.
+Every $u\in\tau$ can be written as $u=\sum_{i}\lambda_{i}w_{i}$ with $\lambda_{i}\geq 0$.
+Since $\alpha$ and $\beta$ are linear on $\tau$, we have $\alpha(u)=\sum_{i}\lambda_{i}\alpha(w_{i})$ and $\beta(u)=\sum_{i}\lambda_{i}\beta(w_{i})$.
+If $u\in\tau^{\circ}$, then $\beta(u)>0$ by assumption. Hence,
+$\frac{\alpha(u)}{\beta(u)}=\frac{\sum_{i}\lambda_{i}\alpha(w_{i})}{\sum_{i}\lambda_{i}\beta(w_{i})}=\sum_{\beta(w_{i})>0}\frac{\lambda_{i}\beta(w_{i})}{\sum_{k}\lambda_{k}\beta(w_{k})}\cdot\frac{\alpha(w_{i})}{\beta(w_{i})}.$
+Thus $\alpha(u)/\beta(u)$ is a convex combination of the finitely many numbers $\alpha(w_{i})/\beta(w_{i})$ with $\beta(w_{i})>0$. Therefore,
+$\frac{\alpha(u)}{\beta(u)}\leq\max_{\{i:\beta(w_{i})>0\}}\frac{\alpha(w_{i})}{\beta(w_{i})}$
+for every $u\in\tau^{\circ}$, and hence the same bound holds for $\sup_{u\in\tau^{\circ}}$.
+Conversely, let $w_{i}$ be a ray generator of $\tau$ with $\beta(w_{i})>0$.
+Choose $z\in\tau^{\circ}$ and set $u_{\varepsilon}=(1-\varepsilon)w_{i}+\varepsilon z$ for $0<\varepsilon<1$.
+Then, $u_{\varepsilon}\in\tau^{\circ}$ and $u_{\varepsilon}\to w_{i}$ as $\varepsilon\to 0^{+}$.
+By continuity of $\alpha$ and $\beta$ and since $\beta(w_{i})>0$, we have $\alpha(u_{\varepsilon})/\beta(u_{\varepsilon})\to\alpha(w_{i})/\beta(w_{i})$.
+Hence, $\sup_{u\in\tau^{\circ}}\alpha(u)/\beta(u)\geq\alpha(w_{i})/\beta(w_{i})$.
+Combining the two inequalities shows that
+$\sup_{u\in\tau^{\circ}}\frac{\alpha(u)}{\beta(u)}=\max_{\{i:\beta(w_{i})>0\}}\frac{\alpha(w_{i})}{\beta(w_{i})}.$
+Since $\Sigma$ has finitely many cones and every ray of $\Sigma$ occurs among the rays of some cone, taking the maximum over all cones gives
+$\sup_{u\in\sigma^{\circ}}\frac{\alpha(u)}{\beta(u)}=\max_{\{j:\beta(u_{j})>0\}}\frac{\alpha(u_{j})}{\beta(u_{j})}.$
+Finally, because $\alpha/\beta$ is homogeneous of degree 0, it suffices to approximate interior points by interior lattice points.
+Rational points are dense in $\sigma^{\circ}$, and any interior rational point becomes a lattice point after multiplying by a positive integer.
+Homogeneity implies that the ratio $\alpha/\beta$ is unchanged by such scaling, so the supremum over $\sigma^{\circ}$ equals the supremum over $\sigma^{\circ}\cap N$.
 ∎
-Corollary 10.16 (Polyhedral Newton data $\Rightarrow$ finite toric candidate set).
-Assume that the toric valuation functions
-$u\longmapsto v_{u}(\mathfrak{a}_{\bullet}),\quad u\longmapsto v_{u}(\mathfrak{b}_{\bullet})$
-are piecewise-linear on a common finite fan $\Sigma$ supported on $\sigma$ (for example, when the relevant Newton data is rational polyhedral and one takes $\Sigma$ a common refinement of the associated normal fans).
-Then, the toric supremum is a finite maximum:
-$\sup_{u\in\sigma^{\circ}\cap N}\frac{v_{u}(\mathfrak{a}_{\bullet})}{v_{u}(\mathfrak{b}_{\bullet})}\;=\;\max_{u\in\mathrm{Ray}(\Sigma)}\frac{v_{u}(\mathfrak{a}_{\bullet})}{v_{u}(\mathfrak{b}_{\bullet})}.$
+In the next few results, if for each monomial filtration $\mathfrak{a}_{\bullet}$}, the function $u\mapsto v_{u}({\mathfrak{a}_{\bullet}})$ on $\sigma^{\circ}\cap N$ admits a continuous positively homogeneous extension to $\sigma$, then we shall denote the value of this extension at $u\in\sigma$ by $\widetilde{v_{u}}({\mathfrak{a}_{\bullet}})$.
+Corollary 10.16.
+Let $\mathfrak{a}_{\bullet}$} and $\mathfrak{b}_{\bullet}$} be monomial filtrations.
+Assume that the functions $\sigma^{\circ}\cap N\to\mathbb{R}_{\geq 0}$, $u\mapsto v_{u}({\mathfrak{a}_{\bullet}})$ and $u\mapsto v_{u}({\mathfrak{b}_{\bullet}})$, admit continuous extensions to $\sigma$, positively homogeneous of degree $1$, and that there exists a finite rational polyhedral fan $\Sigma$ supported on $\sigma$ such that both extended functions are linear on every cone $\tau\in\Sigma$.
+Assume also that $v_{u}({\mathfrak{b}_{\bullet}})>0$ for all $u\in\sigma^{\circ}$.
+Let $\rho_{1},\dots,\rho_{M}$ be the rays of $\Sigma$ and let $u_{1},\dots,u_{M}\in N$ be their primitive generators. Then,
+$\sup_{u\in\sigma^{\circ}\cap N}\frac{v_{u}({\mathfrak{a}_{\bullet}})}{v_{u}({\mathfrak{b}_{\bullet}})}=\max_{\{j:\widetilde{v}_{u_{j}}({\mathfrak{b}_{\bullet}})>0\}}\frac{\widetilde{v}_{u_{j}}({\mathfrak{a}_{\bullet}})}{\widetilde{v}_{u_{j}}({\mathfrak{b}_{\bullet}})}.$
 Proof.
-Apply Lemma 10.15 with $\alpha(u)=v_{u}(a_{\bullet})$ and $\beta(u)=v_{u}(b_{\bullet})$. ∎
+Apply Lemma 10.15 with $\alpha(u)=\widetilde{v}_{u}({\mathfrak{a}_{\bullet}})$ and $\beta(u)=\widetilde{v}_{u}({\mathfrak{b}_{\bullet}})$.
+Since $\widetilde{v}_{u}({\mathfrak{a}_{\bullet}})=v_{u}({\mathfrak{a}_{\bullet}})$ and $\widetilde{v}_{u}({\mathfrak{b}_{\bullet}})=v_{u}({\mathfrak{b}_{\bullet}})$ for $u\in\sigma^{\circ}\cap N$, the claimed equality follows.
+∎
 Remark 10.17.
@@
 Theorem 10.18 (Toric valuative formula for monomial filtrations).
-Let $\mathfrak{a}_{\bullet}=\{\mathfrak{a}_{p}\}_{p\geq 1}$ and $\mathfrak{b}_{\bullet}=\{\mathfrak{b}_{q}\}_{q\geq 1}$ be graded families of monomial ideals in $R$.
-(1) (Universal toric lower bound) One has
-$\mathcal{L}_{\mathfrak{b}_{\bullet}}(\mathfrak{a}_{\bullet})\ \geq\ \sup_{u\in\sigma^{\circ}\cap N}\frac{v_{u}(\mathfrak{a}_{\bullet})}{v_{u}(\mathfrak{b}_{\bullet})}=\sup_{u\in\sigma^{\circ}\cap N}\frac{\inf_{y\in\operatorname{NP}(\mathfrak{a}_{\bullet})}\langle u,y\rangle}{\inf_{x\in\operatorname{NP}(\mathfrak{b}_{\bullet})}\langle u,x\rangle}.$
-(2) (Finite–max among toric valuations) Suppose that the functions $u\mapsto v_{u}(\mathfrak{a}_{\bullet})$ and $u\mapsto v_{u}(\mathfrak{b}_{\bullet})$ are piecewise-linear on a common finite rational fan $\Sigma$ supported on $\sigma$.
-Then
-$\sup_{u\in\sigma^{\circ}\cap N}\frac{v_{u}(\mathfrak{a}_{\bullet})}{v_{u}(\mathfrak{b}_{\bullet})}=\max_{u\in\mathrm{Ray}(\Sigma)}\frac{v_{u}(\mathfrak{a}_{\bullet})}{v_{u}(\mathfrak{b}_{\bullet})}.$
-(3) (Equality with $\mathcal{L}_{\mathfrak{b}_{\bullet}}(\mathfrak{a}_{\bullet})$ under toric testing) Assume in addition that for all $p,q$ the containment $\mathfrak{b}_{q}\subseteq\overline{\mathfrak{a}_{p}}$ is equivalent to the system of inequalities
-$v_{u}(\mathfrak{b}_{q})\geq v_{u}(\mathfrak{a}_{p})\quad\text{for all }u\in\mathrm{Ray}(\Sigma).$
-Then,
-$\mathcal{L}_{\mathfrak{b}_{\bullet}}(\mathfrak{a}_{\bullet})=\max_{u\in\mathrm{Ray}(\Sigma)}\frac{v_{u}(\mathfrak{a}_{\bullet})}{v_{u}(\mathfrak{b}_{\bullet})}.$
+Let $\mathfrak{a}_{\bullet}$} and $\mathfrak{b}_{\bullet}$} be $\mathfrak{m}_{S}$-primary monomial filtrations, and assume that $\mathcal{R}({\mathfrak{a}_{\bullet}})$ is Noetherian.
+Suppose moreover that there exist an $\mathfrak{m}_{S}$-primary monomial ideal $\mathfrak{q}$ and a finite rational polyhedral fan $\Sigma$ supported on $\sigma$ such that:
+  1. (1)
+    $\mathfrak{a}_{\bullet}$} and $\mathfrak{b}_{\bullet}$} are linearly bounded with respect to $\mathfrak{q}$;
+  2. (2)
+    the functions $u\mapsto v_{u}({\mathfrak{a}_{\bullet}})$ and $u\mapsto v_{u}({\mathfrak{b}_{\bullet}})$ on $\sigma^{\circ}\cap N\to\mathbb{R}_{\geq 0}$ admit continuous extensions to $\sigma$, positively homogeneous of degree $1$, and linear on every cone $\tau\in\Sigma$.
+Let $\rho_{1},\dots,\rho_{M}$ be the rays of $\Sigma$, and let $u_{1},\dots,u_{M}\in N$ be their primitive generators. Then
+$\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})=\sup_{u\in\sigma^{\circ}\cap N}\frac{v_{u}({\mathfrak{a}_{\bullet}})}{v_{u}({\mathfrak{b}_{\bullet}})}=\max_{\{j:\widetilde{v}_{u_{j}}({\mathfrak{b}_{\bullet}})>0\}}\frac{\widetilde{v}_{u_{j}}({\mathfrak{a}_{\bullet}})}{\widetilde{v}_{u_{j}}({\mathfrak{b}_{\bullet}})}.$
+In particular, $\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})$ is computed by finitely many toric divisorial valuations.
 Proof.
-(1) By Theorem 5.6 we have the universal bound $\mathcal{L}_{\mathfrak{b}_{\bullet}}(\mathfrak{a}_{\bullet})\geq\sup_{v}\frac{v(\mathfrak{a}_{\bullet})}{v(\mathfrak{b}_{\bullet})}$ over divisorial valuations $v$ centered at $\mathfrak{m}_{S}$.
-Restricting to toric divisorial valuations $v=v_{u}$ gives the first inequality.
-The expression in terms of $\operatorname{NP}(\mathfrak{a}_{\bullet})$ and $\operatorname{NP}(\mathfrak{b}_{\bullet})$ follows from Lemma 10.14.
-(2) This is exactly Lemma 10.15 applied to the functions $u\mapsto v_{u}(\mathfrak{a}_{\bullet})$ and $u\mapsto v_{u}(\mathfrak{b}_{\bullet})$.
-(3) Under the stated testing equivalence, the finite testing hypothesis (H1) of Theorem 5.8 holds with $V=\{v_{u}:u\in\mathrm{Ray}(\Sigma)\}$, hence Theorem 5.8 yields the desired equality. ∎
+By Theorem 5.6, we have
+$\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\geq\sup\limits_{u\in\sigma^{\circ}\cap N}\frac{v_{u}({\mathfrak{a}_{\bullet}})}{v_{u}({\mathfrak{b}_{\bullet}})}.$
+By Corollary 10.16, the latter supremum is equal to $\max\limits_{\{j:\widetilde{v}_{u_{j}}({\mathfrak{b}_{\bullet}})>0\}}\frac{\widetilde{v}_{u_{j}}({\mathfrak{a}_{\bullet}})}{\widetilde{v}_{u_{j}}({\mathfrak{b}_{\bullet}})}.$ Thus, it remains to prove the reverse inequality.
+Since $\mathcal{R}({\mathfrak{a}_{\bullet}})$ is finitely generated, Theorem 5.9 gives a finite set of divisorial valuations $\mathcal{V}({\mathfrak{a}_{\bullet}})=\{v_{1},\dots,v_{r}\}$ such that
+$\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})=\max_{1\leq i\leq r}\frac{v_{i}({\mathfrak{a}_{\bullet}})}{v_{i}({\mathfrak{b}_{\bullet}})},$
+provided that $v_{i}({\mathfrak{b}_{\bullet}})>0$ for all $i$. We shall verify this positivity condition.
+Indeed, since $\mathcal{R}({\mathfrak{a}_{\bullet}})$ is Noetherian, there exists $d\geq 1$ such that the $d$-th Veronese subalgebra $\mathcal{R}({\mathfrak{a}_{\bullet}})^{(d)}$ is standard graded.
+Equivalently, $\mathcal{R}({\mathfrak{a}_{\bullet}})^{(d)}$ is the ordinary Rees algebra of the monomial ideal $\mathfrak{a}_{d}$.
+Hence, the valuations in $\mathcal{V}({\mathfrak{a}_{\bullet}})$ are exactly the Rees valuations of $\mathfrak{a}_{d}$.
+By Theorem 10.11, these are toric divisorial valuations.
+Thus, there exist primitive lattice vectors $w_{1},\dots,w_{r}\in\sigma^{\circ}\cap N$ such that $v_{i}=v_{w_{i}}$ for all $i$.
+Since $\mathfrak{b}_{\bullet}$} is linearly bounded with respect to $\mathfrak{q}$, there exists $c>0$ such that $v_{u}({\mathfrak{b}_{\bullet}})\geq c\,h_{\mathfrak{q}}(u)$ for all $u\in\sigma^{\circ}$.
+Therefore, $v_{w_{i}}({\mathfrak{b}_{\bullet}})\geq c\,h_{\mathfrak{q}}(w_{i})>0$ for every $i$, verifying the desired positive condition.
+Now, each $w_{i}$ lies in $\sigma^{\circ}\cap N$, so $\frac{v_{w_{i}}({\mathfrak{a}_{\bullet}})}{v_{w_{i}}({\mathfrak{b}_{\bullet}})}\leq\sup\limits_{u\in\sigma^{\circ}\cap N}\frac{v_{u}({\mathfrak{a}_{\bullet}})}{v_{u}({\mathfrak{b}_{\bullet}})}$ for every $i$.
+Taking the maximum over $i$, we obtain the reverse inequality
+$\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\leq\sup_{u\in\sigma^{\circ}\cap N}\frac{v_{u}({\mathfrak{a}_{\bullet}})}{v_{u}({\mathfrak{b}_{\bullet}})}.$
+The theorem is proved. ∎
 Remark 10.19 (Mechanism behind finite testing in the toric case).
@@
 Let $\Sigma_{\Gamma}$ denote the normal fan of $\Gamma(F)$ and let $\Sigma_{\mathbb{P}^{n}}$ denote the standard fan of $\mathbb{P}^{n}$.
-Choose a smooth complete fan $\Sigma$ refining both $\Sigma_{\Gamma}$ and $\Sigma_{\mathbb{P}^{n}}$ (such a refinement exists by toric resolution of singularities; see [33, 42], [11, Chapter 11] and [18, Chapter 2]).
-Then, the associated toric variety $X_{\Sigma}$ is a smooth projective toric compactification of the torus $T=(k^{*})^{n}$, and the refinement $\Sigma\succ\Sigma_{\mathbb{P}^{n}}$ induces a proper toric morphism $\pi_{\Sigma}:X_{\Sigma}\longrightarrow\mathbb{P}^{n}$ extending the standard open embedding $T\hookrightarrow\mathbb{P}^{n}$ (see [11, Chapter 3] and [18, Chapters 1–2]).
+Choose a smooth complete fan $\Sigma$ refining both $\Sigma_{\Gamma}$ and $\Sigma_{\mathbb{P}^{n}}$ (such a refinement exists by toric resolution of singularities; see [37, 46], [14, Chapter 11] and [22, Chapter 2]).
+Then, the associated toric variety $X_{\Sigma}$ is a smooth projective toric compactification of the torus $T=(k^{*})^{n}$, and the refinement $\Sigma\succ\Sigma_{\mathbb{P}^{n}}$ induces a proper toric morphism $\pi_{\Sigma}:X_{\Sigma}\longrightarrow\mathbb{P}^{n}$ extending the standard open embedding $T\hookrightarrow\mathbb{P}^{n}$ (see [14, Chapter 3] and [22, Chapters 1–2]).
 Moreover, since $\Sigma$ also refines the normal fan of $\Gamma(F)$, the toric boundary divisors of $X_{\Sigma}$ correspond to weight vectors governing the face systems (initial forms) of $F$, which control the asymptotic behavior at infinity as in Khovanskii’s compactification.
@@
 Proof.
-We work on a toric compactification attached to $\Gamma(F)$ and then descend to divisorial valuations of $R_{\xi}$.
+We will work with the notations in Definition 11.4 and Section 4.3.
 Step 1: toric boundary data and face systems.
@@
 Let $\Gamma_{F}^{\Sigma}\subseteq X_{\Sigma}\times\mathbb{P}^{r}$ be as before, and let $\mu_{\Sigma}:\Gamma_{F}^{\Sigma}\to\overline{\Gamma}_{F}$ be the induced proper birational morphism.
-Fix a torus–invariant divisor $E\subseteq D_{\operatorname{Tor}}$ and consider the locally closed stratum $O(E)\subseteq X_{\Sigma}$.
-Work on the target chart $\{W_{0}\neq 0\}\subseteq\mathbb{P}^{r}$, with affine coordinates $y_{j}:=W_{j}/W_{0}$.
-Restricting to $O(E)\times\mathbb{A}^{r}$, we work in an affine toric chart $U_{\sigma}=\operatorname{Spec}k[\sigma^{\vee}\cap M]$ containing $O(E)$, where $\sigma\in\Sigma$ is a cone whose rays include $w_{E}$.
-In the corresponding toric coordinates $u_{1},\dots,u_{n}$, each coordinate $x_{i}$ is a monomial in the $u_{j}$, and each polynomial $f_{j}(x)$ admits a decomposition of the form
-$f_{j}(x)=u^{m_{j}}\Bigl((f_{j})_{w_{E}}(u)+\sum_{\langle w_{E},\alpha\rangle>m_{j}}c_{j,\alpha}u^{\alpha}\Bigr),$
-where $m_{j}=\min_{\alpha\in\Gamma(f_{j})}\langle w_{E},\alpha\rangle$, and $(f_{j})_{w_{E}}$ is the face polynomial consisting of those terms achieving this minimum.
-Consequently, the strict transform of the graph in $U_{\sigma}\times\mathbb{A}^{r}$ meets $O(E)\times\mathbb{A}^{r}$ if and only if the face system
-$(f_{1})_{w_{E}}=\cdots=(f_{r})_{w_{E}}=0$
-admits a solution in the torus $O(E)\cong(k^{*})^{n}$. This is the standard initial-form criterion for strict transforms in toric modifications; see, for example, [33, 42]. In particular,
-(11.7) $\Gamma_{F}^{\Sigma}\cap\bigl(O(E)\times\{W_{0}\neq 0\}\bigr)=\varnothing\Leftrightarrow(f_{1})_{w_{E}}=\cdots=(f_{r})_{w_{E}}=0\ \text{has no solution in }(k^{*})^{n}.$
-Since $F$ is Newton nondegenerate at infinity, the right-hand condition holds for every $w_{E}\in\mathbb{R}_{>0}^{n}$, and so (11.7) implies: for every boundary divisor $E$ with $w_{E}\in\mathbb{R}_{>0}^{n}$,
-$\Gamma_{F}^{\Sigma}\cap\bigl(O(E)\times\{W_{0}\neq 0\}\bigr)=\varnothing,$
-i.e.
-every point of $\Gamma_{F}^{\Sigma}$ lying over $O(E)$ maps into the hyperplane at infinity $H_{Y}=\{W_{0}=0\}$.
-Equivalently, $\rho_{\Sigma}^{-1}(H_{Y})$ contains $\mu_{\Sigma}^{-1}(\pi_{\Sigma}^{-1}(H_{X}))$ along all such boundary strata.
+We first claim that near $\mu_{\Sigma}^{-1}(\xi)$, one has
+(11.7) $\operatorname{\mathrm{Supp}}(\mathscr{I}_{X}\mathcal{O}_{\Gamma_{F}^{\Sigma}})\subseteq\operatorname{\mathrm{Supp}}(\mathscr{I}_{Y}\mathcal{O}_{\Gamma_{F}^{\Sigma}}).$
+Indeed, consider any point $z\in\operatorname{\mathrm{Supp}}(\mathscr{I}\mathcal{O}_{\Gamma_{F}^{\Sigma}})$.
+Then, $\mu_{\Sigma}(z)\in D_{X}$, so the image of $z$ in $X_{\Sigma}$ lies over the hyperplane at infinity $H_{X}$.
+Let $O(\sigma)\subseteq X_{\Sigma}$ be the torus orbit containing the image of $z$, where $\sigma\in\Sigma$.
+Since $O(\sigma)$ maps into $H_{X}$, the cone $\sigma$ is not contained in any coordinate face of $\mathbb{R}_{\geq 0}^{n}$.
+Equivalently, $\text{relint}(\sigma)\cap\mathbb{R}_{>0}^{n}\neq\varnothing.$ Choose $w\in\text{relint}(\sigma)\cap\mathbb{R}_{>0}^{n}.$
+We shall work in the affine toric chart $U_{\sigma}=\operatorname{Spec}k[\sigma^{\vee}\cap M]\subseteq X_{\Sigma}$ containing $O(\sigma)$.
+In toric coordinates on $U_{\sigma}$, each polynomial $f_{j}$ has an expansion of the form
+$f_{j}=u^{m_{j}}\Bigl((f_{j})_{w}(u)+\sum_{\langle w,\alpha\rangle>m_{j}}c_{j,\alpha}u^{\alpha}\Bigr),$
+where $m_{j}=m_{w}(f_{j})=\min_{\alpha\in\Gamma(f_{j})}\langle w,\alpha\rangle$ and $(f_{j})_{w}$ is the face polynomial consisting of those terms achieving this minimum.
+By the standard initial-form criterion for strict transforms in toric compactifications (see, for example, [37, 46]), the graph closure and its normalization in $U_{\sigma}\times\mathbb{A}^{r}$ meet $O(\sigma)\times\{W_{0}\neq 0\}$ if and only if the face system
+$(f_{1})_{w}=\cdots=(f_{r})_{w}=0$
+has a solution in the torus $O(\sigma)\cong(k^{*})^{\,n-\dim\sigma}.$ Since $w\in\mathbb{R}_{>0}^{n}$ and $F$ is Newton nondegenerate at infinity, this system has no solution.
+Therefore, we conclude that
+$\Gamma_{F}^{\Sigma}\cap\bigl(O(\sigma)\times\{W_{0}\neq 0\}\bigr)=\varnothing.$
+[…]
+$z\in\operatorname{\mathrm{Supp}}(\mathscr{I}_{Y}\mathcal{O}_{\Gamma_{F}^{\Sigma}}).$
+Since $z$ was arbitrary, this proves (11.7) near $\mu_{\Sigma}^{-1}(\xi)$.
 Step 3: descend to divisorial valuations over $\xi$.
-Let $v$ be a divisorial valuation of the local domain $R_{\xi}=\mathcal{O}_{\overline{\Gamma}_{F},\xi}$.
-Choose a normal modification $\nu:W\to\overline{\Gamma}_{F}$ and a prime divisor $E^{\prime}\subseteq W$ with center $\xi$ such that $v=\operatorname{\mathrm{ord}}_{E^{\prime}}$ on $R_{\xi}$.
-Take a common normal refinement $W^{\prime}\to W$ and $W^{\prime}\to\Gamma_{F}^{\Sigma}$ dominating both models, and let $\widetilde{E}^{\prime}\subseteq W^{\prime}$ be the strict transform of $E^{\prime}$.
-Then
-[…]
-Assume $v(\mathscr{I}_{X,\xi})>0$.
-This means that $\widetilde{E}^{\prime}$ lies over $D_{X}=\pi^{-1}(H_{X})$, so its image in $X_{\Sigma}$ is contained in the toric boundary lying over $H_{X}$.
-Hence, $\widetilde{E}^{\prime}$ dominates (after possibly replacing $\Sigma$ by a further refinement) a torus–invariant boundary divisor $E\subseteq D_{\operatorname{Tor}}$ with ray vector $w_{E}\in\mathbb{R}_{>0}^{n}$.
-Indeed, divisors corresponding to rays lying on coordinate faces of the cone $\mathbb{R}_{\geq 0}^{n}$ map into the coordinate boundary of $\mathbb{A}^{n}_{k}\subseteq\mathbb{P}^{n}$ rather than into the hyperplane at infinity $H_{X}=\{X_{0}=0\}$.
-Since $\widetilde{E}^{\prime}$ lies over $D_{X}=\pi^{-1}(H_{X})$, it must dominate a torus-invariant divisor corresponding to a ray in the relative interior $\mathbb{R}_{>0}^{n}$ of the cone.
-This is a standard property of toric compactifications attached to Newton polyhedra; see [33, 42].
-By Step 2, $\Gamma_{F}^{\Sigma}$ over $O(E)$ maps into $H_{Y}$, so $\widetilde{E}^{\prime}$ is contained in the pullback of $D_{Y}=\rho^{-1}(H_{Y})$.
-Equivalently, $\operatorname{\mathrm{ord}}_{\widetilde{E}^{\prime}}(\mathscr{I}_{Y}\mathcal{O}_{W^{\prime}})>0$, i.e.
-$v(\mathscr{I}_{Y,\xi})>0$.
-This proves $v(\mathscr{I}_{X,\xi})>0\Rightarrow v(\mathscr{I}_{Y,\xi})>0$. The second statement is immediate. ∎
+Let $v$ be a divisorial valuation of $R_{\xi}$ with $v(\mathscr{I}_{X,\xi})>0.$ Choose a normal modification $\nu:W\to\overline{\Gamma}_{F}$ and a prime divisor $E^{\prime}\subseteq W$ with center $\xi$ such that $v=\operatorname{\mathrm{ord}}_{E^{\prime}}$ on $R_{\xi}$.
+Take a common normal refinement $\eta:W^{\prime}\to W$ and $\tau:W^{\prime}\to\Gamma_{F}^{\Sigma}$ dominating both models, and let $\widetilde{E}^{\prime}\subseteq W^{\prime}$ be the strict transform of $E^{\prime}$.
+Since $\mathscr{I}_{X}$ is the ideal sheaf of the Cartier divisor $D_{X}$ on $\overline{\Gamma}_{F}$, its pullback to $W^{\prime}$ is invertible, and
+$\operatorname{\mathrm{ord}}_{\widetilde{E}^{\prime}}(\tau^{*}\mathscr{I}_{X})=\operatorname{\mathrm{ord}}_{E^{\prime}}(\mathscr{I}_{X,\xi})=v(\mathscr{I}_{X,\xi})>0.$
+Therefore, the generic point of $\widetilde{E}^{\prime}$ maps into $\operatorname{\mathrm{Supp}}(\mathscr{I}_{X}\mathcal{O}_{\Gamma_{F}^{\Sigma}}).$ By (11.7), near $\mu_{\Sigma}^{-1}(\xi)$ this support is contained in $\operatorname{\mathrm{Supp}}(\mathscr{I}_{Y}\mathcal{O}_{\Gamma_{F}^{\Sigma}}).$ Hence, the generic point of $\widetilde{E}^{\prime}$ also maps into $\operatorname{\mathrm{Supp}}(\mathscr{I}_{Y}\mathcal{O}_{\Gamma_{F}^{\Sigma}}).$ This implies that $\operatorname{\mathrm{ord}}_{\widetilde{E}^{\prime}}(\tau^{*}\mathscr{I}_{Y})>0.$
+Moreover, $\mathscr{I}_{Y}$ is also the ideal sheaf of a Cartier divisor on $\overline{\Gamma}_{F}$. Thus,
+$\operatorname{\mathrm{ord}}_{\widetilde{E}^{\prime}}(\tau^{*}\mathscr{I}_{Y})=\operatorname{\mathrm{ord}}_{E^{\prime}}(\mathscr{I}_{Y,\xi})=v(\mathscr{I}_{Y,\xi}).$
+It follows that $v(\mathscr{I}_{Y,\xi})>0$, as desired. The second statement of the lemma is immediate. ∎
 Corollary 11.7.
@@
   2. (2)
-    for every $\xi\in D_{X}\cap D_{Y}$, after base change to $\operatorname{Spec}(R_{\xi})$ the ideals $\mathscr{I}_{X,\xi}$ and $\mathscr{I}_{Y,\xi}$ pull back to invertible ideals supported on the strict transforms of those $E_{i}$ meeting $\mu^{-1}(\xi)$.
+    for every $\xi\in D_{X}\cap D_{Y}$, after base change to $\operatorname{Spec}(R_{\xi})$ the ideals $\mathscr{I}_{X,\xi}$ and $\mathscr{I}_{Y,\xi}$ pull back to invertible ideals supported on the pullbacks of those $E_{i}$ meeting $\mu^{-1}(\xi)$.
 Proof.
@@
 Denote its irreducible components by $E_{1},\dots,E_{N}$, proving (1).
-For (2), fix $\xi\in D_{X}\cap D_{Y}$ and base change $Z\to\overline{\Gamma}_{F}$ to $\operatorname{Spec}(R_{\xi})$.
-On the smooth scheme $Z_{\xi}:=Z\times_{\overline{\Gamma}_{F}}\operatorname{Spec}(R_{\xi})$, the pullbacks of $\mathscr{I}_{X,\xi}$ and $\mathscr{I}_{Y,\xi}$ define Cartier divisors supported on the pullback of $(D_{X})_{\text{red}}+(D_{Y})_{\text{red}}$ and, hence, on the strict transforms of those $E_{i}$ meeting $\mu^{-1}(\xi)$.
-In particular, both ideals become invertible on $Z_{\xi}$ with support contained in those components, as claimed.
+For (2), fix $\xi\in D_{X}\cap D_{Y}$ and set $Z_{\xi}:=Z\times_{\overline{\Gamma}_{F}}\operatorname{Spec}(R_{\xi})$.
+Since $D_{X}$ and $D_{Y}$ are Cartier divisors on $\overline{\Gamma}_{F}$, their ideal sheaves $\mathscr{I}_{X}$ and $\mathscr{I}_{Y}$ are invertible.
+Hence, their pullbacks to $Z_{\xi}$ remain invertible ideal sheaves and define Cartier divisors supported on the pullback of $(D_{X})_{\mathrm{red}}+(D_{Y})_{\mathrm{red}}$.
+By (1), the divisor $(D_{X})_{\mathrm{red}}+(D_{Y})_{\mathrm{red}}$ is simple normal crossings on $Z$, with irreducible components $E_{1},\dots,E_{N}$.
+Thus, locally on $Z$, it is defined by an equation $z_{1}\cdots z_{r}=0$, where $z_{j}=0$ cuts out one of the components $E_{i}$.
+After base change to $\operatorname{Spec}(R_{\xi})$, the pullback divisor is still locally defined by $z_{1}\cdots z_{r}=0$, so its irreducible components are exactly the pullbacks $E_{i}\times_{\overline{\Gamma}_{F}}\operatorname{Spec}(R_{\xi})$ for those $i$ with $E_{i}\cap\mu^{-1}(\xi)\neq\varnothing$.
+Therefore the pullbacks of $\mathscr{I}_{X}$ and $\mathscr{I}_{Y}$ to $Z_{\xi}$ are invertible ideal sheaves whose supports are contained in those components.
 ∎
@@
 where $\overline{(\cdot)}$ denotes integral closure. Then, on the same model $\pi:Y\to\operatorname{Spec}R$ of Theorem 12.1, one has for every $t$:
-$\mathfrak{a}_{t}\ =\ \pi_{*}\mathcal{O}_{Y}\Big(-\sum_{i=1}^{N}\Big\lfloor t\cdot\sum_{j=1}^{r}\lambda_{j}a_{ij}\Big\rfloor\,E_{i}\Big).$
+$\mathfrak{a}_{t}\ =\ \pi_{*}\mathcal{O}_{Y}\Big(-\sum_{i=1}^{N}d_{i}(t)E_{i}\Big),$
+where $d_{i}(t)=\sum_{j=1}^{r}a_{ij}\lceil\lambda_{j}t\rceil$.
 In particular, for any other family $\mathfrak{b}_{\bullet}$} of the same form (possibly with different coefficients), the containment relations $\mathfrak{b}_{q}\subseteq\mathfrak{a}_{p}$ are determined by the finitely many divisorial valuations $\{\operatorname{\mathrm{ord}}_{E_{i}}\}_{i=1}^{N}$, i.e., the hypotheses of the finite-testing criterion in Theorem 5.8 are satisfied with testing set $\{\operatorname{\mathrm{ord}}_{E_{i}}\}$.
@@
 Let $\pi:Y\to\operatorname{Spec}R$ be as in Theorem 12.1. As constructed, $Y$ is regular, so $Y$ is normal. Write $I^{(j)}\mathcal{O}_{Y}=\mathcal{O}_{Y}(-F_{j})$ with $F_{j}=\sum_{i}a_{ij}E_{i}$. Set
-[…]
+$D_{\mathfrak{a}_{\bullet}}(t)\ :=\ \sum_{j=1}^{r}\lceil\lambda_{j}t\rceil\,F_{j}\ =\ \sum_{i=1}^{N}d_{i}(t)E_{i},\text{ with }d_{i}(t):=\sum_{j=1}^{r}a_{ij}\lceil\lambda_{j}t\rceil\in\mathbb{N}.$
 Step 1: compute integral closures on a fixed principalization. Since $I^{(j)}\mathcal{O}_{Y}$ is invertible, so is $\prod_{j}(I^{(j)})^{\lceil\lambda_{j}t\rceil}\mathcal{O}_{Y}$, and in fact
-$\Big(\prod_{j=1}^{r}(I^{(j)})^{\lceil\lambda_{j}t\rceil}\Big)\mathcal{O}_{Y}\ =\ \mathcal{O}_{Y}\big(-D(t)\big).$
+$\Big(\prod_{j=1}^{r}(I^{(j)})^{\lceil\lambda_{j}t\rceil}\Big)\mathcal{O}_{Y}\ =\ \mathcal{O}_{Y}\big(-D_{\mathfrak{a}_{\bullet}}(t)\big).$
 By the fixed-model integral-closure reduction (Lemma 5.4), the integral closure of $\prod_{j}(I^{(j)})^{\lceil\lambda_{j}t\rceil}$ equals the pushforward of the corresponding invertible sheaf:
-$\mathfrak{a}_{t}\ =\ \overline{\prod_{j=1}^{r}(I^{(j)})^{\lceil\lambda_{j}t\rceil}}\ =\ \pi_{*}\mathcal{O}_{Y}\big(-D(t)\big).$
-Step 2: replace $\lceil\lambda_{j}t\rceil$ by linear bounds in $t$.
-Write $D(t)=\sum_{i}d_{i}(t)E_{i}$ as above.
-Since $d_{i}(t)$ is an integer-valued, piecewise linear function in $t$ with slopes $\sum_{j}\lambda_{j}a_{ij}$, we have
-$d_{i}(t)\ =\ \Big\lfloor t\cdot\sum_{j=1}^{r}\lambda_{j}a_{ij}\Big\rfloor\ +\ \epsilon_{i}(t),\quad\epsilon_{i}(t)\in\{0,1,\dots,r\}.$
-Absorbing the bounded error into flooring yields the displayed formula; this is the same bookkeeping already used in Section 5 when passing between integer- and real-indexed families.
-Step 3: finite testing.
+$\mathfrak{a}_{t}\ =\ \overline{\prod_{j=1}^{r}(I^{(j)})^{\lceil\lambda_{j}t\rceil}}\ =\ \pi_{*}\mathcal{O}_{Y}\big(-D_{\mathfrak{a}_{\bullet}}(t)\big).$
+Step 2: finite testing.
 For any $p,q$, the condition $\mathfrak{b}_{q}\subseteq\mathfrak{a}_{p}$ is equivalent to the inequality of divisors on $Y$: $D_{\mathfrak{b}_{\bullet}}(q)\ \geq\ D_{\mathfrak{a}_{\bullet}}(p)$, i.e., coefficientwise inequalities along $E_{i}$.
@@
 We now push the previous verification to a genuine structural statement in parameter families.
-Theorem 12.3 (Polyhedral wall–chamber structure).
+Theorem 12.3 (Wall–chamber stratification).
 Assume the setup of Theorem 12.2, and let $S\subseteq\mathbb{R}^{m}$ be a parameter set. Suppose that for each $s\in S$ we have two families
-${\mathfrak{a}_{\bullet}}(s)=\{\mathfrak{a}(s)_{t}\}_{t\geq 0},\quad{\mathfrak{b}_{\bullet}}(s)=\{\mathfrak{b}(s)_{t}\}_{t\geq 0},$
+${\mathfrak{a}_{\bullet}}(s)=\{\mathfrak{a}(s)_{t}\}_{t\geq 0}\text{ and }{\mathfrak{b}_{\bullet}}(s)=\{\mathfrak{b}(s)_{t}\}_{t\geq 0},$
 each of the form in Theorem 12.2, with coefficients $\lambda_{j}=\lambda_{j}(s)$ depending piecewise linearly on $s$.
 Assume moreover that the Newton polyhedra $\Gamma(I^{(j)}_{s})$ are independent of $s$ and the nondegeneracy is uniform, so that a single toric model $\pi:Y\to\operatorname{Spec}R$ principalizes all $I^{(j)}_{s}$ simultaneously.
-Then, there exists a finite polyhedral decomposition $S=\bigcup_{\alpha}C_{\alpha}$ such that on each chamber $C_{\alpha}$ there is an index $i(\alpha)\in\{1,\dots,N\}$ with
-$\mathcal{L}_{{\mathfrak{b}_{\bullet}}(s)}\big({\mathfrak{a}_{\bullet}}(s)\big)\ =\ \frac{\operatorname{\mathrm{ord}}_{E_{i(\alpha)}}\big({\mathfrak{a}_{\bullet}}(s)\big)}{\operatorname{\mathrm{ord}}_{E_{i(\alpha)}}\big({\mathfrak{b}_{\bullet}}(s)\big)}\quad\text{for all }s\in C_{\alpha}.$
-In particular, $s\mapsto\mathcal{L}_{{\mathfrak{b}_{\bullet}}(s)}({\mathfrak{a}_{\bullet}}(s))$ is piecewise rational-linear and the maximizing divisor is locally constant on chambers.
+Then, there exists a finite stratification $S=\bigsqcup_{\alpha}C_{\alpha}$ by locally closed semialgebraic subsets such that on each stratum $C_{\alpha}$ one of the following holds:
+  1. (i)
+    $\mathcal{L}_{{\mathfrak{b}_{\bullet}}(s)}({\mathfrak{a}_{\bullet}}(s))=\infty$ for all $s\in C_{\alpha}$;
+  2. (ii)
+    there exists a nonempty subset $J(\alpha)\subseteq\{1,\dots,N\}$ such that for every $s\in C_{\alpha}$,
+    $\mathcal{L}_{{\mathfrak{b}_{\bullet}}(s)}({\mathfrak{a}_{\bullet}}(s))=\max_{i\in J(\alpha)}\frac{\operatorname{\mathrm{ord}}_{E_{i}}({\mathfrak{a}_{\bullet}}(s))}{\operatorname{\mathrm{ord}}_{E_{i}}({\mathfrak{b}_{\bullet}}(s))},$
+    and the maximizer set
+    $\left\{i\in J(\alpha)\ \middle|\ \frac{\operatorname{\mathrm{ord}}_{E_{i}}({\mathfrak{a}_{\bullet}}(s))}{\operatorname{\mathrm{ord}}_{E_{i}}({\mathfrak{b}_{\bullet}}(s))}=\mathcal{L}_{{\mathfrak{b}_{\bullet}}(s)}({\mathfrak{a}_{\bullet}}(s))\right\}$
+    is independent of $s\in C_{\alpha}$.
+In particular, on every stratum $C_{\alpha}$ for which the maximizer set is a singleton $\{i(\alpha)\}$, one has
+$\mathcal{L}_{{\mathfrak{b}_{\bullet}}(s)}({\mathfrak{a}_{\bullet}}(s))=\frac{\operatorname{\mathrm{ord}}_{E_{i(\alpha)}}({\mathfrak{a}_{\bullet}}(s))}{\operatorname{\mathrm{ord}}_{E_{i(\alpha)}}({\mathfrak{b}_{\bullet}}(s))}\quad\text{for all }s\in C_{\alpha}.$
+Hence, on such a stratum the function
+$s\longmapsto\mathcal{L}_{{\mathfrak{b}_{\bullet}}(s)}({\mathfrak{a}_{\bullet}}(s))$
+is given by a single ratio of affine linear functions.
 Proof.
-Fix the common model $\pi:Y\to\operatorname{Spec}R$ as in Theorem 12.2 and its exceptional prime divisors $E_{1},\dots,E_{N}$. Write $E=\sum_{i=1}^{N}E_{i}$.
-Step 1: linearity of divisor coefficients. By Theorem 12.2, for each $s$ we can write
-$\mathfrak{a}(s)_{t}=\pi_{*}\mathcal{O}_{Y}\big(-D^{(s)}_{\mathfrak{a}}(t)\big),\quad\mathfrak{b}(s)_{t}=\pi_{*}\mathcal{O}_{Y}\big(-D^{(s)}_{\mathfrak{b}}(t)\big),$
+Fix the common model $\pi:Y\to\operatorname{Spec}R$ as in Theorem 12.2 and its exceptional prime divisors $E_{1},\dots,E_{N}$.
+Step 1: affine linearity of the asymptotic divisor coefficients. By Theorem 12.2, for each $s\in S$ and each $t\geq 0$ we have
+$\mathfrak{a}(s)_{t}=\pi_{*}\mathcal{O}_{Y}\bigl(-D^{(s)}_{\mathfrak{a}}(t)\bigr)\text{ and }\mathfrak{b}(s)_{t}=\pi_{*}\mathcal{O}_{Y}\bigl(-D^{(s)}_{\mathfrak{b}}(t)\bigr),$
 where
-$D^{(s)}_{\mathfrak{a}}(t)=\sum_{i=1}^{N}d^{(s)}_{\mathfrak{a},i}(t)\,E_{i},\quad D^{(s)}_{\mathfrak{b}}(t)=\sum_{i=1}^{N}d^{(s)}_{\mathfrak{b},i}(t)\,E_{i}$
-and each coefficient $d^{(s)}_{\mathfrak{a},i}(t)$ (resp.
-$d^{(s)}_{\mathfrak{b},i}(t)$) is an integer-valued piecewise linear function whose slope in $t$ is a piecewise linear function of $s$.
-Hence, for each fixed $i$ the asymptotic slopes
-$\mu_{\mathfrak{a},i}(s):=\lim_{t\to\infty}\frac{1}{t}\operatorname{\mathrm{ord}}_{E_{i}}\big(\mathfrak{a}(s)_{t}\big),\quad\mu_{\mathfrak{b},i}(s):=\lim_{t\to\infty}\frac{1}{t}\operatorname{\mathrm{ord}}_{E_{i}}\big(\mathfrak{b}(s)_{t}\big)$
-exist and are piecewise linear functions of $s$.
-Step 2: finite-max formula on the fixed testing set.
-By the finite-testing criterion, Theorem 5.8, with testing set $\{\operatorname{\mathrm{ord}}_{E_{i}}\}$ (we can restrict to those with $\mu_{\mathfrak{b},i}(s)>0$),
-$\mathcal{L}_{{\mathfrak{b}_{\bullet}}(s)}({\mathfrak{a}_{\bullet}}(s))\ =\ \max_{1\leq i\leq N}\ \frac{\mu_{\mathfrak{a},i}(s)}{\mu_{\mathfrak{b},i}(s)}.$
-For each ordered pair $(i,j)$, the inequality
-$\frac{\mu_{\mathfrak{a},i}(s)}{\mu_{\mathfrak{b},i}(s)}\ \geq\ \frac{\mu_{\mathfrak{a},j}(s)}{\mu_{\mathfrak{b},j}(s)}$
-is equivalent to the linear inequality
-$\mu_{\mathfrak{a},i}(s)\mu_{\mathfrak{b},j}(s)-\mu_{\mathfrak{a},j}(s)\mu_{\mathfrak{b},i}(s)\ \geq\ 0.$
-By Theorem 12.2, for each fixed exceptional divisor $E_{i}$ the functions
-$\mu_{\mathfrak{a},i}(t)=\operatorname{\mathrm{ord}}_{E_{i}}(\mathfrak{a}_{t})\quad\text{and}\quad\mu_{\mathfrak{b},i}(s)=\operatorname{\mathrm{ord}}_{E_{i}}(\mathfrak{b}_{s})$
-are piecewise affine linear functions of $t$ and $s$, respectively.
-More precisely, there exist real constants $c_{\mathfrak{a},i},c_{\mathfrak{b},i}$ and bounded error terms $\varepsilon_{\mathfrak{a},i}(t),\varepsilon_{\mathfrak{b},i}(s)$ such that
-$\mu_{\mathfrak{a},i}(t)=c_{\mathfrak{a},i}\,t+\varepsilon_{\mathfrak{a},i}(t),\quad\mu_{\mathfrak{b},i}(s)=c_{\mathfrak{b},i}\,s+\varepsilon_{\mathfrak{b},i}(s),$
-where $\varepsilon_{\mathfrak{a},i}(t)$ (resp.
-$\varepsilon_{\mathfrak{b},i}(s)$) is locally constant away from a finite set of jump points determined by the ceiling functions $\lceil\lambda_{j}t\rceil$ (resp.
-$\lceil\lambda_{j}s\rceil$).
-Consequently, on each region of the (s,t)(s,t)–plane where these ceiling functions are fixed, the functions $\mu_{\mathfrak{a},i}(t)$ and $\mu_{\mathfrak{b},i}(s)$ are affine linear.
-Hence, refining $S$ by the finite arrangement of the linearity regions of all $\mu_{\mathfrak{a},i},\mu_{\mathfrak{b},i}$ yields a finite polyhedral decomposition on which all comparisons are given by linear inequalities.
-Step 3: chambers and constancy of maximizers.
-Intersecting the regions defined by the finitely many pairwise comparison inequalities for the index attaining the maximum produces the claimed chambers $C_{\alpha}$, and on each chamber the maximizing index $i(\alpha)$ is constant.
-∎
+$D^{(s)}_{\mathfrak{a}}(t)=\sum_{j}\lceil\lambda_{\mathfrak{a},j}(s)t\rceil\,F_{\mathfrak{a},j}\text{ and }D^{(s)}_{\mathfrak{b}}(t)=\sum_{j}\lceil\lambda_{\mathfrak{b},j}(s)t\rceil\,F_{\mathfrak{b},j},$
+with $F_{\mathfrak{a},j}=\sum_{i=1}^{N}a_{\mathfrak{a},ij}E_{i}\text{ and }F_{\mathfrak{b},j}=\sum_{i=1}^{N}a_{\mathfrak{b},ij}E_{i}.$ Hence,
+$\operatorname{\mathrm{ord}}_{E_{i}}\bigl(\mathfrak{a}(s)_{t}\bigr)=\sum_{j}a_{\mathfrak{a},ij}\,\lceil\lambda_{\mathfrak{a},j}(s)t\rceil\text{ and }\operatorname{\mathrm{ord}}_{E_{i}}\bigl(\mathfrak{b}(s)_{t}\bigr)=\sum_{j}a_{\mathfrak{b},ij}\,\lceil\lambda_{\mathfrak{b},j}(s)t\rceil.$
+Dividing by $t$ and letting $t\to\infty$, we obtain
+$\operatorname{\mathrm{ord}}_{E_{i}}\bigl({\mathfrak{a}_{\bullet}}(s)\bigr)=\sum_{j}a_{\mathfrak{a},ij}\lambda_{\mathfrak{a},j}(s)\text{ and }\operatorname{\mathrm{ord}}_{E_{i}}\bigl({\mathfrak{b}_{\bullet}}(s)\bigr)=\sum_{j}a_{\mathfrak{b},ij}\lambda_{\mathfrak{b},j}(s).$
+Since the coefficient functions $\lambda_{\mathfrak{a},j}(s)$ and $\lambda_{\mathfrak{b},j}(s)$ are piecewise linear, it follows that each function
+[…]
+is piecewise affine linear, and so continuous on each member of a finite decomposition of $S$.
+Step 2: separate the infinite-valued region.
+Refining the above decomposition if necessary, we may assume that for each fixed $i$ the sign of $\operatorname{\mathrm{ord}}_{E_{i}}\bigl({\mathfrak{b}_{\bullet}}(s)\bigr)$ is constant on each piece.
+Since these quantities are nonnegative, on each piece $P$ there is a well-defined subset
+$J_{P}:=\bigl\{\,i\in\{1,\dots,N\}\mid\operatorname{\mathrm{ord}}_{E_{i}}\bigl({\mathfrak{b}_{\bullet}}(s)\bigr)>0\text{ for all }s\in P\,\bigr\}.$
+If on such a piece $P$ there exists $i\notin J_{P}$ such that $\operatorname{\mathrm{ord}}_{E_{i}}\bigl({\mathfrak{a}_{\bullet}}(s)\bigr)>0\text{ for all }s\in P,$ then for every $s\in P$, one has
+$\operatorname{\mathrm{ord}}_{E_{i}}\bigl({\mathfrak{b}_{\bullet}}(s)\bigr)=0\text{ and }\operatorname{\mathrm{ord}}_{E_{i}}\bigl({\mathfrak{a}_{\bullet}}(s)\bigr)>0.$
+By the finite-testing criterion in Theorem 12.2, no finite slope can satisfy the containment inequalities along $E_{i}$. Hence,
+$\mathcal{L}_{{\mathfrak{b}_{\bullet}}(s)}({\mathfrak{a}_{\bullet}}(s))=\infty\text{ for all }s\in P.$
+Step 3: reduction to Theorem 9.9 on the finite-valued region. Now restrict to a piece $P$ on which no such index occurs. Then, for every $i\notin J_{P}$ one has
+$\operatorname{\mathrm{ord}}_{E_{i}}\bigl({\mathfrak{b}_{\bullet}}(s)\bigr)=0\quad\Longrightarrow\quad\operatorname{\mathrm{ord}}_{E_{i}}\bigl({\mathfrak{a}_{\bullet}}(s)\bigr)=0\quad\text{ for all }s\in P.$
+Thus, only the divisors $E_{i}$ with $i\in J_{P}$ contribute to the finite-testing formula of Theorem 12.2.
+For each $s\in P$, define $\alpha_{i}(s):=\operatorname{\mathrm{ord}}_{E_{i}}\bigl({\mathfrak{a}_{\bullet}}(s)\bigr)\text{ and }\beta_{i}(s):=\operatorname{\mathrm{ord}}_{E_{i}}\bigl({\mathfrak{b}_{\bullet}}(s)\bigr)\quad(i\in J_{P}).$ Then,
+$\mathcal{L}_{{\mathfrak{b}_{\bullet}}(s)}({\mathfrak{a}_{\bullet}}(s))=\max_{i\in J_{P}}\frac{\alpha_{i}(s)}{\beta_{i}(s)},$
+and, by construction, each $\alpha_{i}$ and $\beta_{i}$ is continuous on $P$, while $\beta_{i}(s)\in(0,\infty)\text{ for all }i\in J_{P},\ s\in P.$ Therefore, the hypotheses of Theorem 9.9 are satisfied on $P$ with finite candidate set $V_{P}:=\{\operatorname{\mathrm{ord}}_{E_{i}}\mid i\in J_{P}\}.$ Applying Theorem 9.9, we obtain a finite partition of $P$ into locally closed strata on each of which the maximizer set is constant.
+In particular, on each such stratum the function $s\longmapsto\mathcal{L}_{{\mathfrak{b}_{\bullet}}(s)}({\mathfrak{a}_{\bullet}}(s))$ is given by a single valuative ratio whenever the maximizer set is a singleton.
+Taking the common refinement over all pieces $P$ gives the required finite stratification of $S$. ∎
 12.4. Sharpness: why one needs verifiable hypotheses
@@
 Proof of Lemma 4.4.
-Set $I=\mathfrak{a}^{p}$. We use the following one–function statement, which is a specialization of [36, Théorème 7.2] (equivalence of conditions (1),(5),(6)):
+Set $I=\mathfrak{a}^{p}$. We use the following one–function statement, which is a specialization of [40, Théorème 7.2] (equivalence of conditions (1),(5),(6)):
 (LT–S) For $h\in R$,
@@
 $I\mathcal{O}_{Y}(V)=(s).$
-Fix such $V$ and choose $V^{\prime}\subseteq V$.
+Fix such $V$, and choose $V^{\prime}\Subset V$ a relatively compact subset of $V$.
 Step 1: Two–sided comparison. Since $(f_{i}\circ\pi)^{p}\in(s)$, there exist holomorphic functions $u_{i}$ on $V$ such that
@@
 $\|f(\pi(z))\|^{p}\leq A|s(z)|.$
-Since $I=\mathfrak{a}^{p}$ is generated by finitely many degree – $p$ monomials $F_{\nu}=f_{1}^{e_{1,\nu}}\cdots f_{r}^{e_{r,\nu}}$, their pullbacks generate $(s)$.
+Since $I=\mathfrak{a}^{p}$ is generated by finitely many degree – $p$ monomials $F_{\nu}=f_{1}^{e_{1,\nu}}\cdots f_{s}^{e_{s,\nu}}$, their pullbacks generate $(s)$.
 Hence, there exist holomorphic $v_{\nu}$ on $V$ such that
@@
 $A^{-1}\|f\circ\pi\|^{p}\leq|s|\leq B\|f\circ\pi\|^{p}.$
-Step 2: $(LT\text{--}S)$, first direction. Assume $h\in\overline{I}$. By [36, Théorème 7.2], there exists a neighborhood $\tilde{W}$ of $\pi^{-1}(0)$ such that
+Step 2: $(LT\text{--}S)$, first direction. Assume $h\in\overline{I}$. Then, there exists a neighborhood $\tilde{W}$ of $\pi^{-1}(0)$ such that
 $h\circ\pi\in I\mathcal{O}_{Y}\quad\text{on }\tilde{W}$.
@@
 $|h(x)|\leq C^{\prime}\|f(x)\|^{p}.$
-Step 3: $(LT\text{--}S)$, converse direction. Assume $|h(x)|\leq C\|f(x)\|^{p}$ on a neighborhood $U$ of 0. Pulling back,
-$|h\circ\pi|\leq C\|f\circ\pi\|^{p}.$
-On $V^{\prime}\cap\pi^{-1}(U)$, using $\|f\circ\pi\|^{p}\leq A|s|$,
-$|h\circ\pi|\leq CA|s|.$
-Hence, $(h\circ\pi)/s$ is holomorphic and bounded on $(V^{\prime}\cap\pi^{-1}(U))\setminus\{s=0\}$ and extends holomorphically across $\{s=0\}$.
-Thus, $h\circ\pi\in(s)=I\mathcal{O}_{Y}$ near $\pi^{-1}(0)$.
-By Lejeune-Jalabert–Teissier, Théorème 7.2, $h\in\overline{I}$.
+Step 3: $(LT\text{--}S)$, converse direction. Assume $|h(x)|\leq C\|f(x)\|^{p}$ on a neighborhood $U$ of 0. Pulling back, we get
+$|h\circ\pi|\leq C\|f\circ\pi\|^{p}$
+on $\pi^{-1}(U)$.
+Choose finitely many relatively compact open sets $V_{1},\dots,V_{r}\Subset Y$ covering $\pi^{-1}(0)$ such that for each $\alpha$ the invertible ideal $I\mathcal{O}_{Y}|_{V_{\alpha}}$ is principal, say $I\mathcal{O}_{Y}(V_{\alpha})=(s_{\alpha}),$ with $s_{\alpha}\in\mathcal{O}_{Y}(V_{\alpha})$.
+Shrinking if necessary, we may assume that each $V_{\alpha}$ is contained in one of the neighborhoods $V$ used in Step 1.
+Hence, for each $\alpha$, there exists a constant $A_{\alpha}>0$ such that on $V_{\alpha}$ one has $\|f\circ\pi\|^{p}\leq A_{\alpha}|s_{\alpha}|.$ Therefore, on $(V_{\alpha}\cap\pi^{-1}(U))\setminus\{s_{\alpha}=0\}$,
+$|h\circ\pi|\leq C\|f\circ\pi\|^{p}\leq CA_{\alpha}|s_{\alpha}|,$
+so
+$\left|\frac{h\circ\pi}{s_{\alpha}}\right|\leq CA_{\alpha}.$
+Since $Y$ is normal and $s_{\alpha}$ is a local generator of the invertible ideal $I\mathcal{O}_{Y}|_{V_{\alpha}}$, the germ of $s_{\alpha}$ is a non-zero-divisor at every point of $V_{\alpha}$.
+Thus $(h\circ\pi)/s_{\alpha}$ defines a meromorphic function on $V_{\alpha}\cap\pi^{-1}(U)$, and the boundedness above implies that it is holomorphic there.
+Hence
+$h\circ\pi\in s_{\alpha}\mathcal{O}_{Y}(V_{\alpha}\cap\pi^{-1}(U))=I\mathcal{O}_{Y}(V_{\alpha}\cap\pi^{-1}(U))$
+for every $\alpha$.
+Since the sets $V_{\alpha}$ cover $\pi^{-1}(0)$, it follows that $h\circ\pi\in I\mathcal{O}_{Y}$ on a neighborhood of $\pi^{-1}(0)$. Let $E$ be any exceptional divisor of the normalized blow-up. Then
+$\operatorname{\mathrm{ord}}_{E}(h\circ\pi)\geq\operatorname{\mathrm{ord}}_{E}(I\mathcal{O}_{Y}).$
+These divisors correspond to the Rees valuations of $I$, so the above inequalities imply $v(h)\geq v(I)$ for every such valuation $v$. Hence, $h\in\overline{I}$ by Lemma 5.2.
 Step 4: Deduction of the lemma.